### v2

# arXiv:2602.02373v2[gr-qc] 02 Mar 2026

Abstract.
There are three categories of mathematical results concerning quiescent big bang singularities: the derivation of asymptotics in a symmetry class; the construction of spacetimes given initial data on the singularity; and the proof of big bang formation in the absence of symmetries, including the proof of stable big bang formation. In a recent article, the first author demonstrated the existence of developments corresponding to a geometric notion of initial data on a big bang singularity. Moreover, this article, combined with previous articles by the second author, gives a unified and geometric perspective on large classes of seemingly disparate results in the first two categories. Concerning the third category, Oude Groeniger et al recently formulated a general condition on initial data ensuring big bang formation, including curvature blow up. This result, among other things, generalises previous results on stable big bang formation. However, it does not include a statement saying that the solutions induce initial data on the singularity. Here we tie all three categories of results together by demonstrating that the solutions of Oude Groeniger et al induce data on the singularity. However, the results are more general and can potentially be used to derive similar conclusions in other gauges.
1. Introduction
The subject of this article is quiescent big bang singularities. To put this topic into context, it is natural to return to the work of Belinskiǐ, Khalatnikov and Lifschitz (BKL), see, e.g., [11, 12] as well as more modern refinements, e.g., [16, 17, 25, 26]. Very crudely, the idea is to study the asymptotics in the direction of a big bang singularity using a foliation by spacelike hypersurfaces, the mean curvatures of which diverge; i.e., if $K$ denotes the Weingarten map of the leaves of the foliation, $\theta:=\mathrm{tr}K$ diverges. In order to extract a finite limit, it is therefore natural to expansion normalise by dividing by the mean curvature: $\mathcal{K}:=K/\theta$. We refer to $\mathcal{K}$ as the expansion normalised Weingarten map. One suggestion of BKL is that, along a causal curve going into the singularity, the eigenvalues of $\mathcal{K}$ should oscillate according to a chaotic one-dimensional map referred to as the Kasner map or BKL map. This is referred to as the oscillatory setting. However, for certain matter models and for certain symmetry classes, the oscillations can be neutralised and the resulting behaviour is quiescent in the sense that the eigenvalues converge. Oscillatory solutions are very interesting, but, to the best of our knowledge, there are only results in the spatially homogeneous setting; see, e.g., [54, 36, 37, 6, 31, 32, 13, 7]. There is a vast literature on spatially homogeneous solutions, and most of the results concern the quiescent setting. There are also some results in the case of $\mathbb{T}^{3}$-Gowdy symmetry; see, e.g., [14, 38, 40]. However, going beyond these results has turned out to be difficult. One way to avoid this difficulty is to turn the problem around, to start with “data at the singularity” and to prove that there are corresponding solutions. Over the last 20–30 years, a substantial body of work on this topic has appeared; see, e.g., [29, 27, 34, 4, 28, 15, 53, 1, 30, 2, 21, 5]. Unfortunately, these results have certain deficiencies. One deficiency is that many of the results are in the real analytic setting, a regularity class incompatible with finite speed of propagation, a central notion in general relativity. A second deficiency is that the notions of initial data in the different results are strongly connected to the particular choice of gauge. It is therefore not easy to see how the conditions of the different articles relate. Finally, viewing the constructed spacetimes as solutions to the initial value problem given initial data on the singularity, it would be highly desirable to have a geometric formulation of the initial value problem; isometric data on the singularity should give rise to isometric developments etc. In [45], an article based on [42, 44], the second author developed such a geometric notion of initial data. In [23], the first author demonstrated that this notion of initial data can be used to formulate a geometric initial value problem for the Einstein–nonlinear scalar field equations with data on the singularity. In particular, given initial data there is, up to isometry, a unique corresponding development, and isometric initial data on the singularity give rise to isometric developments. We expect this result to give a unified perspective on the results starting with initial data on the singularity. In the spatially homogeneous Bianchi class A setting, this is verified in [46, 47]. In the $\mathbb{T}^{3}$–Gowdy setting, it is verified in [46]. In particular, this means that the notion of initial data used in [29, 34] is a special case of the initial data used in [45, 23]. That the notion of initial data used in [4] is a special case of the notion introduced in [45] is verified in [45]. That the notion of initial data used in [30, 21, 5] is a special case of the notion introduced in [45] is also verified in [45].
Studying the stability of big bang formation is an alternative approach which has been very active in the last 5–10 years. The results began with the work of Rodnianski and Speck, demonstrating stability of spatially homogeneous, isotropic and spatially flat solutions in the Einstein-scalar field and the Einstein-stiff fluid setting; see [49, 50]. Speck demonstrates a similar result in the case of $\mathbb{S}^{3}$-spatial topology; see [52]. In [51], the authors consider the vacuum equations in higher dimensions in situations with intermediate anisotropy; the conditions are more general than requiring proximity to isotropy, but they are still quite far from covering the full subcritical range in which stability is expected. In [19], Fajman and Urban prove a stability result in the case of spatial local homogeneity and isotropy and negative spatial curvature. This result was later generalised to allow the presence of Vlasov matter and any sign of the sectional curvature in [20]. Finally, Fournodavlos, Rodnianski and Speck demonstrate stable big bang formation for the full subcritical range of Kasner-scalar field spacetimes in [22]. This is a very important result for many reasons. First, because it covers the maximal range in which one can obtain stability. Second, because the authors introduce a new formulation of the equations using a Fermi-Walker propagated frame which has turned out to be very fruitful. The extension of this result to stability of the big bang singularity for the Einstein-Maxwell-scalar field-Vlasov system in the full strong subcritical range is discussed in [3].
All of the stability results mentioned so far are obtained using constant mean curvature (CMC) foliations. Such foliations are natural for several reasons. To begin with, they are uniquely determined by the spacetime for large classes of matter models. Moreover, they synchronise the singularity, something which is not to be expected for most gauges. On the other hand, CMC foliations have deficiencies. In particular, the corresponding gauge fixed system includes elliptic equations. This means that it is not possible to localise the analysis in space. This is contrary to the expected spatial localisation of the causal structure. Moreover, thinking ahead, going beyond stability of spatially homogeneous solutions, quite violent spatial variations are to be expected; this can be seen already in the spikes appearing in the $\mathbb{T}^{3}$-Gowdy symmetric setting, see, e.g., [35]. In order to obtain a clear picture of the behaviour of solutions near the big bang, it can therefore be expected to be necessary to localise the analysis. In the spatially homogeneous, isotropic and spatially flat setting, localisation is achieved by Beyer and Oliynyk in [8], a result which is generalised to include a fluid in [9]. This requires the development of a new, local, gauge. The authors achieve this by using the scalar field as a time coordinate. In [10], Beyer, Oliynyk and Zheng generalise the result to cover the full subcritical range in $3$-dimensions, a result which is generalised to higher dimensions in [55]. As in [22], the authors use a Fermi-Walker transported frame in order to obtain their results.
In spite of the importance of the above results, they have one fundamental weakness, namely that they only demonstrate stability of specific spatially locally homogeneous solutions. Due to the ideas of BKL and the results [4, 23] concerning data on the singularity, it is natural to expect that solutions should exhibit quiescent asymptotics more generally. Moreover, due to the perspective developed in [42, 44, 45], it is natural to expect that expansion normalised initial data should play an important role in phrasing the corresponding condition. In fact, in [33], Oude Groeniger, Petersen and the second author provide a general condition on initial data, without any reference to a background solution (spatially homogeneous or otherwise) such that the corresponding solutions exhibit big bang formation with a crushing CMC foliation and curvature blow up. Moreover, the result gives partial asymptotics, corresponding to some of the data on the singularity, as formulated in [45] or [23]. In addition, the results of [33] not only imply the previous stability results in the Einstein–scalar field setting with CMC foliations, they substantially extend them. Since [33] concerns the Einstein–nonlinear scalar field setting, the results also yield future and past global nonlinear stability results for large classes of spatially locally homogeneous solutions.
Considering the above circle of ideas, the only thing missing to tie the results given data on the singularity together with the stability results is a proof that the solutions constructed in [33] induce data on the singularity in the sense of [45] or [23]. In this article, we provide such a proof. Moreover, we do so by providing tools we expect to be applicable more generally. More specifically, we prove that on the basis of appropriate $C^{2}$-assumptions concerning solutions to the FRS equations used in [33], it is possible to deduce $C^{k}$-conclusions for any $k$. We also prove that, given an appropriate foliation (not necessarily a CMC foliation) and appropriate assumptions concerning a Fermi-Walker propagated orthonormal frame of the leaves of the foliation as well as concerning basic geometric quantities expressed with respect this Fermi-Walker propagated orthonormal frame, it follows that the solution induces data on the singularity in the sense of [45] or [23].
1.1. The Einstein–nonlinear scalar field equations
In this article, we are interested in the Einstein–nonlinear scalar field equations. They are determined once we specify a potential; i.e., a function $V\in C^{\infty}(\mathbb{R})$. A solution consists of a time oriented Lorentz manifold $(M,g)$, i.e. a spacetime, and a function $\varphi\in C^{\infty}(M)$, referred to as the scalar field, satisfying
(1a) $\mathrm{Ric}_{g}-\tfrac{1}{2}\mathrm{Scal}_{g}g$ $=T,$
(1b) $\Box_{g}\varphi$ $=V^{\prime}\circ\varphi.$
Here $\mathrm{Ric}_{g}$ and $\mathrm{Scal}_{g}$ denote the Ricci and scalar curvatures of $g$ respectively, $\Box_{g}$ is the wave operator associated with $g$, and
(2) $T={\mathrm{d}}\varphi\otimes{\mathrm{d}}\varphi-\big(\tfrac{1}{2}|{\mathrm{d}}\varphi|_{g}^{2}+V\circ\varphi\big)g$
is the energy-momentum tensor of $\varphi$. Note that adding a cosmological constant term to the left hand side of (1a) would not make the system more general, as the cosmological constant can be absorbed in the potential.
There is a Cauchy problem associated with (1).
Definition 1.
Initial data for the Einstein–nonlinear scalar field equations (1) consist of a Riemannian manifold $(\Sigma,h)$, a symmetric covariant $2$-tensor field $k$ and two smooth functions $\phi$ and $\psi$ satisfying
(3a) $\mathrm{Scal}_{h}-|k|^{2}_{h}+(\mathrm{tr}_{h}k)^{2}$ $=\psi^{2}+|{\mathrm{d}}\phi|_{h}^{2}+2V\circ\phi,$
(3b) $\mathrm{div}_{h}k-{\mathrm{d}}\mathrm{tr}_{h}k$ $=\psi{\mathrm{d}}\phi.$
If $\theta:=\mathrm{tr}_{h}k$ is constant, the initial data are said to be constant mean curvature (CMC) initial data. Given initial data for the Einstein–nonlinear scalar field equations, say $(\Sigma,h,k,\phi,\psi)$ a corresponding development is a solution $(M,g,\varphi)$ to (1), and an embedding $i:\Sigma\rightarrow M$ such that the following hold. First, $i^{*}g=h$, so that, in particular, $i(\Sigma)$ is a spacelike hypersurface in $(M,g)$. Second, if $\bar{k}$ is the second fundamental form induced on $i(\Sigma)$ and $U$ is the future directed unit normal to $i(\Sigma)$, then
$i^{*}\bar{k}=k,\ \ \ \varphi\circ i=\phi,\ \ \ (U\varphi)\circ i=\psi.$
Remark 2.
Equation (3a) is called the Hamiltonian constraint equation and (3b) is called the momentum constraint equation.
Remark 3.
Due to arguments going back to the work of Yvonne Choquet-Bruhat and Robert Geroch [24], there is a unique maximal globally hyperbolic development associated with initial data; see also [41, Chapter 23] for a textbook presentation.
Remark 4.
Let $(M,g,\varphi)$ be a solution to (1) and let $\Sigma$ be a spacelike hypersurface in $(M,g)$ with future directed unit normal $U$. Define $h$ and $k$ to be the associated induced first and second fundamental forms. Let, moreover, $\phi:=\varphi|_{\Sigma}$ and $\psi:=U(\varphi)|_{\Sigma}$. Then $(\Sigma,h,k,\phi,\psi)$ are initial data for the Einstein–nonlinear scalar field equations; see, e.g., [39, Chapter 13].
Given initial data such that the associated mean curvature is strictly positive, it is possible to define associated expansion normalised initial data, a concept which is of central importance in the definition of initial data on the singularity.
Definition 5.
Let $(\Sigma,h,k,\phi,\psi)$ be initial data for the Einstein–nonlinear scalar field equations and $\theta:=\mathrm{tr}_{h}k$. Assume that $\theta>0$. Let $K$ be the $(1,1)$-tensor field obtained by raising one index in $k$ using $h$. Then $\mathcal{K}:=\theta^{-1}K$ is called the expansion normalised Weingarten map. Define
$\mathcal{H}(X,Y):=h(\theta^{\mathcal{K}}X,\theta^{\mathcal{K}}Y),\ \ \ \Psi:=\theta^{-1}\psi,\ \ \ \Phi:=\phi+\Psi\ln\theta,$
for all $X,Y\in T_{p}\Sigma$ and all $p\in\Sigma$, where
$\theta^{\mathcal{K}}X:=\textstyle\sum_{m=0}^{\infty}\tfrac{(\ln\theta)^{m}}{m!}\mathcal{K}^{m}X.$
Then $(\Sigma,\mathcal{H},\mathcal{K},\Phi,\Psi)$ are referred to as expansion normalised initial data associated with the initial data $(\Sigma,h,k,\phi,\psi)$.
1.2. Initial data on the singularity
Next, we define the notion of initial data on the singularity.
Definition 6.
For integers $\ell\geq 1$ and $n\geq 3$, let $(\Sigma,\mathring{\mathcal{H}})$ be an $n$-dimensional $C^{\ell}$ Riemannian manifold, $\mathring{\mathcal{K}}$ a $C^{\ell}$ $(1,1)$-tensor field on $\Sigma$ and $\mathring{\Phi},\mathring{\Psi}\in C^{\ell}(\Sigma)$. Assume that
  1. (1)
[…]
  2. (2)
    $\mathrm{tr}\mathring{\mathcal{K}}^{2}+\mathring{\Psi}^{2}=1$ and $\mathrm{div}_{\mathring{\mathcal{H}}}\mathring{\mathcal{K}}=\mathring{\Psi}{\mathrm{d}}\mathring{\Phi}$.
  3. (3)
    The eigenvalues $\mathring{p}_{1},\ldots,\mathring{p}_{n}$ of $\mathring{\mathcal{K}}$ are everywhere distinct, and satisfy
    (4) $\mathring{p}_{I}+\mathring{p}_{J}-\mathring{p}_{K}<1$
    for all $I,J$ and $K$ such that $I\neq J$.
Then $(\Sigma,\mathring{\mathcal{H}},\mathring{\mathcal{K}},\mathring{\Phi},\mathring{\Psi})$ are $C^{\ell}$ robust nondegenerate quiescent initial data on the singularity for the Einstein–nonlinear scalar field equations. If they are $C^{\ell}$ for any $\ell\in\mathbb{N}$, they are said to be smooth.
Remark 7.
It is possible to omit the nondegeneracy requirement; i.e., the requirement that the eigenvalues be distinct. However, we are mainly interested in the nondegenerate setting in this article. It is also possible to omit the requirement that (4) hold. However, it is then necessary to introduce separate conditions on the structure coefficients associated with an eigenframe of $\mathring{\mathcal{K}}$; see, e.g., [45, Definition 10, p. 9]. Moreover, the solutions inducing initial data violating (4) are expected to be unstable. Since we are here interested in the asymptotics induced by solutions arising in stability results, such a generalisation is therefore not of great interest here.
In order to relate a solution as in Subsection 1.1 to initial data on the singularity in the sense of Definition 6, we first need to introduce a foliation. Let $(M,g,\varphi)$ be a globally hyperbolic solution to the Einstein–nonlinear scalar field equations. Then $M$ is diffeomorphic to $\Sigma\times\mathcal{I}$ for some $n$-manifold $\Sigma$ and some open interval $\mathcal{I}=(t_{-},t_{+})$, where $\Sigma_{t}:=\Sigma\times\{t\}$ are spacelike Cauchy hypersurfaces. When speaking of big bang singularities, we have crushing singularities in mind. In other words, we here assume the mean curvature of the leaves $\Sigma_{t}$ of the foliation to diverge uniformly to $\infty$ as $t\downarrow t_{-}$. The metric $g$ and scalar field $\varphi$ induce initial data, cf. Remark 4, on each leaf: $(\Sigma,h(t),k(t),\phi(t),\psi(t))$; we think of the data as being defined on $\Sigma$. Moreover, these data induce expansion normalised initital data $(\Sigma,\mathcal{H}(t),\mathcal{K}(t),\Phi(t),\Psi(t))$ as explained in Definition 5. We say that the solution induces initial data on the singularity along the foliation $\{\Sigma_{t}\}_{t\in\mathcal{I}}$ if there are $t_{0}\in\mathcal{I}$; constants $\delta>0$ and $C_{\ell}$ for $\ell\in\mathbb{N}_{0}$; and robust quiescent initial data on the singularity for the Einstein–nonlinear scalar field equations, say $(\Sigma,\mathring{\mathcal{H}},\mathring{\mathcal{K}},\mathring{\Phi},\mathring{\Psi})$, such that
(5a) $\|\theta^{\delta}\bar{D}^{\ell}(\mathcal{H}(t)-\mathring{\mathcal{H}})\|_{C^{0}(\Sigma)}+\|\theta^{\delta}\bar{D}^{\ell}(\mathcal{K}(t)-\mathring{\mathcal{K}})\|_{C^{0}(\Sigma)}$ $\leq C_{\ell},$
(5b) $\|\theta^{\delta}\bar{D}^{\ell}(\Phi(t)-\mathring{\Phi})\|_{C^{0}(\Sigma)}+\|\theta^{\delta}\bar{D}^{\ell}(\Psi(t)-\mathring{\Psi})\|_{C^{0}(\Sigma)}$ $\leq C_{\ell}$
for all $t\leq t_{0}$ and all $\ell\in\mathbb{N}_{0}$. Here $\bar{D}$ is the Levi-Civita connection associated with a fixed reference Riemannian metric on $\Sigma$, say $h_{\mathrm{ref}}$, and the $C^{0}$-norm of a tensor field $T$ is defined as follows:
$\|T\|_{C^{0}(\Sigma)}:=\sup_{x\in\Sigma}|T(x)|_{h_{\mathrm{ref}}}.$
In case $\Sigma$ is non-compact, one could relax (5) to the requirement that for each compact subset $K$ of $\Sigma$, there are constants $\delta$ and $C_{\ell}$ such that (5) holds with $\Sigma$ replaced by $K$.
In $3+1$-dimensions, the step from data on the singularity to corresponding developments is discussed in [23]. In particular, due to [23, Theorem 1.21, p. 10], initial data on the singularity give rise to unique (up to isometry) maximal globally hyperbolic developments with a local Gaussian foliation close to the singularity. Next, we turn to the opposite direction. In particular, we prove that the general condition on initial data introduced in [33], ensuring that the corresponding maximal globally hyperbolic developments have crushing CMC singularities with curvature blow up, in fact yields solutions inducing data on the singularity along the CMC foliation.
1.3. The main result
In this subsection, we state the main result, namely that the solutions constructed in [33] induce data on the singularity. However, to begin with, we need to introduce some terminology.
Definition 8.
Let $\mathfrak{I}=(\Sigma,h,k,\phi,\psi)$ be CMC initial data with mean curvature $\theta>0$. Let $\mathcal{K}$ be the associated expansion normalised Weingarten map, see Definition 5, and let ${\sigma_{p}}\in(0,1)$. Then $\mathfrak{I}$ and the eigenvalues $p_{1},\ldots,p_{n}$ of $\mathcal{K}$ are said to be $\sigma_{p}$}-admissible if they satisfy
(6) $p_{I}+p_{J}-p_{K}<1-{\sigma_{p}},$
for all $I,J,K$, such that $I\neq J$.
Concerning the potential, we make the following assumption.
Definition 9.
Fix $\sigma_{V}\in(0,1)$. If $V\in C^{\infty}(\mathbb{R})$ has the property that for each $k\in\mathbb{N}_{0}$, there is a constant $c_{k}>0$ such that
(7) ${\sum}_{\ell\leq k}|V^{(\ell)}(x)|\leq c_{k}e^{2(1-\sigma_{V})|x|}$
for all $x\in\mathbb{R}$, then $V$ is said to be a $\sigma_{V}$-admissible potential.
Remark 10.
In comparison with [33, Definition 1, p. 5], we have here omitted the requirement that $V$ be non-negative. We include this requirement explicitly in Theorem 11 instead.
Theorem 11 (The main theorem).
Fix $3\leq n\in\mathbb{N}$ and admissibility thresholds $\sigma_{V}$, ${\sigma_{p}}\in(0,1)$. Let $(\Sigma,h_{\mathrm{ref}})$ be a closed Riemannian manifold of dimension $n$ and let $V\in C^{\infty}(\mathbb{R})$ be a non-negative $\sigma_{V}$–admissible potential. Then there is a $\kappa\in\mathbb{N}$, depending only on $n$, $\sigma_{V}$ and $\sigma_{p}$}, such that the following holds. For any $\zeta_{0}>0$, there is a $\zeta_{1}>0$ such that:
If $\mathfrak{I}$ are $\sigma_{p}$}-admissible CMC initial data on $\Sigma$ for the Einstein–nonlinear scalar field equations with potential $V$, such that the associated expansion-normalized initial data $(\Sigma,\mathcal{H}_{0},\mathcal{K}_{0},\Phi_{0},\Psi_{0})$ satisfy
(8) $\|\mathcal{H}^{-1}_{0}\|_{C^{0}_{\mathrm{ref}}(\Sigma)}+\|\mathcal{H}_{0}\|_{H^{\kappa}_{\mathrm{ref}}(\Sigma)}+\|\mathcal{K}_{0}\|_{H^{\kappa}_{\mathrm{ref}}(\Sigma)}+\|\Phi_{0}\|_{H^{\kappa}_{\mathrm{ref}}(\Sigma)}+\|\Psi_{0}\|_{H^{\kappa}_{\mathrm{ref}}(\Sigma)}<\zeta_{0};$
$|p_{I}-p_{J}|>\zeta_{0}^{-1}$ for $I\neq J$, where $p_{I}$ are the eigenvalues of $\mathcal{K}_{0}$; and the mean curvature satisfies $\theta_{0}>\zeta_{1}$, then the maximal globally hyperbolic development of $\mathfrak{I}$, say $(M,g,\varphi)$, with associated embedding $\iota:\Sigma\hookrightarrow M$, has a past crushing big bang singularity in the following sense:
CMC foliation: There is a diffeomorphism $\Xi$ from $(0,t_{0}]\times\Sigma$ to $J^{-}\left(\iota(\Sigma)\right)$, i.e. the causal past of the Cauchy hypersurface $\iota(\Sigma)$, such that $\Xi(\Sigma_{t_{0}})=\iota(\Sigma)$ and the hypersurfaces $\Xi(\Sigma_{t})\subset M$ are spacelike Cauchy hypersurfaces with constant mean curvature $\theta(t)=\frac{1}{t}$, for each $t\in(0,t_{0}]$.
Asymptotic data: There are smooth robust nondegenerate quiescent initial data on the singularity for the Einstein–nonlinear scalar field equations, say $(\Sigma,\mathring{\mathcal{H}},\mathring{\mathcal{K}},\mathring{\Phi},\mathring{\Psi})$, and constants $\delta>0$ and $C_{\ell}$ for $\ell\in\mathbb{N}_{0}$, such that
(9) $\|\mathcal{H}(t)-\mathring{\mathcal{H}}\|_{C^{\ell}_{\mathrm{ref}}(\Sigma)}+\|\mathcal{K}(t)-\mathring{\mathcal{K}}\|_{C^{\ell}_{\mathrm{ref}}(\Sigma)}+\|\Phi(t)-\mathring{\Phi}\|_{C^{\ell}_{\mathrm{ref}}(\Sigma)}+\|\Psi(t)-\mathring{\Psi}\|_{C^{\ell}_{\mathrm{ref}}(\Sigma)}\leq C_{\ell}t^{\delta}$
for all $t\leq t_{0}$ and all $\ell\in\mathbb{N}_{0}$, where $\Sigma,\mathcal{H}(t),\mathcal{K}(t),\Phi(t),\Psi(t)]$ are the expansion normalised initial data induced on $\Sigma_{t}$.
Curvature blow-up: Let $\mathring{p}_{I}$ denote the eigenvalues of $\mathring{\mathcal{K}}$. Then there are constants $C_{\ell}>0$, $\ell\in\mathbb{N}_{0}$, such that $\mathfrak{R}_{g}:=\mathrm{Ric}_{g,\mu\nu}\mathrm{Ric}_{g}^{\mu\nu}$ and $\mathfrak{K}_{g}:=\mathrm{Riem}_{g,\mu\nu\xi\rho}\mathrm{Riem}_{g}^{\mu\nu\xi\rho}$ satisfy
(10a) $\big\|t^{4}\mathfrak{K}_{g}(t,\cdot)-4\big[\textstyle{\sum}_{I}\mathring{p}_{I}^{2}(1-\mathring{p}_{I}^{2})+\textstyle{\sum}_{I<J}\mathring{p}_{I}^{2}\mathring{p}_{J}^{2}\big]\big\|_{C^{\ell}_{\mathrm{ref}}(\Sigma)}\leq C_{\ell}t^{\delta},$
(10b) $\|t^{4}\mathfrak{R}_{g}(t,\cdot)-\mathring{\Psi}^{4}\|_{C^{\ell}_{\mathrm{ref}}(\Sigma)}\leq C_{\ell}t^{\delta}$
for all $t\in(0,t_{0}]$ and all $\ell\in\mathbb{N}_{0}$, so that $(M,g)$ is $C^{2}$ past inextendible. Moreover, every past directed causal geodesic in $M$ is incomplete and $\mathfrak{K}_{g}$ blows up along every past inextendible causal curve.
Remark 12.
The constant $\kappa$ is given as follows. In the main theorem of [33], i.e. [33, Theorem 12, pp. 6–7], the constants $k_{0}$ and $k_{1}$ are introduced. Choosing $k_{0}$ by demanding that equality hold in [33, (12a), p. 6]; and choosing $k_{1}$ to be the smallest integer such that [33, (12b), p. 6] holds for this choice of $k_{0}$, the integer $\kappa$ is given by $\kappa=k_{1}+2$.
Remark 13.
The Sobolev spaces $H^{\ell}_{\mathrm{ref}}$ and $C^{\ell}$ spaces $C^{\ell}_{\mathrm{ref}}$ in (8), (9) and (10) are defined using the reference metric $h_{\mathrm{ref}}$ and the associated Levi–Civita connection. In most of the article we use the $C^{\ell}$ and Sobolev spaces introduced in Subsection A.4.
Remark 14.
The requirement that $V$ be non–negative can be relaxed; cf. [33, Remark 16, p. 7].
Remark 15.
In [33], the stability of quiescent model solutions, a notion introduced in [33, Definition 30, p. 11], is also discussed; cf. [33, Theorem 33, p. 12] and [33, Subsection 1.6, pp. 13–19]. The background solutions to the Einstein-scalar field equations considered in, e.g., [52, 22, 19, 20] are examples of quiescent model solutions, but the notion is more general than this. There are quiescent model solutions inducing degenerate data on the singularity; i.e., data such that the $\mathring{p}_{I}$ are not all distinct. For this reason, the solutions obtained by appealing to the stability results are not necessarily such that Theorem 11 applies. However, even in the degenerate setting, one does obtain conclusions. Consider a solution obtained by appealing to [33, Theorem 33, p. 12]. Then the first part of the proof of Theorem 11, which simply consists of an application of Theorem 19, still applies. Next, assume that at some point, say $x\in\Sigma$, the eigenvalues $\mathring{p}_{I}(x)$ are all distinct. Then there is a closed ball, say $\bar{B}_{R}(x)$, of radius $R>0$ with respect to the reference metric $h_{\mathrm{ref}}$ such that, for every $y\in\bar{B}_{R}(x)$, the eigenvalues $\mathring{p}_{I}(y)$ are all distinct. Choosing $T>0$, $\delta>0$ and $r>0$ small enough, it can then be ensured that $g$ and $\varphi$, restricted to $M_{\mathrm{loc}}:=\bar{B}_{R}(x)\times(0,T)$, constitute a solution to the Einstein–nonlinear scalar field equations satisfying the $(\delta,n,k_{0},r)$-assumptions for any $k_{0}$. This means that we can apply Theorem 24 in order to obtain smooth initial data on the singularity in the ball $\bar{B}_{R}(x)$. Moreover, (9) holds with $\Sigma$ replaced by $\bar{B}_{R}(x)$.
Proof.
The proof is to be found in Section 4. ∎
The statement of the result may seem a bit strange at first sight. However, from the point of view of initial data on the singularity, it is quite natural. In fact, consider a solution that induces initial data on the singularity along a CMC foliation. Then it is clear that for any $k\in\mathbb{N}$, the expansion normalised initial data on the leaves of the CMC foliation are uniformly bounded in $H^{k}_{\mathrm{ref}}$ all the way to the singularity; i.e., as the mean curvature diverges to infinity. It is therefore natural to first fix a bound on the size of the expansion normalised initial data, without imposing any restriction on how large the bound may be, and then to require a lower bound on the mean curvature.
1.4. The FRS equations
One topic of interest in this article is solutions to the FRS equations, introduced in [22]. The purpose of the present section is to recall the slight reformulation of these equations introduced in [33, Theorem 12, pp. 5–6]. These equations arise from the Einstein–nonlinear scalar field equations by assuming a CMC foliation, a vanishing shift vector field, and using a Fermi-Walker transported frame. The underlying assumptions are the following. Let $(\Sigma,h_{\mathrm{ref}})$ be a closed parallelisable Riemannian manifold of dimension $n$; let $\{E_{i}\}_{i=1}^{n}$ be a smooth global orthonormal frame with respect to $h_{\mathrm{ref}}$ with dual frame $\{\eta^{i}\}_{i=1}^{n}$; and let $V\in C^{\infty}(\mathbb{R})$. The manifold on which we wish to solve the equations is of the form $M:=\Sigma\times\mathcal{I}$ for some open interval $\mathcal{I}\subset(0,\infty)$ and the basic variables are $N$, $e_{I}^{i}$, $\omega^{I}_{i}$, $\gamma_{IJK}$, $k_{IJ}$ and $\varphi$. Moreover, we assume $N>0$ on $M$. As clarified in the statement of [33, Proposition 70, pp. 23–24], the geometric interpretation of these variables is the following. Let $e_{0}:=N^{-1}\partial_{t}$, where $\partial_{t}$ signifies differentiation with respect to the second variable in $M\times\mathcal{I}$; $e_{I}:=e_{I}^{i}E_{i}$; and $\omega^{I}:=\omega^{I}_{i}\eta^{i}$. Then $\{e_{\alpha}\}_{\alpha=0}^{n}$ is a time oriented orthonormal frame for a Lorentz metric $g$ on $M$ and if $h$ denotes the induced metric on the constant-$t$ hypersurfaces $\Sigma_{t}:=\Sigma\times\{t\}$, then $\{e_{I}\}_{I=1}^{n}$ is an orthonormal frame with respect to $h$ with dual frame $\{\omega^{I}\}_{I=1}^{n}$. Moreover,
$k:=k_{IJ}\omega^{I}\otimes\omega^{J},\ \ \ \gamma_{IJK}=\omega^{K}([e_{I},e_{J}])$
are the induced second fundamental form on $\Sigma_{t}$ and the structure coefficients of the frame $\{e_{I}\}_{I=1}^{n}$, respectively, and the mean curvature satisfies $k_{II}=t^{-1}$ (here and below, we sum over repeated capital Latin indices, even if both are subscripts). Given the above underlying assumptions, the equations, see [33, Proposition 70, pp. 23–24], are the following, using the conventions introduced in Subsection A.2:
The evolution equations for the frame and co-frame:
(11) $e_{0}(e_{I}^{i})$ $=-k_{IJ}e_{J}^{i},$
(12) $e_{0}(\omega^{I}_{i})$ $=k_{IJ}\omega^{J}_{i}.$
The evolution equations for $\gamma$ and $k$:
(13) $\begin{split}e_{0}\gamma_{IJK}&=-2N^{-1}e_{[I}(Nk_{J]K})-k_{IL}\gamma_{LJK}-k_{JL}\gamma_{ILK}+k_{KL}\gamma_{IJL},\end{split}$
(14) $\begin{split}e_{0}k_{IJ}&=N^{-1}e_{(I}e_{J)}(N)-N^{-1}e_{K}(N\gamma_{K(IJ)})-e_{(I}(\gamma_{J)KK})-t^{-1}k_{IJ}\\ &\quad+\gamma_{KLL}\gamma_{K(IJ)}+\gamma_{I(KL)}\gamma_{J(KL)}-\tfrac{1}{4}\gamma_{KLI}\gamma_{KLJ}\\ &\quad+e_{I}(\varphi)e_{J}(\varphi)+\tfrac{2}{n-1}(V\circ\varphi)\delta_{IJ}.\end{split}$
The evolution equations for the derivatives of the scalar field:
(15) $e_{0}\left(e_{I}\varphi\right)$ $=N^{-1}e_{I}\left(Ne_{0}\varphi\right)-k_{IJ}e_{J}(\varphi),$
(16) $e_{0}\left(e_{0}(\varphi)\right)$ $=e_{I}\left(e_{I}(\varphi)\right)-t^{-1}e_{0}(\varphi)+N^{-1}e_{I}(N)e_{I}(\varphi)-\gamma_{JII}e_{J}(\varphi)-V^{\prime}\circ\varphi.$
The Hamiltonian constraint equation:
(17) $\begin{split}2e_{I}(\gamma_{IJJ})&-\tfrac{1}{4}\gamma_{IJK}(\gamma_{IJK}+2\gamma_{IKJ})-\gamma_{IJJ}\gamma_{IKK}-k_{IJ}k_{IJ}+t^{-2}\\ &=(e_{0}\varphi)^{2}+e_{I}(\varphi)e_{I}(\varphi)+2V\circ\varphi.\end{split}$
The momentum constraint equation:
(18) $e_{I}k_{IJ}=\gamma_{LII}k_{LJ}+\gamma_{IJL}k_{IL}+e_{0}(\varphi)e_{J}(\varphi).$
The lapse equation:
(19) $\begin{split}e_{I}e_{I}(N)&=t^{-2}(N-1)+\gamma_{JII}e_{J}(N)-\big(e_{I}(\varphi)e_{I}(\varphi)+\tfrac{2n}{n-1}V\circ\varphi\big)N\\ &\quad+\left(2e_{I}(\gamma_{IJJ})-\tfrac{1}{4}\gamma_{IJK}(\gamma_{IJK}+2\gamma_{IKJ})-\gamma_{IJJ}\gamma_{IKK}\right)N.\end{split}$
It is convenient to recall that combining (17) and (19) yields [33, (85), p. 32]:
(20) $\begin{split}e_{I}e_{I}(N)&=t^{-2}(N-1)+\gamma_{JII}e_{J}(N)\\ &\quad-\big(t^{-2}-k_{IJ}k_{IJ}-e_{0}(\varphi)e_{0}(\varphi)+\tfrac{2}{n-1}V\circ\varphi\big)N.\end{split}$
In the above equations and in what follows, it is useful to recall [33, Lemma 72, p. 25]:
Lemma 16 (Lemma 70, p. 25, [33]).
Let $(\Sigma,h)$ be a Riemannian manifold with a smooth global orthonormal frame $\{e_{I}\}_{I=1}^{n}$. Then the Ricci and scalar curvature are given by
(21) $\begin{split}\mathrm{Ric}_{h}(e_{I},e_{J})&=e_{K}(\gamma_{K(IJ)})+e_{(I}(\gamma_{J)KK})-\gamma_{I(KL)}\gamma_{J(KL)}\\ &\quad+\tfrac{1}{4}\gamma_{KLI}\gamma_{KLJ}-\gamma_{KLL}\gamma_{K(IJ)},\end{split}$
(22) $\mathrm{Scal}_{h}$ $=2e_{I}(\gamma_{IJJ})-\tfrac{1}{4}\gamma_{IJK}(\gamma_{IJK}+2\gamma_{IKJ})-\gamma_{IJJ}\gamma_{IKK},$
where $\gamma_{IJK}=h([e_{I},e_{J}],e_{K})$.
When deriving energy estimates, it is, in the end, useful to consider an energy associated with
(23) $a_{I}:=\gamma_{IJJ}.$
Interestingly, this quantity satisfies an evolution equation which is in some respects better than that of $\gamma$. In fact, due to (13) and (18),
$\begin{split}e_{0}a_{I}=&-2N^{-1}e_{[I}(Nk_{J]J})-k_{IL}\gamma_{LJJ}-k_{JL}\gamma_{ILJ}+k_{JL}\gamma_{IJL}\\ =&-N^{-1}t^{-1}e_{I}N+N^{-1}e_{J}(N)k_{IJ}+e_{J}(k_{JI})-k_{IL}\gamma_{LJJ}\\ =&-N^{-1}t^{-1}e_{I}N+N^{-1}e_{J}(N)k_{IJ}-k_{IL}a_{L}+a_{L}k_{LI}+\gamma_{JIL}k_{JL}+e_{0}(\varphi)e_{I}(\varphi).\end{split}$
In other words,
(24) $\partial_{t}a_{I}=-t^{-1}e_{I}(N)+e_{J}(N)k_{IJ}+N\gamma_{JIL}k_{JL}+Ne_{0}(\varphi)e_{I}(\varphi).$
1.5. Higher order derivative estimates for solutions to the FRS equations
Even though our main application in the present paper is to derive complete asymptotics for the solutions constructed in [33], it is of interest to derive energy estimates using minimal assumptions. In this way, the result might be useful in other contexts. The main assumptions are the following.
Definition 17.
Consider a solution as described in Subsection 1.4. In particular, the mean curvature of $\Sigma_{t}:=\Sigma\times\{t\}$ is $\theta=t^{-1}$. Assume that the existence interval includes $(0,t_{0}]$, where $t_{0}\in(0,1]$. Next, assume that there is an $\varepsilon>0$ and a constant $C$ such that
(25a) $t^{1-3\varepsilon}\|e\|_{C^{1}(\Sigma_{t})}+t^{1-3\varepsilon}\|\omega\|_{C^{1}(\Sigma_{t})}+t^{1-2\varepsilon}\|\gamma\|_{C^{1}(\Sigma_{t})}+t^{1-2\varepsilon}\|\vec{e}\varphi\|_{C^{1}(\Sigma_{t})}$ $\leq C,$
(25b) $t\|k\|_{C^{1}(\Sigma_{t})}+t\|e_{0}\varphi\|_{C^{1}(\Sigma_{t})}$ $\leq C,$
(25c) $\|N-1\|_{C^{2}(\Sigma_{t})}$ $\leq Ct^{\varepsilon}$
for $t\leq t_{0}$. In addition to this, assume, for $t\leq t_{0}$,
(26) $t^{2}\|t^{-2}-k_{IJ}k_{IJ}-e_{0}(\varphi)e_{0}(\varphi)\|_{C^{1}(\Sigma_{t})}\leq Ct^{2\varepsilon}.$
Then the solution is said to satisfy the basic supremum assumptions.
Remark 18.
In the definition we use the notation introduced in Subsection A.4.
Given these assumptions, the following conclusions hold.
Theorem 19.
Consider a solution to (11)–(19) as described in Subsection 1.4. Assume, in addition, that $V$ is a $\sigma_{V}$-admissible potential for some $\sigma_{V}\in(0,1)$ and that the basic supremum assumptions, see Definition 17, are satisfied with $3\varepsilon\leq\sigma_{V}$. Then there is, for every $\ell\in\mathbb{N}_{0}$, a constant $C_{\ell}$ such that
(27a) $t^{1-2\varepsilon}\|e\|_{C^{\ell}(\Sigma_{t})}+t^{1-2\varepsilon}\|\omega\|_{C^{\ell}(\Sigma_{t})}+t^{1-\varepsilon}\|\gamma\|_{C^{\ell}(\Sigma_{t})}+t^{1-\varepsilon}\|\vec{e}\varphi\|_{C^{\ell}(\Sigma_{t})}$ $\leq C_{\ell},$
(27b) $t\|k\|_{C^{\ell}(\Sigma_{t})}+t\|e_{0}\varphi\|_{C^{\ell}(\Sigma_{t})}$ $\leq C_{\ell},$
(27c) $t^{1-3\varepsilon}\|\vec{e}N\|_{C^{\ell}(\Sigma_{t})}+t^{-\varepsilon/2}\|N-1\|_{C^{\ell}(\Sigma_{t})}$ $\leq C_{\ell}$
hold for all $\ell\in\mathbb{N}_{0}$ and all $t\leq t_{0}$. Moreover, there are $\mathring{\Phi},\mathring{\Psi}\in C^{\infty}(\Sigma)$ and, for every $I,J\in\{1,\dots,n\}$, $\mathring{\mathfrak{K}}_{IJ}\in C^{\infty}(\Sigma)$ such that
(28a) $\|tk_{IJ}(\cdot,t)-\mathring{\mathfrak{K}}_{IJ}\|_{C^{\ell}(\Sigma)}$ $\leq C_{\ell}t^{\varepsilon/4},$
(28b) $\|t(e_{0}\varphi)(\cdot,t)-\mathring{\Psi}\|_{C^{\ell}(\Sigma)}+\|\varphi(\cdot,t)-\mathring{\Psi}\ln t-\mathring{\Phi}\|_{C^{\ell}(\Sigma)}$ $\leq C_{\ell}t^{\varepsilon/4}$
hold for all $\ell\in\mathbb{N}_{0}$, all $I,J\in\{1,\dots,n\}$ and all $t\leq t_{0}$.
Remark 20.
The eigenvalues of the expansion normalised Weingarten map are not assumed to be distinct in this theorem.
Proof.
The proof is to be found at the end of Subsection 2.10. ∎
1.6. Deriving asymptotics, given higher order derivative bounds
Our next goal is to derive asymptotics, given higher order derivative bounds. We do so for crushing foliations that are not necessarily CMC, since the results are applicable in a more general setting we now describe. To begin with, we consider a spacetime $(M,g)$ with a smooth foliation by spacelike hypersurfaces, lapse $N>0$ and vanishing shift vector field, so that $M=\Sigma\times\mathcal{I}$ and
(29) $g=-N^{2}{\mathrm{d}}t\otimes{\mathrm{d}}t+h,$
where $h$ denotes the family of induced (Riemannian) metrics on the $\Sigma_{t}:=\Sigma\times\{t\}$; see Subsection A.3 for our conventions concerning the geometry. In this setting, and with the conventions introduced in Subsection A.3, the Einstein equations read as follows.
Proposition 21.
In the setting described above and with the conventions introduced in Subsection A.3, let $\varphi\in C^{\infty}(M)$. Then $(M,g,\varphi)$ solves the Einstein–nonlinear scalar field equations with potential $V\in C^{\infty}(\mathbb{R})$ if and only if the following system of equations holds: the evolution equations for $K$ and $\varphi$,
(30a) $\mathcal{L}_{e_{0}}K+\mathrm{Ric}_{h}^{\sharp}+\theta K-N^{-1}(\overline{\nabla}^{2}N)^{\sharp}=$ ${\mathrm{d}}\varphi\otimes\mathrm{grad}_{h}\varphi+\tfrac{2}{n-1}(V\circ\varphi)\operatorname{Id},$
(30b) $-e_{0}e_{0}\varphi+\Delta_{h}\varphi-\theta e_{0}\varphi+\langle{\mathrm{d}}\varphi,{\mathrm{d}}(\ln N)\rangle_{h}=$ $V^{\prime}\circ\varphi,$
where $\operatorname{Id}$ denotes the family of identity $(1,1)$-tensors on the $\Sigma_{t}$; and the constraint equations:
(31a) $\mathrm{Scal}_{h}+\theta^{2}-\mathrm{tr}K^{2}$ $=e_{0}(\varphi)^{2}+|{\mathrm{d}}\varphi|_{h}^{2}+2V\circ\varphi,$
(31b) $\mathrm{div}_{h}K-{\mathrm{d}}\theta$ $=e_{0}(\varphi){\mathrm{d}}\varphi.$
Remark 22.
Equations (31a) and (31b) are of course the Hamiltonian and momentum constraint equations on the leaves of the foliation; see (3).
Proof.
To begin with, (30a) is, e.g., an immediate consequence of [43, (269), p. 65], keeping in mind that
$\rho=T(e_{0},e_{0}),\ \ \ \bar{p}=n^{-1}h^{ij}T_{ij},$
$\mathcal{P}$ is defined by [43, (17), p. 8] and the fact that $T$ is given by (2). Next, (30b) follows from a straightforward computation and (31) follow from [39, Chapter 13].
Turning to the other direction, (1b) is an immediate consequence of (30b). Due to (31) the normal–normal and the normal–tangential components of Einstein’s equations hold. Next, combining [43, (257), p. 63] and [43, (269), p. 65] with (30a), it is clear that the spatial–spatial components of the spacetime Ricci tensor equal what they would if Einstein’s equations would hold. Combining this information with the fact that the normal–normal and the normal–tangential components of Einstein’s equations hold, it can be deduced that (1a) holds. ∎
Definition 23.
Let $3\leq n\in\mathbb{N}$, $k_{0}\in\mathbb{N}$, $r>0$, $\delta\in(0,1)$ and fix a $\delta$-admissible potential $V\in C^{\infty}(\mathbb{R})$. Let $(M,g,\varphi)$ be a solution to the Einstein–nonlinear scalar field equations with potential $V$ such that $M=\Sigma\times(0,T)$, where $0<T\in\mathbb{R}$; $\Sigma$ is a closed $n$-dimensional manifold; $\Sigma$ has a global frame $\{E_{i}\}_{i=1}^{n}$ with dual frame $\{\eta^{i}\}_{i=1}^{n}$; and the metric $g$ takes the form (29), where $N>0$ and $h$ denotes a family of Riemannian metrics on $\Sigma_{t}:=\Sigma\times\{t\}$. Let $\theta$ denote the mean curvature of $\Sigma_{t}$. Let $\{e_{I}\}_{I=1}^{n}$ be a family of Fermi-Walker transported orthonormal frames on the $\Sigma_{t}$ with respect to $h$, so that
(32) $\nabla_{e_{0}}e_{I}=e_{I}(\ln N)e_{0},$
where $e_{0}$ is the future directed unit normal to the $\Sigma_{t}$ and $\nabla$ denotes the Levi-Civita connection of $(M,g)$. Assume that $\theta>0$ and let $\mathcal{H}$, $\mathcal{K}$, $\Phi$ and $\Psi$ be the expansion normalised initial data induced on $\Sigma_{t}$. Let $p_{I}$ be the eigenvalues of $\mathcal{K}$ and assume them to be distinct everywhere. Assume that there are everywhere distinct functions $\mathring{p}_{1},\ldots,\mathring{p}_{n}\in C^{k_{0}}(\Sigma)$, and functions $\mathring{\Phi},\mathring{\Psi}\in C^{k_{0}}(\Sigma)$ satisfying
(33) ${\sum}_{I}\mathring{p}_{I}=\textstyle{\sum}_{I}\mathring{p}_{I}^{2}+\mathring{\Psi}^{2}=1,$
and
(34) $\mathring{p}_{I}+\mathring{p}_{J}-\mathring{p}_{K}<1-2\delta,\qquad I\neq J.$
Assume that there is a constant $\mathbf{c}_{0}$ such that
(35) $\|p_{I}(t)-\mathring{p}_{I}\|_{C^{k_{0}}(\Sigma)}+\|\Phi(t)-\mathring{\Phi}\|_{C^{k_{0}}(\Sigma)}+\|\Psi(t)-\mathring{\Psi}\|_{C^{k_{0}}(\Sigma)}\leq\mathbf{c}_{0}t^{\delta}$
for all $t<T$. Let $k_{IJ}:=k(e_{I},e_{J})$; $\{\omega^{I}\}_{I=1}^{n}$ denote the frame dual to $\{e_{I}\}_{I=1}^{n}$; and $\gamma_{IJK}:=\omega^{K}([e_{I},e_{J}])$ denote the structure coefficients of the frame $\{e_{I}\}_{I=1}^{n}$. Let $e_{I}=e_{I}^{i}E_{i}$ and $\omega^{I}=\omega_{i}^{I}\eta^{i}$. Define $\Theta:=N\theta$. Assume that the following estimates hold:
(36a) $t\|k\|_{C^{k_{0}}(\Sigma)}+t\|e_{0}\varphi\|_{C^{k_{0}}(\Sigma)}$ $\leq\mathbf{c}_{0},$
(36b) $t^{1-3\delta}\|e\|_{C^{k_{0}}(\Sigma)}+t^{1-3\delta}\|\omega\|_{C^{k_{0}}(\Sigma)}+t^{-\delta}\|t\Theta-1\|_{C^{k_{0}}(\Sigma)}+t^{1-3\delta}\|\gamma\|_{C^{k_{0}}(\Sigma)}$ $\leq\mathbf{c}_{0},$
(36c) $\|N\|_{C^{k_{0}}(\Sigma)}+\|N^{-1}\|_{C^{0}(\Sigma)}+t^{1-\delta}\|\vec{e}(\ln N)\|_{C^{k_{0}}(\Sigma)}+t^{1-\delta}\|\vec{e}\varphi\|_{C^{k_{0}}(\Sigma)}$ $\leq\mathbf{c}_{0},$
(36d) $\|t\Theta-1\|_{C^{0}(\Sigma)}$ $\leq\tfrac{1}{2}$
for $t<T$, where we use the notation introduced in Subsection A.4. Finally, assume that $|p_{I}-p_{J}|\geq r$ for $I\neq J$. Then $(M,g,\varphi)$ is said to be a solution to the Einstein–nonlinear scalar field equations satisfying the $(\delta,n,k_{0},r)$-assumptions.
Theorem 24.
[…]
$\|\mathcal{H}(t)-\mathring{\mathcal{H}}\|_{C^{\ell}(\Sigma)}+\|\mathcal{K}(t)-\mathring{\mathcal{K}}\|_{C^{\ell}(\Sigma)}\leq Ct^{\delta}$
for some constant $C$ and all $t\leq t_{0}$. Moreover, $\mathring{p}_{I}$ are the eigenvalues of $\mathring{\mathcal{K}}$, and $(\Sigma,\mathring{\mathcal{H}},\mathring{\mathcal{K}},\mathring{\Phi},\mathring{\Psi})$ are $C^{\ell}$ robust nondegenerate quiescent initial data on the singularity for the Einstein–nonlinear scalar field equations.
Remark 25.
The same statement holds if we assume $\Sigma$ to be a smooth compact manifold with boundary. In particular, in case $\Sigma$ is non–compact, localising the assumptions to compact balls in $\Sigma$, we obtain correspondingly localised conclusions.
Remark 26.
In the case of the FRS equations, combining Theorem 19 with Theorem 24 yields stronger conclusions (smooth initial data on the singularity) under weaker assumptions. The reader is encouraged to formulate the corresponding result.
Proof.
The proof is to be found at the end of Section 3. ∎
1.7. Outline of the arguments
The main result of this article is Theorem 11. However, Theorem 11 is an immediate consequence of Theorems 19 and 24. Moreover, the latter theorems have an independent interest in their own right. In particular, it is to be expected that Theorem 24 can be used to obtain data on the singularity in other settings. For these reasons, we here focus on a description of the arguments required to prove Theorems 19 and 24.
Higher order derivative estimates. In [33], the authors derive partial information concerning the asympotics of solutions in the $C^{k}$-norm for fixed finite $k$. However, as $k$ tends to infinity, the number of derivatives you need to control initially tends to infinity. This is unfortunate. In the present article, we therefore begin by deriving partial information concerning the asymptotics for all derivatives, given assumptions concerning the initial data in a fixed finite degree of regularity. This is the purpose of Theorem 19. The basic idea is to prove energy estimates inductively on the order $\ell$ of the energy $\mathcal{E}_{\mathrm{tot},\ell}$ introduced in (43). For the first order energy, we have estimates by assumption. For a higher order, say $\ell$, the goal is to prove an estimate of the form (108). In this estimate, we effectively control the last term on the right hand side due to the induction hypothesis and we control the second term on the right hand side because the factor in front of the energy is integrable. For these reasons, we can effectively ignore the last two terms on the right hand side of (108), with a slight loss. This means, inductively, that the energy only grows at the rate $t^{-9}$ in the direction of the singularity; see Corollary 36. Moreover, the rate is independent of the number of derivatives and of the dimension. Combining this estimate with interpolation yields good enough estimates for most of the quantities of interest. However, the estimates for $k_{IJ}$ and $e_{0}\varphi$ are not immediately optimal. Nevertheless, combining our improved estimates with the equations, optimal estimates for these quantities can also be derived. Moreover, we obtain the conclusions of Theorem 19.
To conclude, the main point is to derive an estimate of the form (108). We begin by deriving some basic consequences of the assumptions of Theorem 19 in Subsections 2.1 and 2.3. Moreover, in Subsection 2.2, we introduce notation for our basic energies. It is of particular importance to note that we include additional time dependent weights (that tend to infinity in the direction of the singularity) in the energies associated with the frame and co-frame; see (41a). In Subsection 2.4 we derive an algorithm for estimating expressions that occur in the energy estimates. The main point here is to develop a calculus that allows us to divide the expressions that appear in the energy estimates into terms that can be estimated by the last two terms on the right hand side of (108) and terms that contribute to the first term on the right hand side. In the development of the algorithm, it is crucial that we have the extra time dependent weights in the energies for the frame and co-frame. With the algorithm at hand, we derive elliptic estimates for the lapse function in Subsection 2.5. The identity (57) may look odd at first sight. However, it turns out to be important since the second term on the left hand side appears in the energy estimates. We also need (58) and (59). It is important to note here that the coefficients in front of the squared Sobolev norms on the right hand sides of (58) and (59) are numerical. Deriving energy estimates for the frame and co-frame is straightforward; see Subsection 2.6. Estimating the second fundamental form and the structure coefficients is more complicated. However, integrating by parts and appealing to the algorithm yields (74). In this estimate, it is important to note that the terms appearing in the integrands have the form $aE_{\mathbf{I}}(f)E_{\mathbf{I}}(g)$. Moreover, the $f$ and $g$ appearing are basic variables in the FRS equations. In addition, $a$ is either a numerical multiple of $t^{-1}$ or a numerical multiple of $k_{IJ}$ or $e_{0}\varphi$. The reason estimates of this form are useful when deriving (108) is that, due to (26),
$t^{2}k_{IJ}k_{IJ}+t^{2}(e_{0}\varphi)^{2}\leq 1+Ct^{2\varepsilon}.$
There is only one problem with (74): expressions of the form $E_{\mathbf{I}}(\gamma_{IJJ})$ appear. Such terms could potentially give rise to dimensionally dependent constants in the first term on the right hand side of (108). We avoid this possibility by estimating the energy for $a_{I}$, introduced in (42), separately in Subsection 2.9. Note that the energy estimate again has a similar structure to that described above. All that remains is to estimate the energy associated with the scalar field. We do so in Subsection 2.8. Again, the relevant estimate has the form described above. Combining all the above observations, we, in the end, arrive at the estimate (108) and thereby at the desired conclusion.
Asymptotics. In the context of quiescent big bang formation, one would like to show that solutions induce data on the singularity as in [45] or [23]. In particular, if the solution is expected to exhibit stable big bang formation, one would like to show that it induces data on the singularity as in Definition 6. In [33], the authors obtain some of these data. Specifically, if $(\Sigma,\mathcal{H},\mathcal{K},\Phi,\Psi)$ denotes the expansion normalized initial data along the foliation, the authors prove the convergence of the eigenvalues of $\mathcal{K}$, of $\Phi$ and of $\Psi$, in $C^{k_{0}}$, for some fixed $k_{0}\in\mathbb{N}$, to nontrivial limits satisfying the expected constraints, (33) and (34). However, they fail to obtain convergence of $\mathcal{H}$ and $\mathcal{K}$, due to limited information on the asymptotic behavior of the Fermi-Walker frame $\{e_{I}\}_{I=1}^{n}$. The authors obtain estimates of the form
$\|e_{I}\|_{C^{k_{0}}(\Sigma)}\leq Ct^{-1+3\sigma}$
for some $\sigma>0$, along with similar estimates for the corresponding dual frame and structure coefficients, which are not sharp. That this is enough information to prove big bang formation is a remarkable property of the Fermi-Walker frame. One does not need to have full knowledge about the asymptotics to prove curvature blow-up.
Following the results of [33], it makes sense to set as our starting point some $C^{k_{0}}$ bounds on a Fermi-Walker frame and on the components of geometric quantities in terms of this frame, along with $C^{k_{0}}$ convergence of the eigenvalues of $\mathcal{K}$, of $\Phi$ and of $\Psi$. These are precisely the estimates (36) and (35). Note that these estimates correspond to the conclusions obtained in [33] with two exceptions. First, since we do not assume the CMC condition, we need a way to control the mean curvature. This is done through the quantity $\Theta=N\theta$, which is assumed to be asymptotic to $t^{-1}$ as $t\downarrow 0$. Second, in [33], the authors obtain convergence of the lapse, $N\to 1$ in $C^{k_{0}}$ as $t\downarrow 0$. By contrast, here we only assume that $N$ is bounded in $C^{k_{0}}$, and that $N^{-1}$ is bounded in $C^{0}$. However, we point out that if the CMC condition is assumed, then the assumption that $t\Theta-1\to 0$ becomes $N-1\to 0$. We begin by deriving some basic consequences of the assumptions in Subsection 3.1.
[…]
Proof of the main result. Once we have Theorems 19 and 24 at our disposal, we can combine them with the arguments in [33] in order to deduce that Theorem 11 holds. The details are to be found in Section 4.
1.8. Acknowledgements
This research was funded by the Swedish Research Council (Vetenskapsrådet), dnr. 2022-03053. The first author was supported by foundations managed by The Royal Swedish Academy of Sciences.
2. Higher order energy estimates for solutions to the FRS equations
The purpose of the present section is to prove Theorem 19. In particular, we, in the present section, consider solutions to (11)–(19) as described in Subsection 1.4. Moreover, we, unless otherwise stated, assume that $V$ is a $\sigma_{V}$-admissible potential and that the basic supremum assumptions, see Definition 17, are satisfied with $3\varepsilon\leq\sigma_{V}$.
2.1. Elementary consequences of the basic supremum assumptions
Due to (26), we know that $t|e_{0}\varphi|\leq 1+Ct^{2\varepsilon}$ for $t\leq t_{0}$. Combining this estimate with (25c) yields
$t\|\partial_{t}\varphi\|_{C^{0}(\Sigma_{t})}\leq 1+Ct^{\varepsilon}$
for $t\leq t_{0}$. This means that
(37) $\|\varphi\|_{C^{0}(\Sigma_{t})}\leq-\ln t+C$
for all $t\leq t_{0}$. Similarly, combining (25b) and (25c), we conclude that
(38) $\|\varphi\|_{C^{1}(\Sigma_{t})}\leq C\langle\ln t\rangle$
for all $t\leq t_{0}$. Next, since
$\mathrm{div}_{h_{\mathrm{ref}}}e_{I}=\mathrm{div}_{h_{\mathrm{ref}}}(e_{I}^{i}E_{i})=E_{i}(e_{I}^{i})+e_{I}^{i}\mathrm{div}_{h_{\mathrm{ref}}}E_{i},$
it follows from (25a) that
(39) $t^{1-3\varepsilon}\|\mathrm{div}_{h_{\mathrm{ref}}}e_{I}\|_{C^{0}(\Sigma_{t})}\leq C$
for all $t\leq t_{0}$. Finally, note that combining (25a) and (25c) yields
(40) $t^{1-4\varepsilon}\|\vec{e}N\|_{C^{1}(\Sigma_{t})}\leq C$
for all $t\leq t_{0}$.
2.2. Energies and basic estimates
The energies we estimate in the end are, given $\ell\in\mathbb{N}_{0}$, defined as follows
(41a) $E_{(e,\omega),\ell}(t)$ $:=t^{-2\varepsilon}\|e\|_{H^{\ell}(\Sigma_{t})}^{2}+t^{-2\varepsilon}\|\omega\|_{H^{\ell}(\Sigma_{t})}^{2},$
(41b) $E_{(\gamma,k),\ell}(t)$ $:=\tfrac{1}{2}\|\gamma\|_{H^{\ell}(\Sigma_{t})}^{2}+\|k\|_{H^{\ell}(\Sigma_{t})}^{2},$
(41c) $E_{(\varphi),\ell}(t)$ $:=\|e_{0}\varphi\|_{H^{\ell}(\Sigma)}^{2}+\|\vec{e}\varphi\|_{H^{\ell}(\Sigma_{t})}^{2}+t^{-2+2\varepsilon}\|\varphi\|_{H^{\ell}(\Sigma_{t})}^{2},$
(41d) $E_{(N),\ell}(t)$ $:=\|\vec{e}N\|_{H^{\ell}(\Sigma_{t})}^{2}+t^{-2}\|N-1\|_{H^{\ell}(\Sigma_{t})}^{2},$
(41e) $E_{\mathrm{tot},\ell}(t)$ $:=E_{(e,\omega),\ell}(t)+E_{(\gamma,k),\ell}(t)+E_{(\varphi),\ell}(t),$
(41f) $\hat{E}_{\mathrm{tot},\ell}(t)$ $:=E_{\mathrm{tot},\ell}+E_{(N),\ell}(t).$
Here $\varepsilon>0$ is the constant whose existence is assumed in Definition 17. Using notation analogous to the one introduced in (137), we, next, define an energy for $a$ introduced in (23):
(42) $E_{(a),\ell}(t):=\|a\|_{H^{\ell}(\Sigma_{t})}^{2}.$
It turns out to be convenient to include this expression in the total energy, and in the final step of the energy estimates, we therefore focus on
(43) $\mathcal{E}_{\mathrm{tot},\ell}:=E_{\mathrm{tot},\ell}+E_{(a),\ell}.$
In Subsection A.5 we record some of the basic estimates and identities we use when deriving energy estimates.
2.3. Estimating the potential
Next, we record some basic estimates concerning the potential.
Lemma 27.
Consider a solution to (11)–(19) as described in Subsection 1.4. Assume, in addition, that $V$ is a $\sigma_{V}$-admissible potential and that the basic supremum assumptions, see Definition 17, are satisfied with $3\varepsilon\leq\sigma_{V}$. Then, for any $\ell\in\mathbb{N}_{0}$, there is a constant $C_{\ell}$, depending only on $(\Sigma,h_{\mathrm{ref}})$, $c_{\ell}$, $\varepsilon$ and $\ell$, such that the following holds for all $t\leq t_{0}$:
(44) $t^{2}\|V\circ\varphi\|_{H^{\ell}(\Sigma_{t})}+t^{2}\|V^{\prime}\circ\varphi\|_{H^{\ell}(\Sigma_{t})}\leq C_{\ell}t^{5\varepsilon}(1+\|\varphi\|_{H^{\ell}(\Sigma_{t})}).$
Moreover,
(45) $t^{2}\|E_{i}(V\circ\varphi)\|_{H^{\ell-1}(\Sigma_{t})}+t^{2}\|E_{i}(V^{\prime}\circ\varphi)\|_{H^{\ell-1}(\Sigma_{t})}\leq C_{\ell}t^{5\varepsilon}\|\varphi\|_{H^{\ell}(\Sigma_{t})}.$
Finally,
(46) $t^{2}\|V\circ\varphi\|_{C^{1}(\Sigma_{t})}+t^{2}\|V^{\prime}\circ\varphi\|_{C^{1}(\Sigma_{t})}\leq C_{\ell}t^{5\varepsilon}.$
Remark 28.
In particular, using the terminology introduced in (41c),
$t^{2}\|E_{i}(V\circ\varphi)\|_{H^{\ell-1}(\Sigma_{t})}+t^{2}\|E_{i}(V^{\prime}\circ\varphi)\|_{H^{\ell-1}(\Sigma_{t})}\leq C_{\ell}t^{1+4\varepsilon}E_{(\varphi),\ell}^{1/2}.$
Proof.
Combining (37) with (7) yields the conclusion that
(47) ${\sum}_{m\leq k}|V^{(m)}\circ\varphi|\leq c_{k}e^{2C}t^{-2+2\sigma_{V}}\leq C_{k}t^{-2+6\varepsilon}$
for any $k\in\mathbb{N}_{0}$, where we used the fact that $\sigma_{V}\geq 3\varepsilon$. Next, note that $E_{\mathbf{I}}V^{(i)}\circ\varphi$ consists of a linear combination of terms of the form
$V^{(i+m)}\circ\varphi E_{\mathbf{I}_{1}}\varphi\cdots E_{\mathbf{I}_{m}}\varphi.$
The estimates (44) and (45) follow by Gagliardo-Nirenberg estimates, i.e. Lemma 57, (37) and the fact that $\langle\ln t\rangle^{m-1}t^{\varepsilon}$ is bounded for $t\leq 1$. Finally, combining (38) with (47) yields (46). ∎
2.4. Algorithm
Consider a solution to (11)–(19) as described in Subsection 1.4. Assume, in addition, that $V$ is a $\sigma_{V}$-admissible potential and that the basic supremum assumptions, see Definition 17, are satisfied with $3\varepsilon\leq\sigma_{V}$. Next, we formulate an algorithm for estimating expressions that occur frequently.
General terms. Consider a term $T$ consisting of the following factors: $l_{1}$ factors of the form $N$; $l_{2}$ factors of the form N1N-1; $l_{3,a}$ factors of the form $e_{I}^{i}$ or $\omega_{i}^{I}$; $l_{3,b}$ factors of the form $t^{-\varepsilon}e_{I}^{i}$ or $t^{-\varepsilon}\omega_{i}^{I}$; $l_{4}$ factors of the form $e_{I}N$; $l_{5}$ factors of the form $\gamma_{IJK}$; $l_{6}$ factors of the form $k_{IJ}$; $l_{7}$ factors of the form $e_{0}\varphi$; $l_{8}$ factors of the form $e_{I}\varphi$; and $l_{9}$ factors of the form $V\circ\varphi$ or $V^{\prime}\circ\varphi$. In what follows, we refer to such a term as a general term. Let
$m_{\mathrm{int}}$ $:=-l_{3,a}-l_{3,b}-l_{4}-l_{5}-l_{6}-l_{7}-l_{8}-2l_{9},$
$m_{\varepsilon}$ $:=l_{2}+3l_{3,a}+2l_{3,b}+4l_{4}+2l_{5}+2l_{8}+5l_{9},$
$\ell_{\varepsilon}$ $:=l_{2}+l_{3,a}+l_{3,b}+l_{4}+l_{5}+l_{8}+l_{9}.$
By (25), (40), (45), (46) and Gagliardo-Nirenberg estimates, $E_{\mathbf{I}}T$, where $0<|\mathbf{I}|\leq\ell$, can be estimated in $L^{2}$ by
$\begin{split}&C_{\ell}t^{m_{\mathrm{int}}+\varepsilon m_{\varepsilon}}[\iota_{1}tE_{(N),\ell}^{1/2}+\iota_{2}t^{1-\varepsilon}E_{(N),\ell}^{1/2}+\iota_{3,a}t^{1-3\varepsilon}t^{\varepsilon}E_{(e,\omega),\ell}^{1/2}+\iota_{3,b}t^{1-2\varepsilon}E_{(e,\omega),\ell}^{1/2}+\iota_{4}t^{1-4\varepsilon}E_{(N),\ell}^{1/2}]\\ &+C_{\ell}t^{m_{\mathrm{int}}+\varepsilon m_{\varepsilon}}[\iota_{5}t^{1-2\varepsilon}E_{(\gamma,k),\ell}^{1/2}+\iota_{6}tE_{(\gamma,k),\ell}^{1/2}+\iota_{7}tE_{(\varphi),\ell}^{1/2}+\iota_{8}t^{1-2\varepsilon}E_{(\varphi),\ell}^{1/2}+\iota_{9}t^{2-5\varepsilon}t^{-1+4\varepsilon}E_{(\varphi),\ell}^{1/2}],\end{split}$
where $\iota_{i}=1$ if $l_{i}>0$ and $\iota_{i}=0$ if $l_{i}=0$. The quantities $\iota_{3,a}$ and $\iota_{3,b}$ are defined similarly.
The good terms. If $T$ is a general term with $\ell_{\varepsilon}\geq 2$ or $l_{3,a}+l_{9}\geq 1$, then $T$ is called a good term. Then
(48) $\|E_{\mathbf{I}}T\|_{L^{2}(\Sigma_{t})}\leq C_{\ell}t^{m_{\mathrm{int}}+1+\varepsilon}\hat{E}_{\mathrm{tot},\ell}^{1/2}.$
Moreover, if $l_{1}+l_{2}+l_{4}=0$, then $\hat{E}_{\mathrm{tot},\ell}$ can be replaced by $E_{\mathrm{tot},\ell}$ on the right hand side.
The bad terms. If $T$ is a general term with $\ell_{\varepsilon}=1$ and $l_{3,a}+l_{9}=0$, then $T$ is called a bad term. In this case, we can assume $T$ to consist of $l_{1}$ factors of the form $N$; one factor, say $f$, which is one of $t^{-1}(N-1)$, $t^{-\varepsilon}e_{I}^{i}$, $t^{-\varepsilon}\omega^{I}_{i}$, $e_{I}N$, $\gamma_{IJK}$ or $e_{I}\varphi$; $l_{6}$ factors of the form $k_{IJ}$; and $l_{7}$ factors of the form $e_{0}\varphi$. Again, we apply $E_{\mathbf{I}}$ to such a term. If no derivative hits $f$, then, appealing to Lemma 57, (25) and (40), the corresponding expression can be estimated by
(49) $C_{\ell}t^{-l_{6}-l_{7}+1}t^{-1+\varepsilon}\hat{E}_{\mathrm{tot},\ell}^{1/2}.$
Moreover, if $l_{1}=0$, then $\hat{E}_{\mathrm{tot},\ell}$ can be replaced by $E_{\mathrm{tot},\ell}$ on the right hand side. If at least one derivative hits $f$ and there is one other factor that is hit by at least one derivative, then one can retain one derivative on $f$ and one derivative on one of the other factors. This leaves $\ell-2$ derivatives. Again, due to Lemma 57, (25) and (40),
$C_{\ell}t^{-l_{6}-l_{7}+1}t^{-1}\hat{E}_{\mathrm{tot},\ell-1}^{1/2}$
for the corresponding expressions. Moreover, $\hat{E}_{\mathrm{tot},\ell-1}$ can be replaced by $E_{\mathrm{tot},\ell-1}$ if $l_{1}=0$ and $f$ is neither $t^{-1}(N-1)$ nor $e_{I}N$. What remains is when $E_{\mathbf{I}}$ only hits $f$. Since N1N-1 can be bounded by $Ct^{\varepsilon}$ in $C^{0}$, we can replace all the occurrences of $N$ in the corresponding expression with $1$ at the price of an error of the form (49). What we are left with are $E_{\mathbf{I}}f$ times the original $l_{6}$ factors of the form $k_{IJ}$ and the original $l_{7}$ factors of the form $e_{0}\varphi$. In other words,
(50) $E_{\mathbf{I}}T=E_{\mathbf{I}}f\cdot k_{I_{1}J_{1}}\cdots k_{I_{l_{6}}J_{l_{6}}}(e_{0}\varphi)^{l_{7}}+R_{B},$
where
(51) $\|R_{B}\|_{L^{2}(\Sigma_{t})}\leq C_{\ell}t^{-l_{6}-l_{7}+1}t^{-1+\varepsilon}\hat{E}_{\mathrm{tot},\ell}^{1/2}+C_{\ell}t^{-l_{6}-l_{7}+1}t^{-1}\hat{E}_{\mathrm{tot},\ell-1}^{1/2}.$
Again, the hats can be removed on the right hand side if $l_{1}=0$ and $f$ is neither $t^{-1}(N-1)$ nor $e_{I}N$. For future reference, it is also of interest to note that
$\|E_{\mathbf{I}}T\|_{L^{2}(\Sigma_{t})}\leq C_{\ell}t^{-l_{6}-l_{7}+1}t^{-1}\hat{E}_{\mathrm{tot},\ell}^{1/2}.$
Again, if $l_{1}=0$ and $f$ is neither $t^{-1}(N-1)$ nor $e_{I}N$, the hat can be removed on the right hand side.
The leading order terms. If $T$ is a general term with $\ell_{\varepsilon}=0$, then $T$ is called a leading order term. Then, by arguments similar to the above,
(52) $\begin{split}&E_{\mathbf{I}}[N^{l_{1}}k_{I_{1}J_{1}}\cdots k_{I_{l_{6}}J_{l_{6}}}(e_{0}\varphi)^{l_{7}}]\\ ={}&l_{1}E_{\mathbf{I}}(N)k_{I_{1}J_{1}}\cdots k_{I_{l_{6}}J_{l_{6}}}(e_{0}\varphi)^{l_{7}}+E_{\mathbf{I}}(k_{I_{1}J_{1}})\cdots k_{I_{l_{6}}J_{l_{6}}}(e_{0}\varphi)^{l_{7}}\\ &\negmedspace+\dots+k_{I_{1}J_{1}}\cdots E_{\mathbf{I}}(k_{I_{l_{6}}J_{l_{6}}})(e_{0}\varphi)^{l_{7}}+l_{7}k_{I_{1}J_{1}}\cdots k_{I_{l_{6}}J_{l_{6}}}E_{\mathbf{I}}(e_{0}\varphi)\cdot(e_{0}\varphi)^{l_{7}-1}+R_{B},\end{split}$
where the first term is absent if $l_{1}=0$; the terms involving derivatives of the $k_{IJ}$ are absent if $l_{6}=0$; the second last term is absent if $l_{7}=0$; and $R_{B}$ can be estimated in $L^{2}$ by
(53) $\|R_{B}\|_{L^{2}(\Sigma_{t})}\leq C_{\ell}t^{-l_{6}-l_{7}+1}t^{\varepsilon}\hat{E}_{\mathrm{tot},\ell}^{1/2}+C_{\ell}t^{-l_{6}-l_{7}+1}\hat{E}_{\mathrm{tot},\ell-1}^{1/2}.$
Moreover, if $l_{1}=0$, the hats can be removed from the right hand side. For future reference, it is also of interest to keep in mind that
(54) $\|E_{\mathbf{I}}T\|_{L^{2}(\Sigma_{t})}\leq C_{\ell}t^{-l_{6}-l_{7}+1}\hat{E}_{\mathrm{tot},\ell}^{1/2}.$
Again, if $l_{1}=0$, the hat can be removed from the right hand side.
Commutators. Next, consider $E_{\mathbf{I}},e_{I}]\psi$. It can be written as $e_{I}^{i}[E_{\mathbf{I}},E_{i}]\psi$, plus a linear combination of terms of the form
(55) $E_{\mathbf{J}}(e_{I}^{i})E_{\mathbf{K}}\psi,$
where $|\mathbf{J}|+|\mathbf{K}|=|\mathbf{I}|+1$, $|\mathbf{J}|\geq 1$ and $|\mathbf{K}|\geq 1$. Due to (25a),
$\|e_{I}^{i}[E_{\mathbf{I}},E_{i}]\psi\|_{L^{2}(\Sigma_{t})}\leq C_{\ell}t^{-1+3\varepsilon}\|\psi\|_{H^{\ell}(\Sigma_{t})}$
for some constant $C_{\ell}$, where $|\mathbf{I}|\leq\ell$. Next, note that (55) can be written
$E_{\mathbf{J}_{1}}E_{j_{1}}(e_{I}^{i})E_{\mathbf{K}_{1}}E_{k_{1}}\psi,$
where $|\mathbf{J}_{1}|+|\mathbf{K}_{1}|\leq|\mathbf{I}|-1$. By Gagliardo-Nirenberg estimates, we conclude that
$\|E_{\mathbf{J}}(e_{I}^{i})E_{\mathbf{K}}\psi\|_{L^{2}(\Sigma_{t})}\leq C_{\ell}\|e\|_{C^{1}(\Sigma_{t})}\|\psi\|_{H^{\ell}(\Sigma_{t})}+C_{\ell}\|\psi\|_{C^{1}(\Sigma_{t})}\|e\|_{H^{\ell}(\Sigma_{t})}.$
Note that the first term on the right hand side can be estimated by appealing to (25a). This yields, assuming $|\mathbf{I}|\leq\ell$ and keeping (41a) in mind,
(56) $\|[E_{\mathbf{I}},e_{I}]\psi\|_{L^{2}(\Sigma_{t})}\leq C_{\ell}t^{-1+3\varepsilon}\|\psi\|_{H^{\ell}(\Sigma_{t})}+C_{\ell}t^{\varepsilon}\|\psi\|_{C^{1}(\Sigma_{t})}E_{(e,\omega),\ell}^{1/2}.$
2.5. Estimating the lapse function
Concerning the lapse function, the following estimates hold.
Lemma 29.
Consider a solution to (11)–(19) as described in Subsection 1.4. Assume, in addition, that $V$ is a $\sigma_{V}$-admissible potential and that the basic supremum assumptions, see Definition 17, are satisfied with $3\varepsilon\leq\sigma_{V}$. Then, for each $\ell\in\mathbb{N}$, there is a $t_{\ell}>0$ such that for $t\leq t_{\ell}$,
(57) $\begin{split}&E_{(N),\ell}+2\textstyle{\sum}_{0<|\mathbf{I}|\leq\ell}\textstyle{\int}_{\Sigma}\big(k_{IJ}E_{\mathbf{I}}k_{IJ}+e_{0}(\varphi)E_{\mathbf{I}}e_{0}(\varphi)\big)E_{\mathbf{I}}(N-1){\mathrm{d}}\mu_{h_{\mathrm{ref}}}=\mathcal{R},\end{split}$
where
$|\mathcal{R}|\leq C_{\ell}t^{\varepsilon}E_{\mathrm{tot},\ell}+C_{\ell}E_{\mathrm{tot},\ell-1}^{1/2}E_{\mathrm{tot},\ell}^{1/2}.$
Moreover,
(58) $\begin{split}\|\vec{e}N\|_{H^{\ell}(\Sigma_{t})}^{2}\leq\|k\|_{H^{\ell}(\Sigma_{t})}^{2}+\|e_{0}\varphi\|_{H^{\ell}(\Sigma_{t})}^{2}+C_{\ell}t^{\varepsilon}E_{\mathrm{tot},\ell}+C_{\ell}E_{\mathrm{tot},\ell-1}^{1/2}E_{\mathrm{tot},\ell}^{1/2}\end{split}$
for all $t\leq t_{\ell}$. Finally,
(59) $\begin{split}\|\vec{e}N\|_{H^{\ell}(\Sigma_{t})}^{2}+\tfrac{1}{2}t^{-2}\|N-1\|_{H^{\ell}(\Sigma_{t})}^{2}\leq{}&2\big(\|k\|_{H^{\ell}(\Sigma_{t})}^{2}+\|e_{0}\varphi\|_{H^{\ell}(\Sigma_{t})}^{2}\big)\\ &\negmedspace+C_{\ell}t^{\varepsilon}E_{\mathrm{tot},\ell}+C_{l}E_{\mathrm{tot},\ell-1}^{1/2}E_{\mathrm{tot},\ell}^{1/2}\end{split}$
for all $t\leq t_{\ell}$.
Remark 30.
One crude special case of (59) is the following:
$E_{(N),\ell}=\|\vec{e}N\|_{H^{\ell}(\Sigma_{t})}^{2}+t^{-2}\|N-1\|_{H^{\ell}(\Sigma_{t})}^{2}\leq C_{\ell}E_{\mathrm{tot},\ell}.$
In particular,
(60) $\hat{E}_{\mathrm{tot},\ell}\leq C_{\ell}E_{\mathrm{tot},\ell}.$
Proof.
As in the proof of [33, Lemma 112, p. 52], applying $E_{\mathbf{I}}$ to (20) and then multiplying with $E_{\mathbf{I}}(N-1)$ yields
$\begin{split}&e_{I}E_{\mathbf{I}}e_{I}(N-1)E_{\mathbf{I}}(N-1)+[E_{\mathbf{I}},e_{I}]e_{I}(N-1)E_{\mathbf{I}}(N-1)-t^{-2}\big(E_{\mathbf{I}}(N-1)\big)^{2}\\ ={}&E_{\mathbf{I}}\big(\gamma_{JII}e_{J}(N)-N\big\{t^{-2}-k_{IJ}k_{IJ}-e_{0}(\varphi)e_{0}(\varphi)+\tfrac{2}{n-1}V\circ\varphi\big\}\big)E_{\mathbf{I}}(N-1).\end{split}$
Integrating the result over $\Sigma$ yields
(61) $\begin{split}&-\textstyle{\int}_{\Sigma}\big(E_{\mathbf{I}}e_{I}(N-1)E_{\mathbf{I}}e_{I}(N-1)+t^{-2}\big(E_{\mathbf{I}}(N-1)\big)^{2}\big){\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ &-\textstyle{\int}_{\Sigma}\big(E_{\mathbf{I}}e_{I}(N-1)[e_{I},E_{\mathbf{I}}](N-1)-[E_{\mathbf{I}},e_{I}]e_{I}(N-1)E_{\mathbf{I}}(N-1)\big){\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ &-\textstyle{\int}_{\Sigma}E_{\mathbf{I}}e_{I}(N-1)E_{\mathbf{I}}(N-1)\mathrm{div}_{h_{\mathrm{ref}}}e_{I}{\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ ={}&\textstyle{\int}_{\Sigma}\big(E_{\mathbf{I}}\big(\gamma_{JII}e_{J}(N)-N\big\{t^{-2}-k_{IJ}k_{IJ}-e_{0}(\varphi)e_{0}(\varphi)+\tfrac{2}{n-1}V\circ\varphi\big\}\big)E_{\mathbf{I}}(N-1)\big){\mathrm{d}}\mu_{h_{\mathrm{ref}}},\end{split}$
where we used (141). Next, note that, due to (39),
(62) $\begin{split}&\textstyle{\sum}_{I}\sum_{|\mathbf{I}|\leq\ell}\big|\textstyle{\int}_{\Sigma}E_{\mathbf{I}}e_{I}(N-1)E_{\mathbf{I}}(N-1)\mathrm{div}_{h_{\mathrm{ref}}}e_{I}{\mathrm{d}}\mu_{h_{\mathrm{ref}}}\big|\leq Ct^{3\varepsilon}E_{(N),\ell}.\end{split}$
Next, (25c) and (56) yield
$\|[E_{\mathbf{I}},e_{I}](N-1)\|_{L^{2}(\Sigma_{t})}\leq C_{\ell}t^{3\varepsilon}E_{(N),\ell}^{1/2}+C_{\ell}t^{2\varepsilon}E_{(e,\omega),\ell}^{1/2}.$
Thus
(63) $\begin{split}&\big|\textstyle{\sum}_{|\mathbf{I}|\leq\ell}\sum_{I}\int_{\Sigma_{t}}E_{\mathbf{I}}e_{I}(N-1)[e_{I},E_{\mathbf{I}}](N-1){\mathrm{d}}\mu_{h_{\mathrm{ref}}}\big|\leq C_{\ell}t^{2\varepsilon}(E_{(N),\ell}+E_{(e,\omega),\ell}).\end{split}$
Next, combining (40) and (56) yields
$\|[E_{\mathbf{I}},e_{I}]e_{I}(N-1)\|_{L^{2}(\Sigma_{t})}\leq C_{\ell}t^{-1+3\varepsilon}\|\vec{e}N\|_{H^{\ell}(\Sigma_{t})}+C_{\ell}t^{-1+5\varepsilon}E_{(e,\omega),\ell}^{1/2}.$
Thus
(64) $\begin{split}&\big|\textstyle{\sum}_{|\mathbf{I}|\leq\ell}\sum_{I}\textstyle{\int}_{\Sigma}[E_{\mathbf{I}},e_{I}]e_{I}(N-1)E_{\mathbf{I}}(N-1){\mathrm{d}}\mu_{h_{\mathrm{ref}}}\big|\leq C_{\ell}t^{3\varepsilon}(E_{(N),\ell}+E_{(e,\omega),\ell}).\end{split}$
Next, note that $\gamma_{JII}e_{J}(N)$ and $\tfrac{2N}{n-1}V\circ\varphi$ are good terms. In particular, (48) yields
(65) $\|E_{\mathbf{I}}(\gamma_{JII}e_{J}(N))\|_{L^{2}(\Sigma_{t})}+\|E_{\mathbf{I}}\big(\tfrac{2N}{n-1}V\circ\varphi\big)\|_{L^{2}(\Sigma_{t})}\leq C_{\ell}t^{-1+\varepsilon}\hat{E}_{\mathrm{tot},\ell}^{1/2}.$
Combining (61), (62), (63), (64) and (65) yields
(66) $\begin{split}&E_{(N),\ell}+\textstyle{\sum}_{|\mathbf{I}|\leq\ell}\textstyle{\int}_{\Sigma}E_{\mathbf{I}}\big(N\big\{-t^{-2}+k_{IJ}k_{IJ}+e_{0}(\varphi)e_{0}(\varphi)\big\}\big)E_{\mathbf{I}}(N-1){\mathrm{d}}\mu_{h_{\mathrm{ref}}}=R_{1},\end{split}$
where
$|R_{1}|\leq C_{\ell}t^{\varepsilon}\hat{E}_{\mathrm{tot},\ell}.$
Next, we wish to estimate $E_{\mathbf{I}}(NR)$ in $L^{2}$, where
$R:=-t^{-2}+k_{IJ}k_{IJ}+e_{0}(\varphi)e_{0}(\varphi).$
Note that $E_{\mathbf{I}}(NR)$ can be written as a sum of $NE_{\mathbf{I}}R$ and a linear combination of terms of the form
$E_{\mathbf{J}}E_{j}N\cdot E_{\mathbf{K}}R,$
where $|\mathbf{J}|+|\mathbf{K}|=|\mathbf{I}|-1$. Due to Lemma 57, (25c) and (26),
$\begin{split}\|E_{\mathbf{J}}E_{j}N\cdot E_{\mathbf{K}}R\|_{L^{2}(\Sigma_{t})}&\leq C_{\ell}\|N-1\|_{C^{1}(\Sigma_{t})}\textstyle{\sum}_{i}\|E_{i}R\|_{H^{\ell-2}(\Sigma_{t})}+C_{\ell}\|R\|_{C^{1}(\Sigma_{t})}\|N-1\|_{H^{\ell}(\Sigma_{t})}\\ &\leq C_{\ell}t^{\varepsilon}\textstyle{\sum}_{i}\|E_{i}R\|_{H^{\ell-2}(\Sigma_{t})}+C_{\ell}t^{-2+2\varepsilon}\|N-1\|_{H^{\ell}(\Sigma_{t})},\end{split}$
where the first term on the right hand sides can be omitted if $|\mathbf{K}|=0$. Moreover,
$\|E_{i}R\|_{H^{\ell-2}(\Sigma_{t})}\leq C_{\ell}t^{-1}(\|k\|_{H^{\ell-1}(\Sigma_{t})}+\|e_{0}\varphi\|_{H^{\ell-1}(\Sigma_{t})})\leq C_{\ell}t^{-1}E_{\mathrm{tot},\ell-1}^{1/2},$
where we appealed to (25b) and Gagliardo-Nirenberg estimates. Summing up,
$\begin{split}\big|\textstyle{\int}_{\Sigma}E_{\mathbf{J}}E_{j}N\cdot E_{\mathbf{K}}R\cdot E_{\mathbf{I}}(N-1){\mathrm{d}}\mu_{h_{\mathrm{ref}}}\big|\leq C_{\ell}t^{\varepsilon}E_{\mathrm{tot},\ell-1}^{1/2}E_{(N),\ell}^{1/2}+C_{\ell}t^{2\varepsilon}E_{(N),\ell}.\end{split}$
Note also that, due to (25c) and (26),
$\big|\textstyle{\int}_{\Sigma}NR(N-1){\mathrm{d}}\mu_{h_{\mathrm{ref}}}\big|\leq Ct^{-2+3\varepsilon}.$
What remains is thus to consider expressions of the form
$\begin{split}&\textstyle{\int}_{\Sigma}NE_{\mathbf{I}}\big(k_{IJ}k_{IJ}+e_{0}(\varphi)e_{0}(\varphi)\big)E_{\mathbf{I}}(N-1){\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ ={}&\textstyle{\int}_{\Sigma}2\big(k_{IJ}E_{\mathbf{I}}k_{IJ}+e_{0}(\varphi)E_{\mathbf{I}}e_{0}(\varphi)\big)E_{\mathbf{I}}(N-1){\mathrm{d}}\mu_{h_{\mathrm{ref}}}+R_{2},\end{split}$
where $\mathbf{I}\neq 0$. Similarly to the discussion of the leading order terms, see Subsection 2.4,
$|R_{2}|\leq C_{\ell}t^{\varepsilon}\hat{E}_{\mathrm{tot},\ell}+C_{\ell}E_{\mathrm{tot},\ell-1}^{1/2}E_{(N),\ell}^{1/2}.$
Summing up, we conclude that
(67) $\begin{split}&E_{(N),\ell}+2\textstyle{\sum}_{0<|\mathbf{I}|\leq\ell}\textstyle{\int}_{\Sigma}\big(k_{IJ}E_{\mathbf{I}}k_{IJ}+e_{0}(\varphi)E_{\mathbf{I}}e_{0}(\varphi)\big)E_{\mathbf{I}}(N-1){\mathrm{d}}\mu_{h_{\mathrm{ref}}}=R_{3},\end{split}$
where
(68) $|R_{3}|\leq C_{\ell}t^{\varepsilon}\hat{E}_{\mathrm{tot},\ell}+C_{\ell}E_{\mathrm{tot},\ell-1}^{1/2}E_{(N),\ell}^{1/2}+Ct^{-2+3\varepsilon}.$
At this point, it is convenient to remember that there is a $t_{a}>0$ such that for $t\leq t_{a}$,
$t^{2}k_{IJ}k_{IJ}+t^{2}e_{0}(\varphi)e_{0}(\varphi)\geq\tfrac{1}{2};$
this is an immediate consequence of (26). Integrating this inequality over $\Sigma$ yields the existence of a constant $c_{a}>0$ such that
$t^{-2}\leq c_{a}\textstyle{\int}_{\Sigma}[k_{IJ}k_{IJ}+e_{0}(\varphi)e_{0}(\varphi)]{\mathrm{d}}\mu_{h_{\mathrm{ref}}}\leq c_{a}E_{\mathrm{tot},\ell}.$
In particular, this means that for $t\leq t_{a}$, we can remove the last term on the right hand side of (68). Next, note that
$|2k_{IJ}E_{\mathbf{I}}k_{IJ}+2e_{0}\varphi\cdot E_{\mathbf{I}}e_{0}\varphi|\leq 2\big(\textstyle{\sum}_{I,J}k_{IJ}^{2}+(e_{0}\varphi)^{2}\big)^{1/2}\big(\textstyle{\sum}_{I,J}(E_{\mathbf{I}}k_{IJ})^{2}+(E_{\mathbf{I}}e_{0}\varphi)^{2}\big)^{1/2}$
This means that
$\begin{split}&\big|\textstyle{\sum}_{0<|\mathbf{I}|\leq\ell}\textstyle{\int}_{\Sigma}\big(2k_{IJ}E_{\mathbf{I}}k_{IJ}+2e_{0}\varphi\cdot E_{\mathbf{I}}e_{0}\varphi\big)E_{\mathbf{I}}(N-1){\mathrm{d}}\mu_{h_{\mathrm{ref}}}\big|\\ \leq{}&2\big\|\big(\textstyle{\sum}_{I,J}k_{IJ}^{2}+(e_{0}\varphi)^{2}\big)^{1/2}\big\|_{C^{0}(\Sigma_{t})}\big(\|k\|_{H^{\ell}(\Sigma_{t})}^{2}+\|e_{0}\varphi\|_{H^{\ell}(\Sigma_{t})}^{2}\big)^{1/2}\|N-1\|_{H^{\ell}(\Sigma_{t})}.\end{split}$
On the other hand, (26) yields
(69) $\begin{split}t\big\|\big(\textstyle{\sum}_{I,J}k_{IJ}^{2}+(e_{0}\varphi)^{2}\big)^{1/2}\big\|_{C^{0}(\Sigma_{t})}&\leq 1+\big\|\big(t^{2}\textstyle{\sum}_{I,J}k_{IJ}^{2}+t^{2}(e_{0}\varphi)^{2}\big)^{1/2}-1\big\|_{C^{0}(\Sigma_{t})}\\ &\leq 1+Ct^{2\varepsilon}.\end{split}$
To conclude
$\begin{split}&\big|\textstyle{\sum}_{0<|\mathbf{I}|\leq\ell}\textstyle{\int}_{\Sigma}\big(2k_{IJ}E_{\mathbf{I}}k_{IJ}+2e_{0}\varphi\cdot E_{\mathbf{I}}e_{0}\varphi\big)E_{\mathbf{I}}(N-1){\mathrm{d}}\mu_{h_{\mathrm{ref}}}\big|\\ \leq{}&2t^{-1}(1+Ct^{\varepsilon})\big(\|k\|_{H^{\ell}(\Sigma_{t})}^{2}+\|e_{0}\varphi\|_{H^{\ell}(\Sigma_{t})}^{2}\big)^{1/2}\|N-1\|_{H^{\ell}(\Sigma_{t})}.\end{split}$
Combining this estimate with (67) yields the conclusion that, for $t\leq t_{a}$,
(70) $\begin{split}E_{(N),\ell}&\leq 2t^{-1}\big(\|k\|_{H^{\ell}(\Sigma_{t})}^{2}+\|e_{0}\varphi\|_{H^{\ell}(\Sigma_{t})}^{2}\big)^{1/2}\|N-1\|_{H^{\ell}(\Sigma_{t})}\\ &\quad+C_{\ell}t^{\varepsilon}(E_{(N),\ell}+E_{\mathrm{tot},\ell})+C_{\ell}E_{\mathrm{tot},\ell-1}^{1/2}E_{(N),\ell}^{1/2}.\end{split}$
Using the estimate $2ab\leq a^{2}/2+2b^{2}$ on the first term on the right hand side, it follows that
$\begin{split}\|\vec{e}N\|_{H^{\ell}(\Sigma_{t})}^{2}+\tfrac{1}{2}t^{-2}\|N-1\|_{H^{\ell}(\Sigma_{t})}^{2}&\leq 2\big(\|k\|_{H^{\ell}(\Sigma_{t})}^{2}+\|e_{0}\varphi\|_{H^{\ell}(\Sigma_{t})}^{2}\big)\\ &\quad+C_{\ell}t^{\varepsilon}(E_{(N),\ell}+E_{\mathrm{tot},\ell})+C_{\ell}E_{\mathrm{tot},\ell-1}^{1/2}E_{(N),\ell}^{1/2}\end{split}$
for $t\leq t_{a}$. By a similar argument, it is, from this estimate, straightforward to deduce that there is, for each $\ell$, a $t_{\ell}>0$ and a $C_{\ell}$ such that
(71) $E_{(N),\ell}\leq C_{\ell}E_{\mathrm{tot},\ell}$
for all $t\leq t_{\ell}$. Inserting this information into (70) yields
(72) $\begin{split}E_{(N),\ell}&\leq 2t^{-1}\big(\|k\|_{H^{\ell}(\Sigma_{t})}^{2}+\|e_{0}\varphi\|_{H^{\ell}(\Sigma_{t})}^{2}\big)^{1/2}\|N-1\|_{H^{\ell}(\Sigma_{t})}\\ &\quad+C_{\ell}t^{\varepsilon}E_{\mathrm{tot},\ell}+C_{\ell}E_{\mathrm{tot},\ell-1}^{1/2}E_{\mathrm{tot},\ell}^{1/2}.\end{split}$
Finally, combining (67) with (71) yields (57). Moreover, (72) and arguments similar to the above yield (58) and (59). The estimate (60) follows from (71). The lemma follows ∎
2.6. Estimating the frame
Next, we estimate the time derivative of $E_{(e,\omega),\ell}$ from below.
Lemma 31.
Consider a solution to (11)–(19) as described in Subsection 1.4. Assume, in addition, that $V$ is a $\sigma_{V}$-admissible potential and that the basic supremum assumptions, see Definition 17, are satisfied with $3\varepsilon\leq\sigma_{V}$. Then, for each $\ell\in\mathbb{N}$, there are constants $C_{\ell}$ and $t_{\ell}>0$ such that for $t<t_{\ell}$,
$\partial_{t}E_{(e,\omega),\ell}\geq-2(1+\varepsilon)t^{-1}E_{(e,\omega),\ell}-C_{\ell}t^{-1+\varepsilon}E_{\mathrm{tot},\ell}-C_{\ell}t^{-1}E_{\mathrm{tot},\ell-1}^{1/2}E_{\mathrm{tot},\ell}^{1/2}.$
Proof.
When the time derivative hits the factor $t^{-2\varepsilon}$, the result is $2\varepsilon t^{-1}E_{(e,\omega),\ell}$. Next, let us focus on the energy associated with the frame. This means we have to estimate
(73) $2t^{-2\varepsilon}\textstyle{\sum}_{I,i}\textstyle{\int}_{\Sigma}E_{\mathbf{I}}e_{I}^{i}\cdot E_{\mathbf{I}}\partial_{t}e_{I}^{i}{\mathrm{d}}\mu_{h_{\mathrm{ref}}}=2t^{-2\varepsilon}\textstyle{\sum}_{I,J,i}\textstyle{\int}_{\Sigma}E_{\mathbf{I}}e_{I}^{i}\cdot E_{\mathbf{I}}(-Nk_{IJ}e_{J}^{i}){\mathrm{d}}\mu_{h_{\mathrm{ref}}}.$
The expression $Nk_{IJ}t^{-\varepsilon}e_{J}^{i}$ is a bad term. In particular, due to the arguments presented in Subsection 2.4, it can be written
$E_{\mathbf{I}}(Nk_{IJ}t^{-\varepsilon}e_{J}^{i})=k_{IJ}t^{-\varepsilon}E_{\mathbf{I}}e_{J}^{i}+R_{B},$
where $R_{B}$ satisfies the estimate
$\|R_{B}\|_{L^{2}(\Sigma_{t})}\leq C_{\ell}t^{-1+\varepsilon}E_{\mathrm{tot},\ell}^{1/2}+C_{\ell}t^{-1}E_{\mathrm{tot},\ell-1}^{1/2};$
cf. (51) and (60). Next, due to (69),
$2t^{-2\varepsilon}\textstyle{\sum}_{I,J,i}\textstyle{\int}_{\Sigma}k_{IJ}E_{\mathbf{I}}e_{I}^{i}\cdot E_{\mathbf{I}}e_{J}^{i}{\mathrm{d}}\mu_{h_{\mathrm{ref}}}\geq-2(1+Ct^{\varepsilon})t^{-1}t^{-2\varepsilon}\textstyle{\sum}_{I,i}\textstyle{\int}_{\Sigma}|E_{\mathbf{I}}e_{I}^{i}|^{2}{\mathrm{d}}\mu_{h_{\mathrm{ref}}}$
Due to this estimate and an analogous estimate concerning the energy for $\omega$, the lemma follows. ∎
2.7. Estimating the second fundamental form and the structure coefficients
Next, we estimate the energy associated with the second fundamental form and the structure coefficients.
Lemma 32.
Consider a solution to (11)–(19) as described in Subsection 1.4. Assume, in addition, that $V$ is a $\sigma_{V}$-admissible potential and that the basic supremum assumptions, see Definition 17, are satisfied with $3\varepsilon\leq\sigma_{V}$. Then, for each $\ell\in\mathbb{N}$, there is a $t_{\ell}>0$ such that for $t\leq t_{\ell}$,
(74) $\begin{split}\partial_{t}E_{(\gamma,k),\ell}=&-\textstyle{\sum}_{I,J,K,L}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}k_{IL}E_{\mathbf{I}}(\gamma_{IJK})E_{\mathbf{I}}(\gamma_{LJK}){\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ &-\textstyle{\sum}_{I,J,K,L}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}k_{JL}E_{\mathbf{I}}(\gamma_{IJK})E_{\mathbf{I}}(\gamma_{ILK}){\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ &+\textstyle{\sum}_{I,J,K,L}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}k_{KL}E_{\mathbf{I}}(\gamma_{IJK})E_{\mathbf{I}}(\gamma_{IJL}){\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ &-2t^{-1}\textstyle{\sum}_{I,J}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}|E_{\mathbf{I}}(k_{IJ})|^{2}{\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ &-2t^{-1}\textstyle{\sum}_{I,J}\sum_{0<|\mathbf{I}|\leq\ell}\int_{\Sigma}k_{IJ}E_{\mathbf{I}}(N)E_{\mathbf{I}}(k_{IJ}){\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ &-2\textstyle{\sum}_{I,J,L}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}k_{LJ}E_{\mathbf{I}}(\gamma_{LII})E_{\mathbf{I}}(e_{J}(N)){\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ &-2\textstyle{\sum}_{J}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}e_{0}(\varphi)E_{\mathbf{I}}[e_{J}(\varphi)]E_{\mathbf{I}}(e_{J}(N)){\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ &+2\textstyle{\sum}_{I,J,L,K}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}(k_{LJ}E_{\mathbf{I}}(\gamma_{LII})+k_{IL}E_{\mathbf{I}}(\gamma_{IJL}))E_{\mathbf{I}}(\gamma_{JKK}){\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ &+2\textstyle{\sum}_{J,K}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}e_{0}(\varphi)E_{\mathbf{I}}[e_{J}(\varphi)]E_{\mathbf{I}}(\gamma_{JKK}){\mathrm{d}}\mu_{h_{\mathrm{ref}}}+\mathcal{R},\end{split}$
where
$|\mathcal{R}|\leq C_{\ell}t^{-1+\varepsilon}E_{\mathrm{tot},\ell}+C_{\ell}t^{-1}E_{\mathrm{tot},\ell-1}^{1/2}E_{\mathrm{tot},\ell}^{1/2}.$
Proof.
Note, to begin with, that (13) and (14) can be written
(75) $\partial_{t}\gamma_{IJK}$ $=-2e_{[I}(Nk_{J]K})+R_{\gamma,IJK},$
(76) $\partial_{t}k_{IJ}$ $=e_{(I}e_{J)}(N)-e_{K}(N\gamma_{K(IJ)})-Ne_{(I}(\gamma_{J)KK})+R_{k,IJ},$
where
(77a) $R_{\gamma,IJK}:=$ $-Nk_{IL}\gamma_{LJK}-Nk_{JL}\gamma_{ILK}+Nk_{KL}\gamma_{IJL},$
(77b) $R_{k,IJ}:=$ $-t^{-1}Nk_{IJ}+N\gamma_{KLL}\gamma_{K(IJ)}+N\gamma_{I(KL)}\gamma_{J(KL)}-\tfrac{1}{4}N\gamma_{KLI}\gamma_{KLJ}$
$+Ne_{I}(\varphi)e_{J}(\varphi)+\tfrac{2N}{n-1}(V\circ\varphi)\delta_{IJ}.$
Differentiating the energy for $k$ and $\gamma$ then yields
$\begin{split}\partial_{t}E_{(\gamma,k),\ell}:={}&\textstyle{\sum}_{I,J,K}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}E_{\mathbf{I}}(\gamma_{IJK})E_{\mathbf{I}}(-2e_{[I}(Nk_{J]K})+R_{\gamma,IJK}){\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ &\negmedspace+2\textstyle{\sum}_{I,J}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}E_{\mathbf{I}}(k_{IJ})E_{\mathbf{I}}(e_{(I}e_{J)}(N)-e_{K}(N\gamma_{K(IJ)})){\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ &\negmedspace+2\textstyle{\sum}_{I,J}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}E_{\mathbf{I}}(k_{IJ})E_{\mathbf{I}}(-Ne_{(I}(\gamma_{J)KK})+R_{k,IJ}){\mathrm{d}}\mu_{h_{\mathrm{ref}}}.\end{split}$
The terms arising from $R_{\gamma,IJK}$ and $R_{k,IJ}$ can be estimated directly in terms of the energy. However, the remaining terms contain too many derivatives. In order to estimate the corresponding expressions, we need to integrate by parts, appeal to the momentum constraint, etc. In order to emphasise this division it is useful to, using the symmetries of $\gamma_{IJK}$ and $k_{IJ}$, rewrite the time derivative of the energy as
$\begin{split}\partial_{t}E_{(\gamma,k),\ell}:={}&\textstyle{\sum}_{I,J,K}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}E_{\mathbf{I}}(\gamma_{IJK})E_{\mathbf{I}}(-2e_{I}(Nk_{JK})){\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ &\negmedspace+2\textstyle{\sum}_{I,J}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}E_{\mathbf{I}}(k_{IJ})E_{\mathbf{I}}(e_{I}e_{J}(N)-e_{K}(N\gamma_{KIJ})){\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ &\negmedspace+2\textstyle{\sum}_{I,J}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}E_{\mathbf{I}}(k_{IJ})E_{\mathbf{I}}(-Ne_{I}(\gamma_{JKK})){\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ &\negmedspace+\textstyle{\sum}_{I,J,K}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}E_{\mathbf{I}}(\gamma_{IJK})E_{\mathbf{I}}(R_{\gamma,IJK}){\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ &\negmedspace+2\textstyle{\sum}_{I,J}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}E_{\mathbf{I}}(k_{IJ})E_{\mathbf{I}}(R_{k,IJ}){\mathrm{d}}\mu_{h_{\mathrm{ref}}}.\end{split}$
Let us begin by considering
(78) ${\sum}_{I,J,K}\sum_{|\mathbf{I}|\leq\ell}\textstyle{\int}_{\Sigma}E_{\mathbf{I}}(\gamma_{IJK})E_{\mathbf{I}}(-2e_{I}(Nk_{JK})){\mathrm{d}}\mu_{h_{\mathrm{ref}}}.$
Due to (25b), (25c), (54), (56) and (60), this expression equals
(79) $2\textstyle{\sum}_{I,J,K}\sum_{|\mathbf{I}|\leq\ell}\textstyle{\int}_{\Sigma}E_{\mathbf{I}}(\gamma_{IJK})e_{I}E_{\mathbf{I}}(Nk_{JK}){\mathrm{d}}\mu_{h_{\mathrm{ref}}}$
up to terms that can be estimated by $C_{\ell}t^{-1+\varepsilon}E_{\mathrm{tot},\ell}$. Next, note that $e_{I}E_{\mathbf{I}}(Nk_{JK})$ can be written as a sum of
(80) $e_{I}[NE_{\mathbf{I}}(k_{JK})],\ \ E_{\mathbf{I}}(e_{I}N)k_{JK},$
of
(81) $e_{I},E_{\mathbf{I}}](N)k_{JK},\ \ E_{\mathbf{I}}(N)e_{I}k_{JK}$
and linear combinations of terms of the form
(82) $e_{I}[E_{\mathbf{I}_{1}}E_{i_{1}}(N)E_{\mathbf{I}_{2}}E_{i_{2}}(k_{JK})].$
Moreover, if $\mathbf{I}=0$, then the only term we need to handle is the first expression in (80). The contribution to (79) from the first expression in (80) can be rewritten by integrating by parts; cf. (141). When doing so, there is one term involving the divergence of $e_{I}$. Due to (39), this term can be estimated in absolute value by $C_{\ell}t^{-1+3\varepsilon}E_{\mathrm{tot},\ell}$. What remains to be considered is
$2\textstyle{\sum}_{I,J,K}\sum_{|\mathbf{I}|\leq\ell}\textstyle{\int}_{\Sigma}e_{I}E_{\mathbf{I}}(\gamma_{IJK})NE_{\mathbf{I}}(k_{JK}){\mathrm{d}}\mu_{h_{\mathrm{ref}}}.$
Appealing to (25a) and (56), we can commute $e_{I}$ and $E_{\mathbf{I}}$ in the first factor in the integrand; the commutator term can be estimated by $C_{\ell}t^{-1+3\varepsilon}E_{\mathrm{tot},\ell}$. What remains is then
(83) $2\textstyle{\sum}_{I,J,K}\sum_{|\mathbf{I}|\leq\ell}\textstyle{\int}_{\Sigma}E_{\mathbf{I}}(e_{I}\gamma_{IJK})NE_{\mathbf{I}}(k_{JK}){\mathrm{d}}\mu_{h_{\mathrm{ref}}}.$
The contribution (79) from the second expression in (80), we keep as it is:
(84) $2\textstyle{\sum}_{I,J,K}\sum_{|\mathbf{I}|\leq\ell}\textstyle{\int}_{\Sigma}E_{\mathbf{I}}(\gamma_{IJK})E_{\mathbf{I}}(e_{I}N)k_{JK}{\mathrm{d}}\mu_{h_{\mathrm{ref}}}.$
In order to estimate the contribution to (79) from the first expression in (81), it is sufficient to appeal to (25b), (25c), (56) and (60). The contribution can be estimated by
(85) $C_{\ell}t^{-1+\varepsilon}E_{\mathrm{tot},\ell}$
(in fact, a better estimate holds). The contribution to (79) from the second expression in (81) can, keeping in mind that we can assume that $\mathbf{I}\neq 0$, also be bounded by (85); cf. (25a), (25b) and (60). Finally, due to Gagliardo-Nirenberg estimates, (25) and (60), the contribution from (82) can be bounded by (85). To summarise, up to terms that can be estimated in absolute value by (85), the expression (78) can be written as a sum of (83) and (84).
Next, consider
(86) $2\textstyle{\sum}_{I,J}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}E_{\mathbf{I}}(k_{IJ})E_{\mathbf{I}}(-e_{K}(N\gamma_{KIJ})){\mathrm{d}}\mu_{h_{\mathrm{ref}}}.$
Note that $E_{\mathbf{I}}(-e_{K}(N\gamma_{KIJ}))$ can be written as a sum of
(87) $E_{\mathbf{I}}(-e_{K}(N)\gamma_{KIJ}),\ \ \ E_{\mathbf{I}}(-Ne_{K}\gamma_{KIJ})+NE_{\mathbf{I}}(e_{K}\gamma_{KIJ}),$
and
$NE_{\mathbf{I}}(e_{K}\gamma_{KIJ}).$
Since the first expression in (87) is a good term, the contribution to (86) from the first expression in (87) can be bounded by (85); cf. (48) and (60). Appealing to Gagliardo-Nirenberg estimates, (25) and (60), the contribution to (86) from the second expression in (87) can be bounded by (85) as well. What remains is
$2\textstyle{\sum}_{I,J,K}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}E_{\mathbf{I}}(k_{IJ})NE_{\mathbf{I}}(e_{K}\gamma_{KIJ}){\mathrm{d}}\mu_{h_{\mathrm{ref}}}.$
Clearly, this term cancels (83). To summarise,
(88) $\begin{split}&\textstyle{\sum}_{I,J,K}\sum_{|\mathbf{I}|\leq\ell}\textstyle{\int}_{\Sigma}E_{\mathbf{I}}(\gamma_{IJK})E_{\mathbf{I}}(-2e_{I}(Nk_{JK})){\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ &+2\textstyle{\sum}_{I,J,K}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}E_{\mathbf{I}}(k_{IJ})E_{\mathbf{I}}(-e_{K}(N\gamma_{KIJ})){\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ =&-2\textstyle{\sum}_{I,J,K}\sum_{|\mathbf{I}|\leq\ell}\textstyle{\int}_{\Sigma}E_{\mathbf{I}}(\gamma_{IJK})E_{\mathbf{I}}(e_{I}N)k_{JK}{\mathrm{d}}\mu_{h_{\mathrm{ref}}}+\dots,\end{split}$
where the dots signify terms that can be estimated in absolute value by (85).
Next, let us turn to
(89) $2\textstyle{\sum}_{I,J}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}E_{\mathbf{I}}(k_{IJ})E_{\mathbf{I}}(e_{I}e_{J}(N)){\mathrm{d}}\mu_{h_{\mathrm{ref}}}.$
Note that $E_{\mathbf{I}}(e_{I}e_{J}(N))$ can be written as a sum of
(90) $E_{\mathbf{I}},e_{I}](e_{J}N),\ \ \ e_{I}E_{\mathbf{I}}(e_{J}N).$
Due to (40), (56) and (60), the contribution to (89) from the first expression in (90) can be bounded by (85). The contribution of the second expression in (90) to (89) can be estimated by first integrating by parts. This results in two terms. The term which involves the divergence of $e_{I}$ can be bounded by (85). The remaining term is
$2\textstyle{\sum}_{I,J}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}e_{I}E_{\mathbf{I}}(k_{IJ})E_{\mathbf{I}}(e_{J}(N)){\mathrm{d}}\mu_{h_{\mathrm{ref}}}.$
Commuting $e_{I}$ and $E_{\mathbf{I}}$ in the first factor in the integrand leads to terms that can be estimated in absolute value by (85); cf. (25b) and (56). This means that we are left with the problem of estimating
(91) $2\textstyle{\sum}_{I,J}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}E_{\mathbf{I}}(e_{I}k_{IJ})E_{\mathbf{I}}(e_{J}(N)){\mathrm{d}}\mu_{h_{\mathrm{ref}}}.$
In order to estimate this expression, we appeal to (18). However, we only write down the details later.
Next, consider
(92) $2\textstyle{\sum}_{I,J,K}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}E_{\mathbf{I}}(k_{IJ})E_{\mathbf{I}}(-Ne_{I}(\gamma_{JKK})){\mathrm{d}}\mu_{h_{\mathrm{ref}}}.$
Note that
(93) $Ne_{I}(\gamma_{JKK})=e_{I}(N)\gamma_{JKK}-e_{I}(N\gamma_{JKK}).$
The first term on the right hand side of (93) is a good term, and the corresponding contribution to (92) can be estimated by (85) due to (48) and (60). What remains is
$2\textstyle{\sum}_{I,J,K}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}E_{\mathbf{I}}(k_{IJ})E_{\mathbf{I}}e_{I}(N\gamma_{JKK}){\mathrm{d}}\mu_{h_{\mathrm{ref}}}.$
At this point we can commute $E_{\mathbf{I}}$ and $e_{I}$, integrate by parts, and commute $e_{I}$ and $E_{\mathbf{I}}$ again. This results, by arguments similar to the above, in
(94) $2\textstyle{\sum}_{I,J,K}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}E_{\mathbf{I}}(e_{I}k_{IJ})E_{\mathbf{I}}(N\gamma_{JKK}){\mathrm{d}}\mu_{h_{\mathrm{ref}}}$
plus terms that can be estimated in absolute value by (85). Again, we can appeal to (18) in order to estimate this expression.
Next, let us turn to
(95) ${\sum}_{I,J,K}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}E_{\mathbf{I}}(\gamma_{IJK})E_{\mathbf{I}}(R_{\gamma,IJK}){\mathrm{d}}\mu_{h_{\mathrm{ref}}},$
where $R_{\gamma,IJK}$ is introduced in (77a). The three terms appearing in $R_{\gamma,IJK}$ are all bad terms; cf. Subsection 2.4. In fact, due to (50), (51) and (60),
$\begin{split}&\textstyle{\sum}_{I,J,K}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}E_{\mathbf{I}}(\gamma_{IJK})E_{\mathbf{I}}(R_{\gamma,IJK}){\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ =&-\textstyle{\sum}_{I,J,K,L}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}k_{IL}E_{\mathbf{I}}(\gamma_{IJK})E_{\mathbf{I}}(\gamma_{LJK}){\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ &-\textstyle{\sum}_{I,J,K,L}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}k_{JL}E_{\mathbf{I}}(\gamma_{IJK})E_{\mathbf{I}}(\gamma_{ILK}){\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ &+\textstyle{\sum}_{I,J,K,L}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}k_{KL}E_{\mathbf{I}}(\gamma_{IJK})E_{\mathbf{I}}(\gamma_{IJL}){\mathrm{d}}\mu_{h_{\mathrm{ref}}}+\dots,\end{split}$
where the dots represent terms that can be estimated in absolute value by
(96) $C_{\ell}t^{-1+\varepsilon}E_{\mathrm{tot},\ell}+C_{\ell}t^{-1}E_{\mathrm{tot},\ell-1}^{1/2}E_{\mathrm{tot},\ell}^{1/2}.$
Next, consider
(97) $2\textstyle{\sum}_{I,J}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}E_{\mathbf{I}}(k_{IJ})E_{\mathbf{I}}(R_{k,IJ}){\mathrm{d}}\mu_{h_{\mathrm{ref}}},$
where $R_{k,IJ}$ is given by (77b). The first term appearing in (77b) is a leading order term; see Subsection 2.4. In fact, due to (52), (53) and (60), it gives the following contribution to (97):
$\begin{split}&-2t^{-1}\textstyle{\sum}_{I,J}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}|E_{\mathbf{I}}(k_{IJ})|^{2}{\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ &-2t^{-1}\textstyle{\sum}_{I,J}\sum_{0<|\mathbf{I}|\leq\ell}\int_{\Sigma}k_{IJ}E_{\mathbf{I}}(N)E_{\mathbf{I}}(k_{IJ}){\mathrm{d}}\mu_{h_{\mathrm{ref}}}+\dots,\end{split}$
where the dots represent terms that can be estimated in absolute value by (96). The remaining terms appearing in $R_{k,IJ}$ are all good terms, and their contribution can, due to (48) and (60), be estimated by (96). To summarise,
$\begin{split}&2\textstyle{\sum}_{I,J}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}E_{\mathbf{I}}(k_{IJ})E_{\mathbf{I}}(R_{k,IJ}){\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ =&-2t^{-1}\textstyle{\sum}_{I,J}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}|E_{\mathbf{I}}(k_{IJ})|^{2}{\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ &-2t^{-1}\textstyle{\sum}_{I,J}\sum_{0<|\mathbf{I}|\leq\ell}\int_{\Sigma}k_{IJ}E_{\mathbf{I}}(N)E_{\mathbf{I}}(k_{IJ}){\mathrm{d}}\mu_{h_{\mathrm{ref}}}+\dots,\end{split}$
where the dots represent terms that can be estimated in absolute value by (96). At this point, we return to (91). Due to (18), we need to estimate
(98) $\begin{split}-2\textstyle{\sum}_{J}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}E_{\mathbf{I}}(\gamma_{LII}k_{LJ}+\gamma_{IJL}k_{IL}+e_{0}(\varphi)e_{J}(\varphi))E_{\mathbf{I}}(e_{J}(N)){\mathrm{d}}\mu_{h_{\mathrm{ref}}}.\end{split}$
Since all the terms appearing on the right hand side of (18) are bad terms, (50), (51), (60) and the arguments preceding (91) yield the conclusion that
(99) $\begin{split}&2\textstyle{\sum}_{I,J}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}E_{\mathbf{I}}(k_{IJ})E_{\mathbf{I}}(e_{I}e_{J}(N)){\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ =&-2\textstyle{\sum}_{J}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}(k_{LJ}E_{\mathbf{I}}(\gamma_{LII})-k_{IL}E_{\mathbf{I}}(\gamma_{JIL})+e_{0}(\varphi)E_{\mathbf{I}}[e_{J}(\varphi)])E_{\mathbf{I}}(e_{J}(N)){\mathrm{d}}\mu_{h_{\mathrm{ref}}}+\dots,\end{split}$
where the dots represent terms that can be estimated in absolute value by (96).
Finally, we need to estimate the contribution from (94). Again, due to (18), we need to estimate
$2\textstyle{\sum}_{J,K}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}E_{\mathbf{I}}(\gamma_{LII}k_{LJ}+\gamma_{IJL}k_{IL}+e_{0}(\varphi)e_{J}(\varphi))E_{\mathbf{I}}(N\gamma_{JKK}){\mathrm{d}}\mu_{h_{\mathrm{ref}}}.$
Note that $N\gamma_{JKK}$ can be replaced by $\gamma_{JKK}$ up to an error that can be estimated in absolute value by $C_{\ell}t^{-1+\varepsilon}E_{\mathrm{tot},\ell}$. For that reason, we, from now on, effectively set $N$ equal to $1$. The remaining terms can be estimated as in the argument concerning (91). Combining these arguments with the observations made immediately prior to (94), it follows that
(100) $\begin{split}&2\textstyle{\sum}_{I,J,K}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}E_{\mathbf{I}}(k_{IJ})E_{\mathbf{I}}(-Ne_{I}(\gamma_{JKK})){\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ ={}&2\textstyle{\sum}_{J,K}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}(k_{LJ}E_{\mathbf{I}}(\gamma_{LII})+k_{IL}E_{\mathbf{I}}(\gamma_{IJL})+e_{0}(\varphi)E_{\mathbf{I}}[e_{J}(\varphi)])E_{\mathbf{I}}(\gamma_{JKK}){\mathrm{d}}\mu_{h_{\mathrm{ref}}}+\dots,\end{split}$
where the dots represent terms that can be estimated in absolute value by (96). At this point, we can sum up the estimates in order to obtain (74), where we used the fact that there is a cancellation between the first term on the right hand side of (88) and one of the terms arising from (99). ∎
2.8. The scalar field
Next, we turn to energy estimates for the scalar field.
Lemma 33.
Consider a solution to (11)–(19) as described in Subsection 1.4. Assume, in addition, that $V$ is a $\sigma_{V}$-admissible potential and that the basic supremum assumptions, see Definition 17, are satisfied with $3\varepsilon\leq\sigma_{V}$. Then, for each $\ell\in\mathbb{N}$, there is a $t_{\ell}>0$ such that for $t\leq t_{\ell}$,
(101) $\begin{split}\partial_{t}E_{(\varphi),\ell}=&-2t^{-1}\textstyle{\sum}_{|\mathbf{I}|\leq\ell}\int_{\Sigma}|E_{\mathbf{I}}(e_{0}\varphi)|^{2}{\mathrm{d}}\mu_{h_{\mathrm{ref}}}-2\textstyle{\sum}_{|\mathbf{I}|\leq\ell}\textstyle{\sum}_{I,J}\int_{\Sigma}k_{IJ}(E_{\mathbf{I}}e_{I}\varphi)(E_{\mathbf{I}}e_{J}\varphi){\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ &-2t^{-1}(1-\varepsilon)t^{-2+2\varepsilon}\textstyle{\sum}_{|\mathbf{I}|\leq\ell}\int_{\Sigma}|E_{\mathbf{I}}\varphi|^{2}{\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ &+2\textstyle{\sum}_{|\mathbf{I}|\leq\ell}\textstyle{\sum}_{I}\int_{\Sigma}e_{0}(\varphi)(E_{\mathbf{I}}e_{I}\varphi)[E_{\mathbf{I}}(e_{I}N)]{\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ &-2t^{-1}\textstyle{\sum}_{0<|\mathbf{I}|\leq\ell}\int_{\Sigma}e_{0}(\varphi)E_{\mathbf{I}}(N)E_{\mathbf{I}}(e_{0}\varphi){\mathrm{d}}\mu_{h_{\mathrm{ref}}}+\mathcal{R},\end{split}$
where
$|\mathcal{R}|\leq C_{\ell}t^{-1+\varepsilon}E_{\mathrm{tot},\ell}+C_{\ell}t^{-1}E_{\mathrm{tot},\ell-1}^{1/2}E_{\mathrm{tot},\ell}^{1/2}.$
Proof.
Let us begin by considering the last term on the right hand side of (41c). We wish to differentiate this term with respect to time. When the derivative hits the factor $t^{-2(1-\varepsilon)}$, we simply obtain $2(1-\varepsilon)t^{-1}$ times the last term. Let us therefore focus on
$2t^{-2+2\varepsilon}\textstyle{\sum}_{|\mathbf{I}|\leq\ell}\int_{\Sigma}(E_{\mathbf{I}}\varphi)[E_{\mathbf{I}}(Ne_{0}\varphi)]{\mathrm{d}}\mu_{h_{\mathrm{ref}}}.$
Since $Ne_{0}\varphi$ is a leading order term, (54) and (60) imply that the absolute value of this expression can be estimated by (85). Next, consider
(102) $2\textstyle{\int}_{\Sigma}(E_{\mathbf{I}}e_{0}\varphi)[E_{\mathbf{I}}(Ne_{0}^{2}\varphi)]{\mathrm{d}}\mu_{h_{\mathrm{ref}}}+2\sum_{I}\int_{\Sigma}(E_{\mathbf{I}}e_{I}\varphi)[E_{\mathbf{I}}(\partial_{t}e_{I}\varphi)]{\mathrm{d}}\mu_{h_{\mathrm{ref}}}.$
Due to (15), the second term can be written
(103) $2\textstyle{\sum}_{I}\int_{\Sigma}(E_{\mathbf{I}}e_{I}\varphi)[E_{\mathbf{I}}(e_{I}(Ne_{0}\varphi)-Nk_{IJ}e_{J}(\varphi)]{\mathrm{d}}\mu_{h_{\mathrm{ref}}}.$
Since $Nk_{IJ}e_{J}(\varphi)$ is a bad term, the contribution from the second term in the second factor in the integrand can be written
$2\textstyle{\sum}_{I,J}\int_{\Sigma}k_{IJ}(E_{\mathbf{I}}e_{I}\varphi)(E_{\mathbf{I}}e_{J}\varphi){\mathrm{d}}\mu_{h_{\mathrm{ref}}}+\dots,$
where the dots represent terms that can be estimated in absolute value by (96). Next, consider the contribution from the first term in the second factor in the integrand of (103). It can be written
(104) $2\textstyle{\sum}_{I}\int_{\Sigma}(E_{\mathbf{I}}e_{I}\varphi)[E_{\mathbf{I}}(e_{I}(N)e_{0}\varphi)+E_{\mathbf{I}}(Ne_{I}(e_{0}\varphi))]{\mathrm{d}}\mu_{h_{\mathrm{ref}}}.$
Again, $e_{I}(N)e_{0}\varphi$ is a bad term, and the contribution from the first term in the second factor in the integrand is
$2\textstyle{\sum}_{I}\int_{\Sigma}e_{0}(\varphi)(E_{\mathbf{I}}e_{I}\varphi)[E_{\mathbf{I}}(e_{I}N)]{\mathrm{d}}\mu_{h_{\mathrm{ref}}}+\dots,$
where the dots represent terms that can be estimated in absolute value by (96). The contribution from the second term in the second factor in the integrand of (104) can be written as a sum of
(105) $2\textstyle{\sum}_{I}\int_{\Sigma}(E_{\mathbf{I}}e_{I}\varphi)NE_{\mathbf{I}}[e_{I}(e_{0}\varphi)]{\mathrm{d}}\mu_{h_{\mathrm{ref}}}$
plus a linear combination of terms of the form
(106) $2\textstyle{\sum}_{I}\int_{\Sigma}(E_{\mathbf{I}}e_{I}\varphi)E_{\mathbf{I}_{1}}[E_{i_{1}}(N-1)]E_{\mathbf{I}_{2}}[e_{I}(e_{0}\varphi)]{\mathrm{d}}\mu_{h_{\mathrm{ref}}},$
where $|\mathbf{I}_{1}|+|\mathbf{I}_{2}|=|\mathbf{I}|-1$. By arguments similar to the above, expressions of the form (106) can be estimated in absolute value by (85). Turning to (105), commuting $e_{I}$ and $E_{\mathbf{I}}$ leads to a term that can be estimated in absolute value by (85); cf. (56). We are thus left with
$2\textstyle{\sum}_{I}\int_{\Sigma}(E_{\mathbf{I}}e_{I}\varphi)Ne_{I}[E_{\mathbf{I}}(e_{0}\varphi)]{\mathrm{d}}\mu_{h_{\mathrm{ref}}}.$
Integrating by parts in this expression generates two terms. The term involving the divergence of $e_{I}$ can be estimated in absolute value by $C_{\ell}t^{-1+3\varepsilon}E_{\mathrm{tot},\ell}$. The remaining term can be written
$2\textstyle{\sum}_{I}\int_{\Sigma}e_{I}[N(E_{\mathbf{I}}e_{I}\varphi)]E_{\mathbf{I}}(e_{0}\varphi){\mathrm{d}}\mu_{h_{\mathrm{ref}}}.$
The term that results when $e_{I}$ hits $N$ can be estimated in absolute value by $Ct^{-1+4\varepsilon}E_{\mathrm{tot},\ell}$. Commuting $e_{I}$ and $E_{\mathbf{I}}$ again leads to expressions that can be bounded in absolute value by $C_{\ell}t^{-1+\varepsilon}E_{\mathrm{tot},\ell}$. What remains is then
$2\textstyle{\sum}_{I}\int_{\Sigma}N(E_{\mathbf{I}}e_{I}e_{I}\varphi)E_{\mathbf{I}}(e_{0}\varphi){\mathrm{d}}\mu_{h_{\mathrm{ref}}}.$
Next, $N$ can be commuted through $E_{\mathbf{I}}$ at the price of a term that can be estimated in absolute value by $C_{\ell}t^{-1+\varepsilon}E_{\mathrm{tot},\ell}$. We are left with
$2\textstyle{\sum}_{I}\int_{\Sigma}E_{\mathbf{I}}(Ne_{I}e_{I}\varphi)E_{\mathbf{I}}(e_{0}\varphi){\mathrm{d}}\mu_{h_{\mathrm{ref}}}.$
At this point, we can combine the result with the first term appearing in (102). This leads to
$2\textstyle{\int}_{\Sigma}E_{\mathbf{I}}[N(e_{0}e_{0}\varphi-e_{I}e_{I}\varphi)]E_{\mathbf{I}}(e_{0}\varphi){\mathrm{d}}\mu_{h_{\mathrm{ref}}}.$
Next, we appeal to (16). As a result, we need to estimate
$2\textstyle{\int}_{\Sigma}E_{\mathbf{I}}(-t^{-1}Ne_{0}(\varphi)+e_{I}(N)e_{I}(\varphi)-N\gamma_{JII}e_{J}(\varphi)-NV^{\prime}\circ\varphi)E_{\mathbf{I}}(e_{0}\varphi){\mathrm{d}}\mu_{h_{\mathrm{ref}}}.$
The three last terms inside the paranthesis in the first factor in the integrand are good terms, and their contribution is bounded in absolute value by $C_{\ell}t^{-1+\varepsilon}E_{\mathrm{tot},\ell}$. The first term is a leading order term. Due to (52), (53) and (60), we conclude that
$\begin{split}&2\textstyle{\int}_{\Sigma}E_{\mathbf{I}}[N(e_{0}e_{0}\varphi-e_{I}e_{I}\varphi)]E_{\mathbf{I}}(e_{0}\varphi){\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ =&-2t^{-1}\textstyle{\int}_{\Sigma}|E_{\mathbf{I}}(e_{0}\varphi)|^{2}{\mathrm{d}}\mu_{h_{\mathrm{ref}}}-2t^{-1}\textstyle{\int}_{\Sigma}e_{0}(\varphi)E_{\mathbf{I}}(N)E_{\mathbf{I}}(e_{0}\varphi){\mathrm{d}}\mu_{h_{\mathrm{ref}}}+\dots,\end{split}$
where the dots represent terms that can be estimated in absolute value by (96) and the second term can be dropped if $\mathbf{I}=0$. The lemma follows. ∎
2.9. The energy of $a_{I}$
Next, we consider the evolution of the energy $E_{(a),\ell}$, introduced in (42).
Lemma 34.
Consider a solution to (11)–(19) as described in Subsection 1.4. Assume, in addition, that $V$ is a $\sigma_{V}$-admissible potential and that the basic supremum assumptions, see Definition 17, are satisfied with $3\varepsilon\leq\sigma_{V}$. Then, for each $\ell\in\mathbb{N}$, there is a $t_{\ell}>0$ such that for $t\leq t_{\ell}$,
$\begin{split}\partial_{t}E_{(a),\ell}=&-2t^{-1}\textstyle{\sum}_{|\mathbf{I}|\leq\ell}\sum_{I}\int_{\Sigma}(E_{\mathbf{I}}a_{I})\cdot E_{\mathbf{I}}[e_{I}(N)]{\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ &+2\textstyle{\sum}_{|\mathbf{I}|\leq\ell}\sum_{I,J}\int_{\Sigma}k_{IJ}(E_{\mathbf{I}}a_{I})\cdot E_{\mathbf{I}}[e_{J}(N)]{\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ &+2\textstyle{\sum}_{|\mathbf{I}|\leq\ell}\sum_{I,J,L}\int_{\Sigma}k_{JL}(E_{\mathbf{I}}a_{I})\cdot E_{\mathbf{I}}(\gamma_{JIL}){\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ &+2\textstyle{\sum}_{|\mathbf{I}|\leq\ell}\sum_{I}\int_{\Sigma}(e_{0}\varphi)(E_{\mathbf{I}}a_{I})\cdot E_{\mathbf{I}}(e_{I}\varphi){\mathrm{d}}\mu_{h_{\mathrm{ref}}}+\mathcal{R},\end{split}$
where
(107) $|\mathcal{R}|\leq C_{\ell}t^{-1+\varepsilon}E_{\mathrm{tot},\ell}+C_{\ell}t^{-1}E_{\mathrm{tot},\ell-1}^{1/2}E_{\mathrm{tot},\ell}^{1/2}.$
Proof.
Differentiating $E_{(a),\ell}$ with respect to time yields
$\partial_{t}E_{(a),\ell}=2\textstyle{\sum}_{|\mathbf{I}|\leq\ell}\sum_{I}\int_{\Sigma}(E_{\mathbf{I}}a_{I})E_{\mathbf{I}}(\partial_{t}a_{I}){\mathrm{d}}\mu_{h_{\mathrm{ref}}}.$
The contribution from the first term on the right hand side of (24) is
$2t^{-1}\textstyle{\sum}_{|\mathbf{I}|\leq\ell}\sum_{I}\int_{\Sigma}(E_{\mathbf{I}}a_{I})\cdot E_{\mathbf{I}}[e_{I}(N)]{\mathrm{d}}\mu_{h_{\mathrm{ref}}}.$
The last three terms on the right hand side of (24) are all bad terms. Up to a term which can be estimated in absolute value by (96), the corresponding contributions can therefore be written
$2\textstyle{\sum}_{|\mathbf{I}|\leq\ell}\sum_{I}\int_{\Sigma}(E_{\mathbf{I}}a_{I})[k_{IJ}E_{\mathbf{I}}(e_{J}N)+k_{JL}E_{\mathbf{I}}(\gamma_{JIL})+e_{0}(\varphi)E_{\mathbf{I}}(e_{I}\varphi)]{\mathrm{d}}\mu_{h_{\mathrm{ref}}};$
cf. (50) and (51). The lemma follows. ∎
2.10. Summing up
Finally, we turn to $\mathcal{E}_{\mathrm{tot},\ell}$, introduced in (43).
Lemma 35.
Consider a solution to (11)–(19) as described in Subsection 1.4. Assume, in addition, that $V$ is a $\sigma_{V}$-admissible potential and that the basic supremum assumptions, see Definition 17, are satisfied with $3\varepsilon\leq\sigma_{V}$. Then, for each $\ell\in\mathbb{N}$, there is a $t_{\ell}>0$ such that for $t\leq t_{\ell}$,
(108) $\partial_{t}\mathcal{E}_{\mathrm{tot},\ell}\geq-8t^{-1}\mathcal{E}_{\mathrm{tot},\ell}-C_{\ell}t^{-1+\varepsilon}\mathcal{E}_{\mathrm{tot},\ell}-C_{\ell}t^{-1}\mathcal{E}_{\mathrm{tot},\ell-1}^{1/2}\mathcal{E}_{\mathrm{tot},\ell}^{1/2}.$
Proof.
Combining Lemmas 31, 32 and 33 with (57) yields
$\begin{split}\partial_{t}E_{\mathrm{tot},\ell}\geq&-2t^{-1}(1+\varepsilon)E_{(e,\omega),\ell}-2t^{-1}\textstyle{\sum}_{|\mathbf{I}|\leq\ell}\int_{\Sigma}|E_{\mathbf{I}}(e_{0}\varphi)|^{2}{\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ &-2\textstyle{\sum}_{|\mathbf{I}|\leq\ell}\textstyle{\sum}_{I,J}\int_{\Sigma}k_{IJ}(E_{\mathbf{I}}e_{I}\varphi)(E_{\mathbf{I}}e_{J}\varphi){\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ &-2t^{-1}(1-\varepsilon)t^{-2+2\varepsilon}\textstyle{\sum}_{|\mathbf{I}|\leq\ell}\int_{\Sigma}|E_{\mathbf{I}}\varphi|^{2}{\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ &-\textstyle{\sum}_{I,J,K,L}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}k_{IL}E_{\mathbf{I}}(\gamma_{IJK})E_{\mathbf{I}}(\gamma_{LJK}){\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ &-\textstyle{\sum}_{I,J,K,L}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}k_{JL}E_{\mathbf{I}}(\gamma_{IJK})E_{\mathbf{I}}(\gamma_{ILK}){\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ &+\textstyle{\sum}_{I,J,K,L}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}k_{KL}E_{\mathbf{I}}(\gamma_{IJK})E_{\mathbf{I}}(\gamma_{IJL}){\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ &+2\textstyle{\sum}_{I,J,L,K}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}(k_{LJ}E_{\mathbf{I}}(\gamma_{LII})+k_{IL}E_{\mathbf{I}}(\gamma_{IJL}))E_{\mathbf{I}}(\gamma_{JKK}){\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ &-2t^{-1}\textstyle{\sum}_{I,J}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}|E_{\mathbf{I}}(k_{IJ})|^{2}{\mathrm{d}}\mu_{h_{\mathrm{ref}}}+t^{-1}E_{(N),l}\\ &-2\textstyle{\sum}_{I,J,L}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}k_{LJ}E_{\mathbf{I}}(\gamma_{LII})E_{\mathbf{I}}(e_{J}(N)){\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ &+2\textstyle{\sum}_{J,K}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}e_{0}(\varphi)E_{\mathbf{I}}[e_{J}(\varphi)]E_{\mathbf{I}}(\gamma_{JKK}){\mathrm{d}}\mu_{h_{\mathrm{ref}}}+\mathcal{R},\end{split}$
where $\mathcal{R}$ satisfies an estimate of the form (107). Before proceeding, note that for any $\alpha_{I}$,
(109) $\big|\textstyle{\sum}_{I,J}k_{IJ}\alpha_{I}\alpha_{J}\big|\leq\big(\textstyle{\sum}_{I,J}k_{IJ}^{2}\big)^{1/2}\big(\textstyle{\sum}_{I,J}\alpha_{I}^{2}\alpha_{J}^{2}\big)^{1/2}\leq(1+Ct^{\varepsilon})t^{-1}\textstyle{\sum}_{I}\alpha_{I}^{2},$
where we appealed to (69) in the last step. Keeping this estimate in mind, including $E_{(a),\ell}$ and recalling Lemma 34 then yields
$\begin{split}\partial_{t}\mathcal{E}_{\mathrm{tot},\ell}\geq&-2t^{-1}(1+\varepsilon)E_{(e,\omega),\ell}-2t^{-1}E_{(\varphi),\ell}-2t^{-1}\textstyle{\sum}_{I,J}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}|E_{\mathbf{I}}(k_{IJ})|^{2}{\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ &-\textstyle{\sum}_{I,J,K,L}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}k_{IL}E_{\mathbf{I}}(\gamma_{IJK})E_{\mathbf{I}}(\gamma_{LJK}){\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ &-\textstyle{\sum}_{I,J,K,L}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}k_{JL}E_{\mathbf{I}}(\gamma_{IJK})E_{\mathbf{I}}(\gamma_{ILK}){\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ &+\textstyle{\sum}_{I,J,K,L}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}k_{KL}E_{\mathbf{I}}(\gamma_{IJK})E_{\mathbf{I}}(\gamma_{IJL}){\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ &+2\textstyle{\sum}_{J,L}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}(k_{LJ}E_{\mathbf{I}}(a_{L})+2k_{IL}E_{\mathbf{I}}(\gamma_{IJL}))E_{\mathbf{I}}(a_{J}){\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ &+4\textstyle{\sum}_{J}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}e_{0}(\varphi)E_{\mathbf{I}}[e_{J}(\varphi)]E_{\mathbf{I}}(a_{J}){\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ &-2t^{-1}\textstyle{\sum}_{|\mathbf{I}|\leq\ell}\sum_{I}\int_{\Sigma}(E_{\mathbf{I}}a_{I})\cdot E_{\mathbf{I}}[e_{I}(N)]{\mathrm{d}}\mu_{h_{\mathrm{ref}}}+t^{-1}E_{(N),\ell}+\mathcal{R},\end{split}$
where $\mathcal{R}$ satisfies an estimate of the form (107). Next, note that, due to (109),
$\begin{split}&-\textstyle{\sum}_{I,J,K,L}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}k_{IL}E_{\mathbf{I}}(\gamma_{IJK})E_{\mathbf{I}}(\gamma_{LJK}){\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ &-\textstyle{\sum}_{I,J,K,L}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}k_{JL}E_{\mathbf{I}}(\gamma_{IJK})E_{\mathbf{I}}(\gamma_{ILK}){\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ &+\textstyle{\sum}_{I,J,K,L}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}k_{KL}E_{\mathbf{I}}(\gamma_{IJK})E_{\mathbf{I}}(\gamma_{IJL}){\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ \geq&-\textstyle{\sum}_{I,J,K}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}t^{-1}(3+Ct^{\varepsilon})|E_{\mathbf{I}}(\gamma_{IJK})|^{2}{\mathrm{d}}\mu_{h_{\mathrm{ref}}}.\end{split}$
Moreover,
$\begin{split}&-2t^{-1}\textstyle{\sum}_{|\mathbf{I}|\leq\ell}\sum_{I}\int_{\Sigma}(E_{\mathbf{I}}a_{I})\cdot E_{\mathbf{I}}[e_{I}(N)]{\mathrm{d}}\mu_{h_{\mathrm{ref}}}+t^{-1}E_{(N),\ell}\\ \geq&-t^{-1}\textstyle{\sum}_{|\mathbf{I}|\leq\ell}\sum_{I}\int_{\Sigma}|E_{\mathbf{I}}a_{I}|^{2}{\mathrm{d}}\mu_{h_{\mathrm{ref}}}.\end{split}$
We also have
$\begin{split}&4\textstyle{\sum}_{J}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}e_{0}(\varphi)E_{\mathbf{I}}[e_{J}(\varphi)]E_{\mathbf{I}}(a_{J}){\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ \geq&-t^{-1}\textstyle{\sum}_{J}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}|E_{\mathbf{I}}[e_{J}(\varphi)]|^{2}{\mathrm{d}}\mu_{h_{\mathrm{ref}}}-4t^{-1}\textstyle{\sum}_{J}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}|te_{0}(\varphi)|^{2}|E_{\mathbf{I}}(a_{J})|^{2}{\mathrm{d}}\mu_{h_{\mathrm{ref}}}.\end{split}$
Next,
$\begin{split}&4\textstyle{\sum}_{I,J,L}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}k_{IL}E_{\mathbf{I}}(\gamma_{IJL})E_{\mathbf{I}}(a_{J}){\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ \geq&-t^{-1}\textstyle{\sum}_{I,J,K}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}|E_{\mathbf{I}}(\gamma_{IJK})|^{2}{\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ &-4t^{-1}\textstyle{\sum}_{I,J,L}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}t^{2}k_{IL}k_{IL}|E_{\mathbf{I}}(a_{J})|^{2}{\mathrm{d}}\mu_{h_{\mathrm{ref}}}.\end{split}$
Summing up and recalling (26) and (109) yields
$\begin{split}\partial_{t}\mathcal{E}_{\mathrm{tot},\ell}\geq&-2(1+\varepsilon)t^{-1}E_{(e,\omega),\ell}-3t^{-1}E_{(\varphi),\ell}-2t^{-1}\textstyle{\sum}_{I,J}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}|E_{\mathbf{I}}(k_{IJ})|^{2}{\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ &-4t^{-1}\textstyle{\sum}_{I,J,K}\sum_{|\mathbf{I}|\leq\ell}\int_{\Sigma}|E_{\mathbf{I}}(\gamma_{IJK})|^{2}{\mathrm{d}}\mu_{h_{\mathrm{ref}}}\\ &-7t^{-1}\textstyle{\sum}_{|\mathbf{I}|\leq\ell}\sum_{I}\int_{\Sigma}|E_{\mathbf{I}}a_{I}|^{2}{\mathrm{d}}\mu_{h_{\mathrm{ref}}}+\mathcal{R}.\end{split}$
Since $3\varepsilon\leq\sigma_{V}$ and $\sigma_{V}<1$, we conclude that (108) holds. ∎
Due to the above estimates, we are in a position to deduce energy estimates for all orders.
Corollary 36.
Consider a solution to (11)–(19) as described in Subsection 1.4. Assume, in addition, that $V$ is a $\sigma_{V}$-admissible potential and that the basic supremum assumptions, see Definition 17, are satisfied with $3\varepsilon\leq\sigma_{V}$. Then, for each $\ell\in\mathbb{N}$, there is a $C_{\ell}$ such that, for $t\leq t_{0}$,
(110) $E_{\mathrm{tot},\ell}(t)\leq C_{\ell}t^{-9}.$
Proof.
We prove a slightly improved version of (110) by induction. The inductive hypothesis is that there are constants $C_{\ell}$ and $b_{\ell}$ such that
(111) $\mathcal{E}_{\mathrm{tot},\ell}\leq C_{\ell}\langle\ln t\rangle^{b_{\ell}}t^{-8}$
for $t\leq t_{0}$. That (111) holds for $\ell\leq 1$ is an immediate consequence of (25) and (38).
Given that (111) holds for some $\ell$, consider (108) with $\ell$ replaced by $\ell+1$. Let $C_{\ell+1}$ be the constant appearing in the second term on the right hand side of (108) (with $\ell$ replaced by $\ell+1$) and let
(112) $r_{\ell+1}(t):=8\ln t+\varepsilon^{-1}C_{\ell+1}t^{\varepsilon}.$
Compute
$\partial_{t}\Big(\exp(r_{\ell+1})\mathcal{E}_{\mathrm{tot},\ell+1}\Big)\geq-C_{\ell+1}t^{-1}\exp(r_{\ell+1})\mathcal{E}_{\mathrm{tot},\ell}^{1/2}\mathcal{E}_{\mathrm{tot},\ell+1}^{1/2}.$
Introducing
(113) $F_{\ell}:=\exp(r_{\ell+1})\mathcal{E}_{\mathrm{tot},\ell+1}$
and noting that, by the inductive hypothesis,
$\exp(r_{\ell+1}(t)/2)\mathcal{E}_{\mathrm{tot},\ell}^{1/2}(t)\leq C_{\ell+1}\langle\ln t\rangle^{b_{\ell}/2},$
we conclude that
$\partial_{t}F_{\ell}^{1/2}\geq-C_{\ell+1}t^{-1}\langle\ln t\rangle^{b_{\ell}/2}.$
Integrating this estimate from $t$ to $t_{\ell+1}$ yields
$F_{\ell}^{1/2}(t_{\ell+1})-F_{\ell}^{1/2}(t)\geq-C_{\ell+1}\langle\ln t\rangle^{b_{\ell}/2+1}.$
Since the solution is smooth, $F_{\ell}^{1/2}(t_{\ell+1})$ is bounded by some constant, so that
$F_{\ell}^{1/2}(t)\leq C_{\ell+1}\langle\ln t\rangle^{b_{\ell}/2+1}.$
for $t\leq t_{\ell+1}$. Recalling (112) and (113), we conclude that
$\mathcal{E}_{\mathrm{tot},\ell+1}(t)\leq C_{\ell+1}\langle\ln t\rangle^{b_{\ell}+2}t^{-8}$
for $t\leq t_{\ell+1}$. Since this estimate also holds for $t_{\ell+1}\leq t\leq t_{0}$, due to the fact that the solution is smooth, we conclude that (111) holds with $\ell$ replaced by $\ell+1$. By induction, we conclude that it holds for all $\ell$. Since $t\langle\ln t\rangle^{b_{\ell}}$ is bounded for $t\leq t_{0}$, the estimate (110) holds. ∎
Finally, we are in a position to prove Theorem 19.
Proof of Theorem 19.
Due to (140), (25a) and (110), it follows that
$\|e\|_{H^{\ell}(\Sigma_{t})}\leq C_{\ell,k}t^{(-1+3\varepsilon)(1-\ell/k)}t^{-9\ell/(2k)}.$
Taking a large enough $k$, depending only on $\ell$ and $\varepsilon$, it follows that
$\|e\|_{H^{\ell}(\Sigma_{t})}\leq C_{\ell,\varepsilon}t^{-1+2\varepsilon}.$
Due to Sobolev embedding, the same holds for all $C^{\ell}$ norms. Due to this and similar arguments for $\omega$, $\gamma$ and $e_{I}\varphi$, we deduce from (25a) and (110) that (27a) holds. Similarly, using (25a) and (25c), we can argue that (27c) holds. Next, due to (140), (25b) and (110), it follows that
$t^{1+\varepsilon/4}\|k\|_{C^{\ell}(\Sigma_{t})}+t^{1+\varepsilon/4}\|e_{0}\varphi\|_{C^{\ell}(\Sigma_{t})}\leq C_{\ell}$
for all $\ell\in\mathbb{N}_{0}$ and all $t\leq t_{0}$. Moreover, due to (140), (38) and (110), it follows that
$t^{\varepsilon}\|\varphi\|_{C^{\ell}(\Sigma_{t})}\leq C_{\ell}$
for all $\ell\in\mathbb{N}_{0}$ and all $t\leq t_{0}$. Combining this observation with (44) yields
$t^{2}\|V\circ\varphi\|_{C^{\ell}(\Sigma_{t})}+t^{2}\|V^{\prime}\circ\varphi\|_{C^{\ell}(\Sigma_{t})}\leq C_{\ell}t^{4\varepsilon}$
for all $\ell\in\mathbb{N}_{0}$ and all $t\leq t_{0}$. On the other hand, due to (76),
$\begin{split}\partial_{t}(tk_{IJ})=&-(N-1)k_{IJ}+te_{(I}e_{J)}(N)-te_{K}(N\gamma_{K(IJ)})-Nte_{(I}(\gamma_{J)KK})+Nt\gamma_{KLL}\gamma_{K(IJ)}\\ &+Nt\gamma_{I(KL)}\gamma_{J(KL)}-\tfrac{1}{4}Nt\gamma_{KLI}\gamma_{KLJ}+Nte_{I}(\varphi)e_{J}(\varphi)+\tfrac{2N}{n-1}t(V\circ\varphi)\delta_{IJ}.\end{split}$
Combining this equality with the estimates derived so far yields the conclusion that
$\|\partial_{t}(tk_{IJ})\|_{C^{\ell}(\Sigma_{t})}\leq C_{\ell}t^{-1+\varepsilon/4}$
for all $\ell\in\mathbb{N}_{0}$ and all $t\leq t_{0}$. This estimate can be integrated to yield the existence of a $\mathring{\mathfrak{K}}_{IJ}\in C^{\infty}(\Sigma)$ such that (28a) holds. Next, note that (16) implies that
$\partial_{t}\left(te_{0}(\varphi)\right)=-(N-1)e_{0}(\varphi)+tNe_{I}\left(e_{I}(\varphi)\right)+te_{I}(N)e_{I}(\varphi)-tN\gamma_{JII}e_{J}(\varphi)-tNV^{\prime}\circ\varphi.$
By an argument similar to the derivation of (28a), we conclude that there is a $\mathring{\Psi}\in C^{\infty}(\Sigma)$ such that
(114) $\|t(e_{0}\varphi)(\cdot,t)-\mathring{\Psi}\|_{C^{\ell}(\Sigma)}+\|(t\partial_{t}\varphi)(\cdot,t)-\mathring{\Psi}\|_{C^{\ell}(\Sigma)}\leq C_{\ell}t^{\varepsilon/4}$
for all $\ell\in\mathbb{N}_{0}$ and all $t\leq t_{0}$; the bound on the second term on the left hand side follows from the bound of the first term, combined with (27c). The bound on the second term can be integrated to conclude that there is a $\mathring{\Phi}\in C^{\infty}(\Sigma)$ such that
(115) $\|\varphi(\cdot,t)-\mathring{\Psi}\ln t-\mathring{\Phi}\|_{C^{\ell}(\Sigma)}\leq C_{\ell}t^{\varepsilon/4}$
for all $\ell\in\mathbb{N}_{0}$ and all $t\leq t_{0}$. Combining (114) and (115) yields (28b). Finally, due to (28), it is clear that (27b) holds. ∎
3. Asymptotics
The first step in the derivation of the asymptotics is to deduce basic consequences of the assumptions of Theorem 24.
3.1. Basic consequences of the assumptions
We begin by deriving bounds for $\mathring{p}_{I}$.
Lemma 37.
Let $3\leq n\in\mathbb{N}$, $k_{0}\in\mathbb{N}$, $r>0$, $\delta\in(0,1)$ and fix a $\delta$-admissible potential $V\in C^{\infty}(\mathbb{R})$. If $(M,g,\varphi)$ is a solution to the Einstein–nonlinear scalar field equations satisfying the $(\delta,n,k_{0},r)$-assumptions, then
$|\mathring{p}_{I}|<1-2\delta$
on $\Sigma$ for all $I$.
Remark 38.
Here and below, $\Sigma$ is assumed to be the closed manifold appearing in Definition 23.
Proof.
The conclusion is an immediate consequence of the assumptions and the argument given in the proof of [33, Lemma 77, p. 29]. ∎
Lemma 39.
Let $\alpha_{I}$, $I=1,\dots,n$, and $\beta$ be constants. Let $3\leq n\in\mathbb{N}$, $k_{0}\in\mathbb{N}$, $r>0$, $\delta\in(0,1)$ and fix a $\delta$-admissible potential $V\in C^{\infty}(\mathbb{R})$. If $(M,g,\varphi)$ is a solution to the Einstein–nonlinear scalar field equations satisfying the $(\delta,n,k_{0},r)$-assumptions, then, using the notation of Subsection A.4, there is a constant $C$ such that
(116) $\big|E_{\mathrm{\mathbf{I}}}\big(\theta^{\sum_{I}\alpha_{I}p_{I}+\beta}\big)\big|\leq C\langle\ln t\rangle^{|\mathrm{\mathbf{I}}|}t^{-\sum_{I}\alpha_{I}\mathring{p}_{I}-\beta},\qquad\|\ln\theta\|_{C^{k_{0}}(\Sigma)}\leq C\langle\ln t\rangle$
on $M$ for $|\mathrm{\mathbf{I}}|\leq k_{0}$; and for $t\in(0,T)$, respectively.
Remark 40.
Here and below, $T>0$ is assumed to be the positive number appearing in Definition 23.
Proof.
We begin with the case without derivatives. Write
$\theta^{\sum_{I}\alpha_{I}p_{I}+\beta}=N^{-\sum_{I}\alpha_{I}p_{I}-\beta}(t\Theta)^{\sum_{I}\alpha_{I}p_{I}+\beta}t^{\sum_{I}\alpha_{I}(\mathring{p}_{I}-p_{I})}t^{-\sum_{I}\alpha_{I}\mathring{p}_{I}-\beta}.$
Note that $|(\ln t)(\mathring{p}_{I}-p_{I})|\leq C|\ln t|t^{\delta}$ is bounded for $t\in(0,T)$. Moreover, due to (36d), it is clear that $(t\Theta)^{-1}\leq 2$. Due to this observation and the assumptions, it is clear that $N$, $N^{-1}$, $t\Theta$ and $(t\Theta)^{-1}$ are bounded. Combining the above observations, it is clear that
$|\theta^{\sum_{I}\alpha_{I}p_{I}+\beta}|\leq Ct^{-\sum_{I}\alpha_{I}\mathring{p}_{I}-\beta}.$
Next,
$|\ln\theta|=|\ln(t\Theta)-\ln N-\ln t|\leq C\langle\ln t\rangle.$
Turning to the derivatives, if $1\leq|\mathrm{\mathbf{I}}|\leq k_{0}$,
$E_{\mathrm{\mathbf{I}}}\ln\theta=E_{\mathrm{\mathbf{I}}}(\ln(t\Theta)-\ln N-\ln t)=E_{\mathrm{\mathbf{I}}}(\ln(t\Theta)-\ln N).$
Thus $|E_{\mathrm{\mathbf{I}}}\ln\theta|\leq C$ and the second estimate in (116) follows. Finally,
$E_{\mathrm{\mathbf{I}}}\big(\theta^{\sum_{I}\alpha_{I}p_{I}+\beta}\big)=\theta^{\sum_{I}\alpha_{I}p_{I}+\beta}\textstyle\sum E_{\mathrm{\mathbf{I}}_{1}}\big((\ln\theta)(\textstyle\sum_{I}\alpha_{I}p_{I}+\beta)\big)\cdots E_{\mathrm{\mathbf{I}}_{r}}\big((\ln\theta)(\textstyle\sum_{I}\alpha_{I}p_{I}+\beta)\big),$
from which the result follows. ∎
Lemma 41.
Let $3\leq n\in\mathbb{N}$, $k_{0}\in\mathbb{N}$, $r>0$, $\delta\in(0,1)$ and fix a $\delta$-admissible potential $V\in C^{\infty}(\mathbb{R})$. If $(M,g,\varphi)$ is a solution to the Einstein–nonlinear scalar field equations satisfying the $(\delta,n,k_{0},r)$-assumptions, then, using the notation of Subsection A.4, there is a constant $C$ such that
(117) $\|\varphi\|_{C^{k_{0}}(\Sigma)}\leq C\langle\ln t\rangle,\qquad\|V\circ\varphi\|_{C^{k_{0}}(\Sigma)}\leq C\langle\ln t\rangle^{k_{0}}t^{-2+2\delta}$
for all $t\in(0,T)$.
Proof.
The first estimate in (117) is an immediate consequence of the assumptions, the second estimate in (116) and
$\varphi=-\mathring{\Psi}\ln\theta+(\mathring{\Psi}-\Psi)\ln\theta+\Phi.$
Next, since
$|\mathring{\Psi}|=\big(1-\textstyle\sum_{I}\mathring{p}_{I}^{2}\big)^{1/2}\leq 1$
and $(\mathring{\Psi}-\Psi)\ln\theta+\Phi$ is bounded,
$|V^{(k)}\circ\varphi|\leq Ce^{2(1-\delta)|\varphi|}\leq Ce^{2(1-\delta)|\mathring{\Psi}\ln\theta|}\leq Ct^{-2+2\delta}$
for $k\leq k_{0}$, where we appealed to $\ln\theta=\ln(t\Theta)-\ln N-\ln t$. Next,
$|E_{\mathrm{\mathbf{I}}}(V\circ\varphi)|=\big|\textstyle\sum(V^{(k)}\circ\varphi)E_{\mathrm{\mathbf{I}}_{1}}(\varphi)\cdots E_{\mathrm{\mathbf{I}}_{k}}(\varphi)\big|\leq C\langle\ln t\rangle^{|\mathrm{\mathbf{I}}|}t^{-2+2\delta}$
for $|\mathrm{\mathbf{I}}|\leq k_{0}$. The lemma follows. ∎
The following lemma is necessary to obtain bounds that are uniform on $\Sigma$ in Theorem 56 below.
Lemma 42.
Let $3\leq n\in\mathbb{N}$, $k_{0}\in\mathbb{N}$, $r>0$, $\delta\in(0,1)$ and fix a $\delta$-admissible potential $V\in C^{\infty}(\mathbb{R})$. If $(M,g,\varphi)$ is a solution to the Einstein–nonlinear scalar field equations satisfying the $(\delta,n,k_{0},r)$-assumptions, then, using the notation of Subsection A.4, there is a finite open cover $\{U_{j}\}$ of $\Sigma$, and constants $0<\alpha_{j}\leq\delta/4$ such that
(118a) $|-1+\mathring{p}_{I}(x)+\mathring{p}_{J}(x)-\mathring{p}_{K}(x)+m\delta+\alpha_{j}|$ $\geq\tfrac{\alpha_{j}}{2},$
(118b) $|-1-\mathring{p}_{I}(x)+m\delta+\alpha_{j}|$ $\geq\tfrac{\alpha_{j}}{2}$
for all $x\in U_{j}$, $m\in\mathbb{Z}$ and $I,J,K\in\{1,\dots,n\}$ with $I\neq J$.
Proof.
Let $a_{1},\ldots,a_{k}$ be real numbers and let $m_{i}\delta$ be the integer multiple of $\delta$ that is closest to $a_{i}$ (if there are two such numbers, either of them may be chosen). By rearranging the $a_{i}$, we may assume that the first $K$ of them are the ones satisfying $a_{i}=m_{i}\delta$. Then there is an $0<\alpha\leq\delta/4$ such that
$|a_{i}-m_{i}\delta-\alpha|\geq\alpha$
for all $i$. Indeed, we can take $\alpha$ to be half the minimum of $|a_{i}-m_{i}\delta|$ for $i\geq K+1$. In fact, the inequality
$|a_{i}-m\delta-\alpha|\geq\alpha$
holds for every integer $m$.
For every $x\in\Sigma$, we think of the $a_{i}$ as the numbers $1-\mathring{p}_{I}(x)-\mathring{p}_{J}(x)+\mathring{p}_{K}(x)$ for $I\neq J$, and the $1+\mathring{p}_{I}(x)$. Then by the previous paragraph, there is an $0<\alpha(x)\leq\delta/4$ such that
$\begin{split}|1-\mathring{p}_{I}(x)-\mathring{p}_{J}(x)+\mathring{p}_{K}(x)-m\delta-\alpha(x)|&\geq\alpha(x),\\ |1+\mathring{p}_{I}(x)-m\delta-\alpha(x)|&\geq\alpha(x).\end{split}$
By continuity, there is an open set $U_{x}\subset\Sigma$ containing $x$, such that
$\begin{split}|1-\mathring{p}_{I}(y)-\mathring{p}_{J}(y)+\mathring{p}_{K}(y)-m\delta-\alpha(x)|&\geq\tfrac{\alpha(x)}{2},\\ |1+\mathring{p}_{I}(y)-m\delta-\alpha(x)|&\geq\tfrac{\alpha(x)}{2},\end{split}$
for all $y\in U_{x}$. Finally, since the $U_{x}$ form an open cover of $\Sigma$, there are finitely many of them that cover $\Sigma$. ∎
3.2. The eigenframe
Due to the assumptions, the eigenvalues $p_{I}$ of $\mathcal{K}$ are distinct everywhere. Due to [42, Lemma A.1, p. 223], we can, by taking a finite covering space of $\Sigma$, if necessary, assume that there are smooth vector fields $\widehat{e}_{I}$ on $M$, for $I=1,\ldots,n$, tangential to the hypersurfaces $\Sigma_{t}$ and such that $\mathcal{K}(\widehat{e}_{I})=p_{I}\widehat{e}_{I}$ and $h(\widehat{e}_{I},\widehat{e}_{J})=\delta_{IJ}$. In particular, $\{\widehat{e}_{I}\}_{I=1}^{n}$ is then an orthonormal frame for $h$ on each $\Sigma_{t}$. Denote the corresponding dual frame $\{\widehat{\omega}^{I}\}_{I=1}^{n}$. Next, we deduce evolution equations for the $\widehat{e}_{I}$.
Lemma 43.
Let $3\leq n\in\mathbb{N}$, $k_{0}\in\mathbb{N}$, $r>0$, $\delta\in(0,1)$ and fix a $\delta$-admissible potential $V\in C^{\infty}(\mathbb{R})$. Given that $(M,g,\varphi)$ is a solution to the Einstein–nonlinear scalar field equations satisfying the $(\delta,n,k_{0},r)$-assumptions, assume that there is an eigenframe $\{\widehat{e}_{I}\}_{I=1}^{n}$ with the properties listed above. Then $\{\widehat{e}_{I}\}_{I=1}^{n}$ and the eigenvalues $p_{I}$ satisfy the system of equations
(119a) $[e_{0},\widehat{e}_{I}]$ $=\widehat{e}_{I}(\ln N)e_{0}-\theta p_{I}\widehat{e}_{I}+\textstyle\sum_{J\neq I}\tfrac{1}{\theta(p_{I}-p_{J})}h\big((\mathcal{L}_{e_{0}}K)(\widehat{e}_{I}),\widehat{e}_{J}\big)\widehat{e}_{J},$
(119b) $e_{0}p_{I}$ $=-\theta^{-1}e_{0}(\theta)p_{I}+\theta^{-1}g\big((\mathcal{L}_{e_{0}}K)(\widehat{e}_{I}),\widehat{e}_{I}\big)$
(no summation on $I$). In particular,
$\partial_{t}\widehat{e}_{I}^{\;i}=-\Theta p_{I}\widehat{e}_{I}^{\;i}+\textstyle\sum_{J\neq I}\tfrac{N^{2}}{\Theta(p_{I}-p_{J})}g\big((\mathcal{L}_{e_{0}}K)(\widehat{e}_{I}),\widehat{e}_{J}\big)\widehat{e}_{J}^{\;i}$
(no summation on $I$), where $\widehat{e}_{I}=\widehat{e}_{I}^{\;i}E_{i}$.
Remark 44.
Recall from Subsection A.3, that with respect to the frame $\{E_{i}\}_{i=1}^{n}$ and dual frame $\{\eta^{i}\}_{i=1}^{n}$,
$\mathcal{L}_{e_{0}}K=N^{-1}\partial_{t}(K^{i}_{\phantom{i}j})E_{i}\otimes\eta^{j}.$
Proof.
Compute
$0=e_{0}\big(g(\widehat{e}_{I},\widehat{e}_{J})\big)=(\mathcal{L}_{e_{0}}g)(\widehat{e}_{I},\widehat{e}_{J})+g\big([e_{0},\widehat{e}_{I}],\widehat{e}_{J}\big)+g\big(\widehat{e}_{I},[e_{0},\widehat{e}_{J}]\big).$
Hence
$2\theta p_{I}\delta_{IJ}=-g\big([e_{0},\widehat{e}_{I}],\widehat{e}_{J}\big)-g\big(\widehat{e}_{I},[e_{0},\widehat{e}_{J}]\big)$
(no summation on $I$ on the left hand side). In particular, it follows that
$g\big([e_{0},\widehat{e}_{I}],\widehat{e}_{I}\big)=-\theta p_{I}$
(no summation on $I$ on the left hand side). For $X\in\mathfrak{X}(M)$, denote by $\overline{X}$ the part of $X$ tangential to the $\Sigma_{t}$. Then, applying $\overline{[e_{0},\,\cdot\,]}$ to both sides of $K(\widehat{e}_{I})=\theta p_{I}\widehat{e}_{I}$ (no summation) yields
$(\mathcal{L}_{e_{0}}K)(\widehat{e}_{I})+K\big(\overline{[e_{0},\widehat{e}_{I}]}\big)=e_{0}(\theta p_{I})\widehat{e}_{I}+\theta p_{I}\overline{[e_{0},\widehat{e}_{I}]}$
(no summation). Note that $K\big(\overline{[e_{0},\widehat{e}_{I}]}\big)=\sum_{J}g\big([e_{0},\widehat{e}_{I}],\widehat{e}_{J}\big)\theta p_{J}\widehat{e}_{J}$. The equation for $p_{I}$ follows by applying $g(\widehat{e}_{I},\,\cdot\,)$ to the resulting equation. On the other hand, applying $g(\widehat{e}_{J},\,\cdot\,)$ with $J\neq I$, yields
$g\big([e_{0},\widehat{e}_{I}],\widehat{e}_{J}\big)=\tfrac{1}{\theta(p_{I}-p_{J})}g\big((\mathcal{L}_{e_{0}}K)(\widehat{e}_{I}),\widehat{e}_{J}\big).$
Next,
$g\big([e_{0},\widehat{e}_{I}],e_{0}\big)=g\big([\widehat{e}_{I},N^{-1}\partial_{t}],e_{0}\big)=g\big(-N^{-2}\widehat{e}_{I}(N)\partial_{t},e_{0}\big)=g\big(-\widehat{e}_{I}(\ln N)e_{0},e_{0}\big)=\widehat{e}_{I}(\ln N).$
Finally, we can write
$e_{0},\widehat{e}_{I}]=-g\big([e_{0},\widehat{e}_{I}],e_{0}\big)e_{0}+\textstyle\sum_{J}g\big([e_{0},\widehat{e}_{I}],\widehat{e}_{J}\big)\widehat{e}_{J}.$
The lemma follows. ∎
The idea is to translate estimates for the Fermi-Walker frame $\{e_{I}\}_{I=1}^{n}$, its dual frame $\{\omega^{I}\}_{I=1}^{n}$ and its structure coefficients, to the eigenframe $\{\widehat{e}_{I}\}_{I=1}^{n}$. To this end, we consider the family of matrices $A$ whose elements are defined by
$\widehat{e}_{I}=A_{I}^{J}e_{J}.$
Note that $A$ is a family of orthogonal matrices, so that $e_{I}=A_{J}^{I}\widehat{e}_{J}$. For the dual frames, $\widehat{\omega}^{I}=A_{I}^{J}\omega^{J}$. Moreover, $A_{I}^{J}=g(\widehat{e}_{I},e_{J})$, so that $A_{I}^{J}$ is smooth, and the bound $|A_{I}^{J}|\leq 1$ holds everywhere on $M$. Therefore, in order to translate estimates for the Fermi-Walker frame, we only need to control the derivatives of $A$. To this end, we deduce an evolution equation for $A$.
Lemma 45.
Let $3\leq n\in\mathbb{N}$, $k_{0}\in\mathbb{N}$, $r>0$, $\delta\in(0,1)$ and fix a $\delta$-admissible potential $V\in C^{\infty}(\mathbb{R})$. Given that $(M,g,\varphi)$ is a solution to the Einstein–nonlinear scalar field equations satisfying the $(\delta,n,k_{0},r)$-assumptions, assume that there is an eigenframe $\{\widehat{e}_{I}\}_{I=1}^{n}$ with the properties listed at the beginning of the present subsection. Then the elements of $A$ satisfy the evolution equation
(120) $\partial_{t}A_{I}^{J}=\textstyle\sum_{K\neq I}\tfrac{N^{2}}{\Theta(p_{I}-p_{K})}g\big((\mathcal{L}_{e_{0}}K)(\widehat{e}_{I}),\widehat{e}_{K}\big)A_{K}^{J}.$
Proof.
First, note that
(121) $\partial_{t}\widehat{e}_{I}^{\;i}=\big(\partial_{t}A_{I}^{J}\big)e_{J}^{i}+A_{I}^{J}\partial_{t}e_{J}^{i}.$
From (32), it follows that
$\partial_{t}e_{I}^{i}=-Nk_{IJ}e_{J}^{i}.$
Therefore,
(122) $\begin{split}A_{I}^{J}\partial_{t}e_{J}^{i}=&-NA_{I}^{J}k_{JK}e_{K}^{i}=-NA_{M}^{K}k(\widehat{e}_{I},\widehat{e}_{M})e_{K}^{i}\\ =&-NA_{I}^{K}\theta p_{I}e_{K}^{i}=-N\theta p_{I}\widehat{e}_{I}^{\;i}\end{split}$
(no summation on $I$). We can then conclude from Lemma 43 that
$\big(\partial_{t}A_{I}^{J}\big)e_{J}=\textstyle\sum_{K\neq I}\tfrac{N^{2}}{\Theta(p_{I}-p_{K})}g\big(\mathcal{L}_{e_{0}}K(\widehat{e}_{I}),\widehat{e}_{K}\big)\widehat{e}_{K}.$
Since $\widehat{e}_{K}=A_{K}^{J}e_{J}$, the result follows. ∎
Since we only have information about the Fermi-Walker frame, we rewrite the evolution equation for $A_{I}^{J}$ as follows. Define $B_{IKLM}$ by
$B_{IKLM}:=\tfrac{N^{2}}{\Theta(p_{I}-p_{K})}\big(-\mathrm{Ric}_{h}(e_{L},e_{M})+N^{-1}(\overline{\nabla}^{2}N)(e_{L},e_{M})+e_{L}(\varphi)e_{M}(\varphi)\big),$
for $I\neq K$, and $B_{IILM}=0$. Then, (30a) implies that $A_{I}^{J}$ satisfies the equation
(123) $\partial_{t}A_{I}^{J}=\textstyle\sum_{K}B_{IKLM}A_{I}^{L}A_{K}^{M}A_{K}^{J}.$
It is thus of interest to deduce estimates for $B_{IKLM}$.
Lemma 46.
Let $3\leq n\in\mathbb{N}$, $k_{0}\in\mathbb{N}$, $r>0$, $\delta\in(0,1)$ and fix a $\delta$-admissible potential $V\in C^{\infty}(\mathbb{R})$. Given that $(M,g,\varphi)$ is a solution to the Einstein–nonlinear scalar field equations satisfying the $(\delta,n,k_{0},r)$-assumptions, assume that there is an eigenframe $\{\widehat{e}_{I}\}_{I=1}^{n}$ with the properties listed at the beginning of the present subsection. Then there is a constant $C$ such that
(124) $\|\mathrm{Ric}_{h}(e_{I},e_{J})\|_{C^{k_{0}-1}(\Sigma)}\leq Ct^{-2+6\delta},\qquad\|N^{-1}(\overline{\nabla}^{2}N)(e_{I},e_{J})\|_{C^{k_{0}-1}(\Sigma)}\leq Ct^{-2+2\delta}$
for $t<T$. Moreover, for $t<T$,
$\begin{split}\|B_{IKLM}\|_{C^{k_{0}-1}(\Sigma)}\leq Ct^{-1+2\delta}.\end{split}$
Proof.
We begin with $\mathrm{Ric}_{h}(e_{I},e_{J})$. Due to Lemma 16,
$\mathrm{Ric}_{h}(e_{I},e_{J})=e_{K}\gamma_{K(IJ)}+e_{(I}\gamma_{J)KK}-\gamma_{I(KL)}\gamma_{J(KL)}+\tfrac{1}{4}\gamma_{KLI}\gamma_{KLJ}-\gamma_{KLL}\gamma_{K(IJ)}.$
Thus the first estimate in (124) follows from (36). Next, note that
$N^{-1}(\overline{\nabla}^{2}N)(e_{I},e_{J})=e_{I}e_{J}(\ln N)+e_{I}(\ln N)e_{J}(\ln N)-\Gamma_{IJK}e_{K}(\ln N),$
where $\Gamma_{IJK}=g(\nabla_{e_{I}}e_{J},e_{K})$ are the connection coefficients of the Fermi-Walker frame. Since $\Gamma_{IJK}=\frac{1}{2}(\gamma_{IJK}+\gamma_{KIJ}-\gamma_{JKI})$ due to the Koszul formula, we conclude that the second estimate in (124) holds. The estimate for $B_{IKLM}$ now follows by its definition, and the estimates for the $p_{I}$, $N$, $\Theta$ and $e_{I}(\varphi)$, including the estimate (36d). ∎
Proposition 47.
[…]
Remark 48.
When we speak of $t_{0}$ in what follows, we take it for granted to be the number fixed in this proposition.
Proof.
For $m<k_{0}$, define the functions
$\rho_{m}:=\textstyle\sum_{I,J}\textstyle\sum_{|\mathrm{\mathbf{I}}|=m}\big(E_{\mathrm{\mathbf{I}}}A_{I}^{J}\big)^{2}.$
Since $|A_{I}^{J}|\leq 1$ everywhere, it follows that $\rho_{0}\leq n^{2}$. We derive uniform bounds for the remaining $\rho_{m}$ inductively. Fix $m<k_{0}$. Suppose that there is a constant $C$, such that $\rho_{k}\leq C$ for all $k\leq m-1$. Then, by Lemma 46 and (123),
$\begin{split}|\partial_{t}\rho_{m}|&=2\textstyle\sum_{I,J}\textstyle\sum_{|\mathrm{\mathbf{I}}|=m}\big(E_{\mathrm{\mathbf{I}}}A_{I}^{J}\big)\big(E_{\mathrm{\mathbf{I}}}\partial_{t}A_{I}^{J}\big)\leq 2\rho_{m}^{1/2}\big(\textstyle\sum_{I,J}\textstyle\sum_{|\mathrm{\mathbf{I}}|=m}\big(E_{\mathrm{\mathbf{I}}}\partial_{t}A_{I}^{J}\big)^{2}\big)^{1/2}\\ &\leq 2\rho_{m}^{1/2}\big(Ct^{-1+2\delta}\rho_{m}^{1/2}+Ct^{-1+2\delta}\big)\leq Ct^{-1+2\delta}\rho_{m}+Ct^{-1+2\delta}\rho_{m}^{1/2}\\ &\leq Ct^{-1+2\delta}\rho_{m}+Ct^{-1+2\delta},\end{split}$
where we used Young’s inequality. Since $t^{-1+2\delta}$ is integrable on $(0,t_{0}]$, it follows that
$\rho_{m}(s)=\rho_{m}(t_{0})+\textstyle\int_{s}^{t_{0}}-\partial_{t}\rho_{m}(t){\mathrm{d}}t\leq C+C\textstyle\int_{s}^{t_{0}}t^{-1+2\delta}\rho_{m}(t){\mathrm{d}}t$
for all $s\in(0,t_{0}]$. By Grönwall’s lemma, we conclude that $\rho_{m}(t)\leq C$ for some constant $C$ and all $t\in(0,t_{0}]$.
[…]
$\big|E_{\mathrm{\mathbf{I}}}A_{I}^{J}(s)-E_{\mathrm{\mathbf{I}}}A_{I}^{J}(u)\big|\leq\textstyle\int_{u}^{s}\big|E_{\mathrm{\mathbf{I}}}\partial_{t}A_{I}^{J}(t)\big|{\mathrm{d}}t\leq C(s^{2\delta}-u^{2\delta}).$
[…]
Now that we know $A_{I}^{J}$ to be bounded in $C^{k_{0}-1}(\Sigma)$, we can obtain estimates in $C^{k_{0}-1}(\Sigma)$ for the $\widehat{e}_{I}$, the $\widehat{\omega}^{I}$, and the corresponding structure coefficients.
Lemma 49.
Let $3\leq n\in\mathbb{N}$, $k_{0}\in\mathbb{N}$, $r>0$, $\delta\in(0,1)$ and fix a $\delta$-admissible potential $V\in C^{\infty}(\mathbb{R})$. Given that $(M,g,\varphi)$ is a solution to the Einstein–nonlinear scalar field equations satisfying the $(\delta,n,k_{0},r)$-assumptions, assume that there is an eigenframe $\{\widehat{e}_{I}\}_{I=1}^{n}$ with the properties listed at the beginning of the present subsection. Let $\lambda_{IJK}:=g\big([\widehat{e}_{I},\widehat{e}_{J}],\widehat{e}_{K}\big)$. Then there is a constant $C$ such that, for all $t\in(0,t_{0}]$,
$\|\widehat{e}_{I}^{\;i}\|_{C^{k_{0}-1}(\Sigma)}+\|\widehat{\omega}_{i}^{I}\|_{C^{k_{0}-1}(\Sigma)}+\|\lambda_{IJK}\|_{C^{k_{0}-2}(\Sigma)}\leq Ct^{-1+3\delta}.$
Proof.
The estimates for $\widehat{e}_{I}$ and $\widehat{\omega}^{I}$ follow immediately from Proposition 47, since $\widehat{e}_{I}=A_{I}^{J}e_{J}$ and $\widehat{\omega}^{I}=A_{I}^{J}\omega^{J}$. For $\lambda_{IJK}$,
$\widehat{e}_{I},\widehat{e}_{J}]=\big[A_{I}^{L}e_{L},A_{J}^{M}e_{M}\big]=A_{I}^{L}e_{L}\big(A_{J}^{M}\big)e_{M}-A_{J}^{M}e_{M}\big(A_{I}^{L}\big)e_{L}+A_{J}^{M}A_{I}^{L}[e_{L},e_{M}].$
Hence,
$\lambda_{IJK}=A_{I}^{L}e_{L}\big(A_{J}^{M}\big)A_{K}^{M}-A_{J}^{M}e_{M}\big(A_{I}^{L}\big)A_{K}^{L}+A_{J}^{M}A_{I}^{L}A_{K}^{N}\gamma_{LMN}.$
The estimates now follow from Proposition 47. ∎
3.3. The expansion normalized eigenframe
We introduce an expansion normalized eigenframe $\{{\check{e}}_{I}\}_{I=1}^{n}$, with dual frame $\{{\check{\omega}}^{I}\}_{I=1}^{n}$, defined by
${\check{e}}_{I}:=\theta^{-p_{I}}\widehat{e}_{I},\qquad{\check{\omega}}^{I}:=\theta^{p_{I}}\widehat{\omega}^{I}$
(no summation). Denote by $\Lambda_{IJ}^{K}$ the corresponding structure coefficients, given by
$\Lambda_{IJ}^{K}:={\check{\omega}}^{K}\big([{\check{e}}_{I},{\check{e}}_{J}]\big).$
Lemma 50.
Let $3\leq n\in\mathbb{N}$, $k_{0}\in\mathbb{N}$, $r>0$, $\delta\in(0,1)$ and fix a $\delta$-admissible potential $V\in C^{\infty}(\mathbb{R})$. Given that $(M,g,\varphi)$ is a solution to the Einstein–nonlinear scalar field equations satisfying the $(\delta,n,k_{0},r)$-assumptions, assume that there is an eigenframe $\{\widehat{e}_{I}\}_{I=1}^{n}$ with the properties listed at the beginning of Subsection 3.2. Then there is a constant $C$ such that
$\big|E_{\mathrm{\mathbf{I}}}\big({\check{e}}_{I}^{i}\big)\big|\leq Ct^{-1+\mathring{p}_{I}+2\delta},\quad\big|E_{\mathrm{\mathbf{I}}}\big({\check{\omega}}_{i}^{I}\big)\big|\leq Ct^{-1-\mathring{p}_{I}+2\delta},\quad\big|E_{\mathrm{\mathbf{J}}}\big(\Lambda_{IJ}^{K}\big)\big|\leq Ct^{-1+\mathring{p}_{I}+\mathring{p}_{J}-\mathring{p}_{K}+2\delta}$
hold for $t\leq t_{0}$, $|\mathrm{\mathbf{I}}|\leq k_{0}-1$ and $|\mathrm{\mathbf{J}}|\leq k_{0}-2$.
Proof.
The estimates for ${\check{e}}_{I}^{i}$ and ${\check{\omega}}_{i}^{I}$ follow directly from the definitions, Lemma 39 and Lemma 49, since $\langle\ln t\rangle^{k_{0}-1}t^{\delta}$ is bounded for $t\in(0,t_{0}]$. For $\Lambda_{IJ}^{K}$,
$\begin{split}[{\check{e}}_{I},{\check{e}}_{J}]&=[\theta^{-p_{I}}\widehat{e}_{I},\theta^{-p_{J}}\widehat{e}_{J}]\\ &=-\theta^{-p_{I}}\widehat{e}_{I}(p_{J}\ln\theta){\check{e}}_{J}+\theta^{-p_{J}}\widehat{e}_{J}(p_{I}\ln\theta){\check{e}}_{I}+\textstyle\sum_{K}\theta^{p_{K}-p_{I}-p_{J}}\lambda_{IJK}{\check{e}}_{K}.\end{split}$
Hence,
$\Lambda_{IJ}^{K}=\theta^{p_{K}-p_{I}-p_{J}}\big(-\widehat{e}_{I}(p_{J}\ln\theta)\delta_{J}^{K}+\widehat{e}_{J}(p_{I}\ln\theta)\delta_{I}^{K}+\lambda_{IJK}\big).$
The last estimate now follows from Lemmas 39 and 49. ∎
Next, we deduce evolution equations for these objects.
Lemma 51.
Let $3\leq n\in\mathbb{N}$, $k_{0}\in\mathbb{N}$, $r>0$, $\delta\in(0,1)$ and fix a $\delta$-admissible potential $V\in C^{\infty}(\mathbb{R})$. Given that $(M,g,\varphi)$ is a solution to the Einstein–nonlinear scalar field equations satisfying the $(\delta,n,k_{0},r)$-assumptions, assume that there is an eigenframe $\{\widehat{e}_{I}\}_{I=1}^{n}$ with the properties listed at the beginning of Subsection 3.2. Then ${\check{e}}_{I}$, ${\check{\omega}}^{I}$ and $\Lambda_{IJ}^{K}$ satisfy the following system of equations
(126a) $\partial_{t}{\check{e}}_{I}^{i}$ $=\mathcal{B}_{I}^{J}{\check{e}}_{J}^{i},$
(126b) $\partial_{t}{\check{\omega}}_{i}^{I}$ $=-\mathcal{B}_{J}^{I}{\check{\omega}}_{i}^{J},$
(126c) $\partial_{t}\Lambda_{IJ}^{K}$ $=-\mathcal{B}_{K}^{K}\Lambda_{IJ}^{K}+\mathcal{B}_{I}^{L}\Lambda_{LJ}^{K}-\mathcal{B}_{J}^{L}\Lambda_{LI}^{K}+{\check{e}}_{I}\big(\mathcal{B}_{J}^{K}\big)-{\check{e}}_{J}\big(\mathcal{B}_{I}^{K}\big)+\textstyle\sum_{L\neq K}\theta^{2(p_{K}-p_{L})}\mathcal{B}_{K}^{L}\Lambda_{IJ}^{L},$
where $\mathcal{B}_{I}^{J}$ is given by
$\mathcal{B}_{I}^{I}:=-(\ln\theta)\partial_{t}p_{I}-\Theta^{-1}N^{2}p_{I}\big(-\mathrm{Scal}_{h}+N^{-1}\Delta_{h}N+|{\mathrm{d}}\varphi|_{h}^{2}+\tfrac{2n}{n-1}V\circ\varphi\big),$
(no summation) and
$\mathcal{B}_{I}^{J}:=\tfrac{N^{2}}{\Theta(p_{I}-p_{J})}(\mathcal{L}_{e_{0}}K)\big({\check{e}}_{I},{\check{\omega}}^{J}\big),$
for $I\neq J$.
Proof.
Due to (120), (121) and (122),
$\begin{split}\partial_{t}{\check{e}}_{I}^{i}&=\partial_{t}(\theta^{-p_{I}}\widehat{e}_{I}^{\;i})=-\partial_{t}(p_{I}\ln\theta){\check{e}}_{I}^{i}+\theta^{-p_{I}}\partial_{t}\widehat{e}_{I}^{\;i}\\ &=-\partial_{t}(p_{I}\ln\theta){\check{e}}_{I}^{i}+\theta^{-p_{I}}\big(-\Theta p_{I}\widehat{e}_{I}^{\;i}+\textstyle\sum_{J\neq I}\tfrac{N^{2}}{\Theta(p_{I}-p_{J})}g\big((\mathcal{L}_{e_{0}}K)(\widehat{e}_{I}),\widehat{e}_{J}\big)\widehat{e}_{J}^{\;i}\big)\\ &=-\big(\partial_{t}(p_{I}\ln\theta)+\Theta p_{I}\big){\check{e}}_{I}^{i}+\textstyle\sum_{J\neq I}\tfrac{N^{2}}{\Theta(p_{I}-p_{J})}(\mathcal{L}_{e_{0}}K)\big({\check{e}}_{I},{\check{\omega}}^{J}\big){\check{e}}_{J}^{i}.\end{split}$
The result for ${\check{e}}_{I}^{i}$ now follows by computing $\partial_{t}\theta$ using (30a). Since $\delta^{I}_{J}={\check{\omega}}^{I}({\check{e}}_{J})$, (126b) follows from (126a). Finally, for $\Lambda_{IJ}^{K}$, note that
$\Lambda_{IJ}^{K}=\theta^{2p_{K}}g\big([{\check{e}}_{I},{\check{e}}_{J}],{\check{e}}_{K}\big).$
Hence, appealing to the Jacobi identity,
$\begin{split}\partial_{t}\Lambda_{IJ}^{K}&=\partial_{t}(\theta^{2p_{K}})g\big([{\check{e}}_{I},{\check{e}}_{J}],{\check{e}}_{K}\big)+N\theta^{2p_{K}}\Big(2k\big([{\check{e}}_{I},{\check{e}}_{J}],{\check{e}}_{K}\big)\\ &\quad+g\big([[e_{0},{\check{e}}_{I}],{\check{e}}_{J}],{\check{e}}_{K}\big)+g\big([{\check{e}}_{I},[e_{0},{\check{e}}_{J}]],{\check{e}}_{K}\big)+g\big([{\check{e}}_{I},{\check{e}}_{J}],[e_{0},{\check{e}}_{K}]\big)\Big).\end{split}$
Note that
$e_{0},{\check{e}}_{I}]=N^{-1}[\partial_{t},{\check{e}}_{I}]+{\check{e}}_{I}(\ln N)e_{0}=N^{-1}\mathcal{B}_{I}^{L}{\check{e}}_{L}+{\check{e}}_{I}(\ln N)e_{0},$
so that
$\begin{split}[[e_{0},{\check{e}}_{I}],{\check{e}}_{J}]&=N^{-1}\mathcal{B}_{I}^{L}[{\check{e}}_{L},{\check{e}}_{J}]+N^{-1}{\check{e}}_{J}(\ln N)\mathcal{B}_{I}^{L}{\check{e}}_{L}-N^{-1}{\check{e}}_{J}(\mathcal{B}_{I}^{L}){\check{e}}_{L}\\ &\quad+{\check{e}}_{I}(\ln N)[e_{0},{\check{e}}_{J}]-{\check{e}}_{J}{\check{e}}_{I}(\ln N)e_{0}.\end{split}$
Going back to $\partial_{t}\Lambda_{IJ}^{K}$, our previous observations yield (no summation on $K$)
$\partial_{t}\Lambda_{IJ}^{K}$ $=2\partial_{t}(p_{K}\ln\theta)\Lambda_{IJ}^{K}+2\Theta p_{K}\Lambda_{IJ}^{K}$
$\quad+\theta^{2p_{K}}\Big(\mathcal{B}_{I}^{L}g\big([{\check{e}}_{L},{\check{e}}_{J}],{\check{e}}_{K}\big)+{\check{e}}_{J}(\ln N)\mathcal{B}_{I}^{L}g({\check{e}}_{L},{\check{e}}_{K})-{\check{e}}_{J}\big(\mathcal{B}_{I}^{L}\big)g({\check{e}}_{L},{\check{e}}_{K})$
$\quad+N{\check{e}}_{I}(\ln N)g\big([e_{0},{\check{e}}_{J}],{\check{e}}_{K}\big)-\mathcal{B}_{J}^{L}g\big([{\check{e}}_{L},{\check{e}}_{I}],{\check{e}}_{K}\big)-{\check{e}}_{I}(\ln N)\mathcal{B}_{J}^{L}g({\check{e}}_{L},{\check{e}}_{K})$
$\quad+{\check{e}}_{I}\big(\mathcal{B}_{J}^{L}\big)g({\check{e}}_{L},{\check{e}}_{K})-N{\check{e}}_{J}(\ln N)g\big([e_{0},{\check{e}}_{I}],{\check{e}}_{K}\big)+\mathcal{B}_{K}^{L}g\big([{\check{e}}_{I},{\check{e}}_{J}],{\check{e}}_{L}\big)\Big)$
$=-2\mathcal{B}_{K}^{K}\Lambda_{IJ}^{K}+\mathcal{B}_{I}^{L}\Lambda_{LJ}^{K}-{\check{e}}_{J}\big(\mathcal{B}_{I}^{K}\big)-\mathcal{B}_{J}^{L}\Lambda_{LI}^{K}+{\check{e}}_{I}\big(\mathcal{B}_{J}^{K}\big)+\textstyle\sum_{L}\theta^{2(p_{K}-p_{L})}\mathcal{B}_{K}^{L}\Lambda_{IJ}^{L},$
where we used that $\mathcal{B}_{K}^{K}=\partial_{t}(p_{K}\ln\theta)+\Theta p_{K}$ (no summation). The lemma follows. ∎
3.4. Asymptotics
In order to use the equations of Lemma 51, we need an expression for the off-diagonal components of the spatial Ricci tensor in terms of ${\check{e}}_{I}$ and $\Lambda_{IJ}^{K}$.
Lemma 52.
Let $3\leq n\in\mathbb{N}$, $k_{0}\in\mathbb{N}$, $r>0$, $\delta\in(0,1)$ and fix a $\delta$-admissible potential $V\in C^{\infty}(\mathbb{R})$. Given that $(M,g,\varphi)$ is a solution to the Einstein–nonlinear scalar field equations satisfying the $(\delta,n,k_{0},r)$-assumptions, assume that there is an eigenframe $\{\widehat{e}_{I}\}_{I=1}^{n}$ with the properties listed at the beginning of Subsection 3.2. For $I\neq J$,
$\mathrm{Ric}_{h}^{\sharp}\big({\check{e}}_{I},{\check{\omega}}^{J}\big)=\mathcal{R}_{1,I}{}^{J}+\mathcal{R}_{2,I}{}^{J}+\mathcal{R}_{3,I}{}^{J}$,
where $\mathcal{R}_{i,I}{}^{J}$ are given schematically by
$\mathcal{R}_{1,I}{}^{J}$ $=\theta^{2p_{J}}*{\check{e}}_{I}*{\check{e}}_{J},$
$\mathcal{R}_{2,I}{}^{J}$ $=\theta^{2p_{J}}*{\check{e}}_{I}*\Lambda_{JK}^{K}+\theta^{2p_{J}}*{\check{e}}_{K}*\Lambda_{IJ}^{K}+\textstyle\sum_{K}\theta^{2(p_{J}+p_{K}-p_{I})}*{\check{e}}_{K}*\Lambda_{KJ}^{I}$
$\quad+\textstyle\sum_{K}\theta^{2p_{K}}*{\check{e}}_{K}*\Lambda_{IK}^{J}+\theta^{2p_{J}}*{\check{e}}_{J}*\Lambda_{IK}^{K},$
$\mathcal{R}_{3,I}{}^{J}$ $=\textstyle\sum_{K,L}\theta^{2(p_{K}+p_{L}-p_{I})}*\Lambda_{KL}^{I}*\Lambda_{KL}^{J}+\theta^{2p_{J}}*\Lambda_{IK}^{L}*\Lambda_{LJ}^{K}+\textstyle\sum_{K}\theta^{2p_{K}}*\Lambda_{IK}^{L}*\Lambda_{LK}^{J}$
$\quad+\textstyle\sum_{K,L}\theta^{2(p_{J}+p_{K}-p_{L})}*\Lambda_{IK}^{L}*\Lambda_{JK}^{L}+\textstyle\sum_{L}\theta^{2p_{L}}*\Lambda_{LK}^{K}*\Lambda_{IL}^{J}+\theta^{2p_{J}}*\Lambda_{LK}^{K}*\Lambda_{IJ}^{L}$
$\quad+\textstyle\sum_{L}\theta^{2(p_{L}+p_{J}-p_{I})}*\Lambda_{LK}^{K}*\Lambda_{LJ}^{I}+\textstyle\sum_{K}\theta^{2(p_{J}+p_{K}-p_{I})}*\Lambda_{KL}^{I}*\Lambda_{KJ}^{L}.$
Here, each type of term may contain up to two factors of the form $E_{\mathrm{\mathbf{I}}}(p_{I}\ln\theta)$, with $|\mathrm{\mathbf{I}}|\leq 2$. Moreover, each factor ${\check{e}}_{I}$ and $\Lambda_{IJ}^{K}$ may have at most one $E_{i}$ derivative applied to it.
Proof.
Define the connection coefficients
$\Gamma_{IJ}^{K}:={\check{\omega}}^{K}\big(\overline{\nabla}_{{\check{e}}_{I}}{\check{e}}_{J}\big).$
Then, by the Koszul formula,
$\begin{split}2\theta^{-2p_{K}}\Gamma_{IJ}^{K}&=2h\big(\overline{\nabla}_{{\check{e}}_{I}}{\check{e}}_{J},{\check{e}}_{K}\big)\\ &={\check{e}}_{I}(\theta^{-2p_{J}}\delta_{JK})+{\check{e}}_{J}(\theta^{-2p_{I}}\delta_{IK})-{\check{e}}_{K}(\theta^{-2p_{I}}\delta_{IJ})\\ &\quad+h\big({\check{e}}_{K},[{\check{e}}_{I},{\check{e}}_{J}]\big)-h\big({\check{e}}_{J},[{\check{e}}_{I},{\check{e}}_{K}]\big)-h\big({\check{e}}_{I},[{\check{e}}_{J},{\check{e}}_{K}]\big)\end{split}$
(no summation on any index). Hence,
(127) $\begin{split}\Gamma_{IJ}^{K}=&-{\check{e}}_{I}(p_{J}\ln\theta)\delta_{JK}-{\check{e}}_{J}(p_{I}\ln\theta)\delta_{IK}+\theta^{2(p_{K}-p_{I})}{\check{e}}_{K}(p_{I}\ln\theta)\delta_{IJ}\\ &+\tfrac{1}{2}\big(\Lambda_{IJ}^{K}-\theta^{2(p_{K}-p_{J})}\Lambda_{IK}^{J}-\theta^{2(p_{K}-p_{I})}\Lambda_{JK}^{I}\big).\end{split}$
Next,
$\begin{split}\mathrm{Ric}_{h}^{\sharp}({\check{e}}_{I})&=\textstyle\sum_{K}\theta^{2p_{K}}\bar{R}({\check{e}}_{I},{\check{e}}_{K}){\check{e}}_{K}\\ &=\textstyle\sum_{K}\theta^{2p_{K}}\big(\overline{\nabla}_{{\check{e}}_{I}}\overline{\nabla}_{{\check{e}}_{K}}{\check{e}}_{K}-\overline{\nabla}_{{\check{e}}_{K}}\overline{\nabla}_{{\check{e}}_{I}}{\check{e}}_{K}-\overline{\nabla}_{[{\check{e}}_{I},{\check{e}}_{K}]}{\check{e}}_{K}\big)\\ &=\textstyle\sum_{K}\theta^{2p_{K}}\big({\check{e}}_{I}\big(\Gamma_{KK}^{L}\big){\check{e}}_{L}-{\check{e}}_{K}\big(\Gamma_{IK}^{L}\big){\check{e}}_{L}+\Gamma_{KK}^{L}\Gamma_{IL}^{M}{\check{e}}_{M}-\Gamma_{IK}^{L}\Gamma_{KL}^{M}{\check{e}}_{M}-\Lambda_{IK}^{L}\Gamma_{LK}^{M}{\check{e}}_{M}\big).\end{split}$
Therefore,
$\mathrm{Ric}_{h}^{\sharp}\big({\check{e}}_{I},{\check{\omega}}^{J}\big)=\textstyle\sum_{K}\theta^{2p_{K}}\big({\check{e}}_{I}\big(\Gamma_{KK}^{J}\big)-{\check{e}}_{K}\big(\Gamma_{IK}^{J}\big)+\Gamma_{KK}^{L}\Gamma_{IL}^{J}-\Gamma_{IK}^{L}\Gamma_{KL}^{J}-\Lambda_{IK}^{L}\Gamma_{LK}^{J}\big).$
Now we expand each term in terms of $\Lambda_{IJ}^{K}$. First,
$\textstyle\sum_{K}\theta^{2p_{K}}{\check{e}}_{I}\big(\Gamma_{KK}^{J}\big)$ $=\theta^{2p_{J}}\big(-2{\check{e}}_{I}{\check{e}}_{J}(p_{J}\ln\theta)+{\check{e}}_{I}{\check{e}}_{J}(\ln\theta)$
$\quad+2\textstyle\sum_{K}{\check{e}}_{J}(p_{K}\ln\theta){\check{e}}_{I}\big((p_{J}-p_{K})\ln\theta\big)$
$\quad+{\check{e}}_{I}\big(\Lambda_{JK}^{K}\big)+2\textstyle\sum_{K}{\check{e}}_{I}\big((p_{J}-p_{K})\ln\theta\big)\Lambda_{JK}^{K}\big).$
Next,
$\textstyle\sum_{K}\theta^{2p_{K}}{\check{e}}_{K}\big(\Gamma_{IK}^{J}\big)$
$=$ $-\theta^{2p_{J}}{\check{e}}_{J}{\check{e}}_{I}(p_{J}\ln\theta)-\textstyle\sum_{K}\theta^{2p_{K}}{\check{e}}_{K}{\check{e}}_{K}(p_{I}\ln\theta)\delta_{IJ}+\theta^{2p_{J}}{\check{e}}_{I}{\check{e}}_{J}(p_{I}\ln\theta)$
$+2\theta^{2p_{J}}{\check{e}}_{J}(p_{I}\ln\theta){\check{e}}_{I}\big((p_{J}-p_{I})\ln\theta\big)+\tfrac{1}{2}\textstyle\sum_{K}\theta^{2p_{K}}{\check{e}}_{K}\big(\Lambda_{IK}^{J}\big)-\tfrac{1}{2}\textstyle\sum_{K}\theta^{2p_{J}}{\check{e}}_{K}\big(\Lambda_{IJ}^{K}\big)$
$-\textstyle\sum_{K}\theta^{2p_{J}}{\check{e}}_{K}\big((p_{J}-p_{K})\ln\theta\big)\Lambda_{IJ}^{K}-\tfrac{1}{2}\textstyle\sum_{K}\theta^{2(p_{K}+p_{J}-p_{I})}{\check{e}}_{K}\big(\Lambda_{KJ}^{I}\big)$
$-\textstyle\sum_{K}\theta^{2(p_{J}+p_{K}-p_{I})}{\check{e}}_{K}\big((p_{J}-p_{I})\ln\theta\big)\Lambda_{KJ}^{I}.$
In addition,
$\textstyle\sum_{K}\theta^{2p_{K}}\Lambda_{IK}^{L}\Gamma_{LK}^{J}=$ $-\theta^{2p_{J}}\Lambda_{IJ}^{L}{\check{e}}_{L}(p_{J}\ln\theta)-\textstyle\sum_{K}\theta^{2p_{K}}\Lambda_{IK}^{J}{\check{e}}_{K}(p_{J}\ln\theta)+\textstyle\sum_{K}\theta^{2p_{J}}{\check{e}}_{J}(p_{K}\ln\theta)\Lambda_{IK}^{K}$
$+\tfrac{1}{2}\textstyle\sum_{K}\theta^{2p_{K}}\Lambda_{IK}^{L}\Lambda_{LK}^{J}-\tfrac{1}{2}\theta^{2p_{J}}\Lambda_{IK}^{L}\Lambda_{LJ}^{K}+\tfrac{1}{2}\textstyle\sum_{K,L}\theta^{2(p_{J}+p_{K}-p_{L})}\Lambda_{IK}^{L}\Lambda_{JK}^{L}.$
Moreover,
$\textstyle\sum_{K}\theta^{2p_{K}}\Gamma_{KK}^{L}\Gamma_{IL}^{J}$ $=\textstyle\sum_{K,L}\theta^{2p_{K}}\big(-2{\check{e}}_{K}(p_{K}\ln\theta)\delta_{KL}+\theta^{2(p_{L}-p_{K})}{\check{e}}_{L}(p_{K}\ln\theta)$
$\quad+\theta^{2(p_{L}-p_{K})}\Lambda_{LK}^{K}\big)\big(-{\check{e}}_{I}(p_{L}\ln\theta)\delta_{LJ}-{\check{e}}_{L}(p_{I}\ln\theta)\delta_{IJ}$
$\quad+\theta^{2(p_{J}-p_{I})}{\check{e}}_{J}(p_{I}\ln\theta)\delta_{IL}+\tfrac{1}{2}\Lambda_{IL}^{J}-\tfrac{1}{2}\theta^{2(p_{J}-p_{L})}\Lambda_{IJ}^{L}-\tfrac{1}{2}\theta^{2(p_{J}-p_{I})}\Lambda_{LJ}^{I}\big).$
Finally,
$\textstyle\sum_{K}\theta^{2p_{K}}\Gamma_{IK}^{L}\Gamma_{KL}^{J}$ $=\textstyle\sum_{K,L}\theta^{2p_{K}}\big(-{\check{e}}_{I}(p_{K}\ln\theta)\delta_{KL}-{\check{e}}_{K}(p_{I}\ln\theta)\delta_{IL}+\theta^{2(p_{L}-p_{I})}{\check{e}}_{L}(p_{I}\ln\theta)\delta_{IK}$
$\quad+\tfrac{1}{2}\Lambda_{IK}^{L}-\tfrac{1}{2}\theta^{2(p_{L}-p_{K})}\Lambda_{IL}^{K}-\tfrac{1}{2}\theta^{2(p_{L}-p_{I})}\Lambda_{KL}^{I}\big)$
$\quad\,\big(-{\check{e}}_{K}(p_{L}\ln\theta)\delta_{LJ}-{\check{e}}_{L}(p_{K}\ln\theta)\delta_{KJ}+\theta^{2(p_{J}-p_{K})}{\check{e}}_{J}(p_{K}\ln\theta)\delta_{KL}$
$\quad+\tfrac{1}{2}\Lambda_{KL}^{J}-\tfrac{1}{2}\theta^{2(p_{J}-p_{L})}\Lambda_{KJ}^{L}-\tfrac{1}{2}\theta^{2(p_{J}-p_{K})}\Lambda_{LJ}^{K}\big).$
It may now be verified that for $I\neq J$, every term that appears in $\mathrm{Ric}_{h}^{\sharp}\big({\check{e}}_{I},{\check{\omega}}^{J}\big)$ is of one of the types described in the statement of the lemma. ∎
In the following two lemmas, we deduce the bounds for $\mathcal{B}_{I}^{J}$ that will be required for improving on the estimates for the expansion normalized eigenframe.
Lemma 53.
Let $3\leq n\in\mathbb{N}$, $k_{0}\in\mathbb{N}$, $r>0$, $\delta\in(0,1)$ and fix a $\delta$-admissible potential $V\in C^{\infty}(\mathbb{R})$. Given that $(M,g,\varphi)$ is a solution to the Einstein–nonlinear scalar field equations satisfying the $(\delta,n,k_{0},r)$-assumptions, assume that there is an eigenframe $\{\widehat{e}_{I}\}_{I=1}^{n}$ with the properties listed at the beginning of Subsection 3.2. Then there is a constant $C$ such that
$\big|E_{\mathrm{\mathbf{I}}}\big(\mathcal{B}_{I}^{I}\big)\big|\leq Ct^{-1+\delta},\qquad\big|E_{\mathrm{\mathbf{I}}}\big(\mathcal{B}_{I}^{J}\big)\big|\leq C\langle\ln t\rangle^{|\mathrm{\mathbf{I}}|}t^{-1+\mathring{p}_{I}-\mathring{p}_{J}+2\delta}$
for $t\leq t_{0}$, where $I\neq J$, there is no summation over $I$, and $|\mathrm{\mathbf{I}}|\leq k_{0}-1$.
Proof.
We begin by estimating $e_{0}p_{I}$ using Lemma 43. By (30a),
$e_{0}\theta$ $=-\mathrm{Scal}_{h}-\theta^{2}+N^{-1}\Delta_{h}N+|{\mathrm{d}}\varphi|_{h}^{2}+\tfrac{2n}{n-1}V\circ\varphi,$
$g\big((\mathcal{L}_{e_{0}}K)(\widehat{e}_{I}),\widehat{e}_{I}\big)$ $=-\mathrm{Ric}_{h}(\widehat{e}_{I},\widehat{e}_{I})-\theta^{2}p_{I}+N^{-1}(\overline{\nabla}^{2}N)(\widehat{e}_{I},\widehat{e}_{I})+\widehat{e}_{I}(\varphi)^{2}+\tfrac{2}{n-1}V\circ\varphi$
(no summation). Note that the terms involving $\theta^{2}$ cancel out in the right-hand side of (119b). Hence, we can estimate the remaining terms using the assumptions, Lemma 41, Lemma 46 and Proposition 47 to obtain
$t\|e_{0}p_{I}\|_{C^{k_{0}-1}(\Sigma)}\leq C\langle\ln t\rangle^{k_{0}-1}t^{2\delta}.$
The result for $\mathcal{B}_{I}^{I}$ now follows from its definition and Lemma 46.
For $\mathcal{B}_{I}^{J}$, we have
$(\mathcal{L}_{e_{0}}K)\big({\check{e}}_{I},{\check{\omega}}^{J}\big)=\theta^{p_{J}-p_{I}}g\big((\mathcal{L}_{e_{0}}K)(\widehat{e}_{I}),\widehat{e}_{J}\big).$
The result follows from the assumptions, the definition of $\mathcal{B}_{I}^{J}$, (30a), Lemmas 39 and 46, and Proposition 47. ∎
Lemma 54.
Let $3\leq n\in\mathbb{N}$, $2\leq k_{0}\in\mathbb{N}$, $r>0$, $\delta\in(0,1)$ and fix a $\delta$-admissible potential $V\in C^{\infty}(\mathbb{R})$. Given that $(M,g,\varphi)$ is a solution to the Einstein–nonlinear scalar field equations satisfying the $(\delta,n,k_{0},r)$-assumptions, assume that there is an eigenframe $\{\widehat{e}_{I}\}_{I=1}^{n}$ with the properties listed at the beginning of Subsection 3.2. Let $0<\mathbf{c}_{1}\in\mathbb{R}$, $1\leq k\leq k_{0}-1$, $x\in\Sigma$, and suppose that the following estimates hold for some $m\in\mathbb{N}$, $t\leq t_{0}$ and all $|\mathrm{\mathbf{I}}|\leq k$: For each ${\check{e}}_{I}$ separately, either
(128) $\big|E_{\mathrm{\mathbf{I}}}\big({\check{e}}_{I}^{i}\big)(x)\big|\leq\mathbf{c}_{1}t^{-1+\mathring{p}_{I}(x)+m\delta},\quad\textrm{or}\quad\big|E_{\mathrm{\mathbf{I}}}\big({\check{e}}_{I}^{i}\big)(x)\big|\leq\mathbf{c}_{1};$
and for each $\Lambda_{IJ}^{K}$ separately, either
$\big|E_{\mathrm{\mathbf{I}}}\big(\Lambda_{IJ}^{K}\big)(x)\big|\leq\mathbf{c}_{1}t^{-1+\mathring{p}_{I}(x)+\mathring{p}_{J}(x)-\mathring{p}_{K}(x)+m\delta},\quad\textrm{or}\quad\big|E_{\mathrm{\mathbf{I}}}\big(\Lambda_{IJ}^{K}\big)(x)\big|\leq\mathbf{c}_{1}.$
Then there is a constant $C_{2}$, depending only on $\mathbf{c}_{1}$ and the constants appearing in the assumptions of Theorem 24 and the previous lemmas of the present section, such that, for $t\leq t_{0}$ and $I\neq J$,
$\big|E_{\mathrm{\mathbf{I}}}\big(\mathcal{B}_{I}^{J}\big)(x)\big|\leq C_{2}t^{-1+\mathring{p}_{I}(x)-\mathring{p}_{J}(x)+(m+1)\delta}+C_{2}t^{-1+2\delta}\min\{1,t^{2(\mathring{p}_{I}(x)-\mathring{p}_{J}(x))}\},$
where $|\mathrm{\mathbf{I}}|\leq k-1$. In particular, if all the $E_{\mathrm{\mathbf{I}}}\big({\check{e}}_{I}^{i}\big)(x)$ and the $E_{\mathrm{\mathbf{I}}}\big(\Lambda_{IJ}^{K}\big)(x)$ are bounded, then
$\big|E_{\mathrm{\mathbf{I}}}\big(\mathcal{B}_{I}^{J}\big)(x)\big|\leq C_{2}t^{-1+2\delta}\min\{1,t^{2(\mathring{p}_{I}(x)-\mathring{p}_{J}(x))}\}.$
Proof.
For each ${\check{e}}_{I}$, we, in this proof, say that case 1 (or 2) holds if the left (or right) inequality in (128) holds. We use similar terminology for the $\Lambda_{IJ}^{K}$. In addition, when we write $\langle\ln t\rangle^{*}$, it is understood that * stands for a positive integer whose exact value is unimportant (but only depends on $k$). In particular, its exact value may change from line to line. Finally, to simplify the notation, we omit the argument $x$ throughout the proof. Thus, it is to be understood that all functions on $\Sigma$ that appear below are evaluated at the point $x\in\Sigma$.
Let $I\neq J$. We illustrate the idea with one of the types of terms appearing in Lemma 52. Let $|\mathrm{\mathbf{I}}|\leq k-1$. Consider $\theta^{2p_{J}}*{\check{e}}_{I}*{\check{e}}_{J}$. If case 1 holds for both ${\check{e}}_{I}$ and ${\check{e}}_{J}$, then, keeping Lemma 50 in mind,
$|E_{\mathrm{\mathbf{I}}}(\theta^{2p_{J}}*{\check{e}}_{I}*{\check{e}}_{J})|\leq C\langle\ln t\rangle^{*}t^{-2+\mathring{p}_{I}-\mathring{p}_{J}+2\max\{m,2\}\delta}.$
If case 2 holds for ${\check{e}}_{I}$, and case 1 holds for ${\check{e}}_{J}$, Lemma 37 implies
$|E_{\mathrm{\mathbf{I}}}(\theta^{2p_{J}}*{\check{e}}_{I}*{\check{e}}_{J})|\leq C\langle\ln t\rangle^{*}t^{-1-\mathring{p}_{J}+m\delta}\leq C\langle\ln t\rangle^{*}t^{-2+\mathring{p}_{I}-\mathring{p}_{J}+(m+2)\delta}.$
If case 1 holds for ${\check{e}}_{I}$ and case 2 holds for ${\check{e}}_{J}$, a similar estimate results. If case 2 holds for both and $I<J$, then
(129) $|E_{\mathrm{\mathbf{I}}}(\theta^{2p_{J}}*{\check{e}}_{I}*{\check{e}}_{J})|\leq C\langle\ln t\rangle^{*}t^{-2\mathring{p}_{J}}\leq C\langle\ln t\rangle^{*}t^{-2+4\delta};$
and if $I>J$, then
(130) $|E_{\mathrm{\mathbf{I}}}(\theta^{2p_{J}}*{\check{e}}_{I}*{\check{e}}_{J})|\leq C\langle\ln t\rangle^{*}t^{-2+2(\mathring{p}_{I}-\mathring{p}_{J})+2(1-\mathring{p}_{I})}\leq C\langle\ln t\rangle^{*}t^{-2+2(\mathring{p}_{I}-\mathring{p}_{J})+4\delta}.$
If $\theta*{\check{e}}*\Lambda$ and $\theta*\Lambda*\Lambda$ are any of the types of terms appearing in $\mathrm{Ric}_{h}^{\,\sharp}\big({\check{e}}_{I},{\check{\omega}}^{J}\big)$, it is readily verified that, whenever both of the non $\theta$ factors are in case 1, or one is in case 1 and the other in case 2, the same pattern as above arises. Thus
$|E_{\mathrm{\mathbf{I}}}(\theta*{\check{e}}*\Lambda)|+|E_{\mathrm{\mathbf{I}}}(\theta*\Lambda*\Lambda)|\leq C\langle\ln t\rangle^{*}t^{-2+\mathring{p}_{I}-\mathring{p}_{J}+(m+2)\delta}$
when both non $\theta$ factors are in case 1, or one is in case 1 and the other in case 2. Note, however, that if one is in case 1 and the other in case 2, we sometimes need to appeal to (34). It remains to check what happens when both factors are in case 2. Note that in that case, the only time dependent contribution to the bound comes from the power of $\theta$.
The following are the possible upper bounds for the terms in $\mathrm{Ric}_{h}^{\,\sharp}\big({\check{e}}_{I},{\check{\omega}}^{J}\big)$, for $I\neq J$, when both of the factors are in case 2:
(131a) $C\langle\ln t\rangle^{*}t^{-2\mathring{p}_{J}};$
(131b) $C\langle\ln t\rangle^{*}t^{2(\mathring{p}_{I}-\mathring{p}_{J}-\mathring{p}_{K})},$ $K$ $\neq J;$
(131c) $C\langle\ln t\rangle^{*}t^{-2\mathring{p}_{K}},$ $K$ $\neq I;$
(131d) $C\langle\ln t\rangle^{*}t^{2(\mathring{p}_{I}-\mathring{p}_{K}-\mathring{p}_{L})},$ $K$ $\neq L;$
(131e) $C\langle\ln t\rangle^{*}t^{2(\mathring{p}_{L}-\mathring{p}_{J}-\mathring{p}_{K})},$ $K$ $\neq J,\quad K\neq I.$
All the expressions in (131) can be estimated by the far right hand side of (129) due to (34) and Lemma 37. Assume now that $I>J\textsterling$. Then the expression (131a) can be estimated as in (130). In the case of (131b),
$2(\mathring{p}_{I}-\mathring{p}_{J}-\mathring{p}_{K})=-2+2(\mathring{p}_{I}-\mathring{p}_{J})+2(1-\mathring{p}_{K})>-2+2(\mathring{p}_{I}-\mathring{p}_{J})+4\delta.$
In the case of (131c), since $K\neq I$,
$2\mathring{p}_{K}=-2+2(\mathring{p}_{I}-\mathring{p}_{J})+2(1+\mathring{p}_{J}-\mathring{p}_{I}-\mathring{p}_{K})>-2+2(\mathring{p}_{I}-\mathring{p}_{J})+4\delta.$
In the case of (131d), since $K\neq L$,
$2(\mathring{p}_{I}-\mathring{p}_{K}-\mathring{p}_{L})=-2+2(\mathring{p}_{I}-\mathring{p}_{J})+2(1+\mathring{p}_{J}-\mathring{p}_{K}-\mathring{p}_{L})>-2+2(\mathring{p}_{I}-\mathring{p}_{J})+4\delta.$
Finally, in the case of (131e), since $K\neq I$,
$2(\mathring{p}_{L}-\mathring{p}_{J}-\mathring{p}_{K})=-2+2(\mathring{p}_{I}-\mathring{p}_{J})+2(1+\mathring{p}_{L}-\mathring{p}_{I}-\mathring{p}_{K})>-2+2(\mathring{p}_{I}-\mathring{p}_{J})+4\delta.$
Summing up, we have shown that
$\big|E_{\mathrm{\mathbf{I}}}\big(\mathrm{Ric}_{h}^{\sharp}\big({\check{e}}_{I},{\check{\omega}}^{J}\big)\big)\big|\leq Ct^{-2+\mathring{p}_{I}-\mathring{p}_{J}+(m+1)\delta}+Ct^{-2+2\delta}\min\{1,t^{2(\mathring{p}_{I}-\mathring{p}_{J})}\},$
for $I\neq J$ and $|\mathrm{\mathbf{I}}|\leq k-1$. In particular, if all the $E_{\mathrm{\mathbf{I}}}\big({\check{e}}_{I}^{i}\big)$ and the $E_{\mathrm{\mathbf{I}}}\big(\Lambda_{IJ}^{K}\big)$ are bounded for $|\mathrm{\mathbf{I}}|\leq k$, then
$\big|E_{\mathrm{\mathbf{I}}}\big(\mathrm{Ric}_{h}^{\sharp}\big({\check{e}}_{I},{\check{\omega}}^{J}\big)\big)\big|\leq Ct^{-2+2\delta}\min\{1,t^{2(\mathring{p}_{I}-\mathring{p}_{J})}\}.$
Next, we consider the remaining terms in $\mathcal{B}_{I}^{J}$, starting with
(132) $({\mathrm{d}}\varphi\otimes\mathrm{grad}_{h}\varphi)\big({\check{e}}_{I},{\check{\omega}}^{J}\big)=\theta^{2p_{J}}{\check{e}}_{I}(\varphi){\check{e}}_{J}(\varphi).$
Moreover,
(133) $\begin{split}(\overline{\nabla}^{2}N)^{\sharp}\big({\check{e}}_{I},{\check{\omega}}^{J}\big)&=\theta^{2p_{J}}(\overline{\nabla}^{2}N)({\check{e}}_{I},{\check{e}}_{J})\\ &=\theta^{2p_{J}}{\check{e}}_{I}{\check{e}}_{J}N+\theta^{2p_{J}}{\check{e}}_{I}(p_{J}\ln\theta){\check{e}}_{J}(N)+\theta^{2p_{J}}{\check{e}}_{J}(p_{I}\ln\theta){\check{e}}_{I}(N)\\ &\quad-\tfrac{1}{2}\theta^{2p_{J}}\Lambda_{IJ}^{K}{\check{e}}_{K}(N)+\tfrac{1}{2}\textstyle\sum_{K}\theta^{2p_{K}}\Lambda_{IK}^{J}{\check{e}}_{K}(N)\\ &\quad+\tfrac{1}{2}\textstyle\sum_{K}\theta^{2(p_{J}+p_{K}-p_{I})}\Lambda_{JK}^{I}{\check{e}}_{K}(N).\end{split}$
This observation finishes the proof, since each of the terms on the far right hand sides of (132) and (133) is qualitatively of the same type as a term that appeared when estimating the components of the spatial Ricci tensor. ∎
Next, we prove the crucial inductive step required to obtain the asymptotics: it is possible to improve on the estimates for the expansion normalized eigenframe, at the cost of losing two derivatives at each step.
Lemma 55.
Let $3\leq n\in\mathbb{N}$, $4\leq k_{0}\in\mathbb{N}$, $r>0$, $\delta\in(0,1)$ and fix a $\delta$-admissible potential $V\in C^{\infty}(\mathbb{R})$. Given that $(M,g,\varphi)$ is a solution to the Einstein–nonlinear scalar field equations satisfying the $(\delta,n,k_{0},r)$-assumptions, assume that there is an eigenframe $\{\widehat{e}_{I}\}_{I=1}^{n}$ with the properties listed at the beginning of Subsection 3.2. Fix an open cover $U_{j}$ of $\Sigma$, with corresponding constants $\alpha_{j}$, as in Lemma 42, and let $m\in\mathbb{N}$. Suppose that for each $j$, the following pointwise estimates hold in $U_{j}$ for all $|\mathrm{\mathbf{I}}|\leq k$, $t\leq t_{0}$ and some $k\leq k_{0}-2$: for ${\check{e}}_{I}$,
$\big|E_{\mathrm{\mathbf{I}}}\big({\check{e}}_{I}^{i}\big)\big|$ $\leq C_{j}t^{-1+\mathring{p}_{I}+m\delta+\alpha_{j}},$ $\text{if}\quad-1+\mathring{p}_{I}+m\delta+\alpha_{j}$ $<0,$
$\big|E_{\mathrm{\mathbf{I}}}\big({\check{e}}_{I}^{i}\big)\big|$ $\leq C_{j},$ $\text{if}\quad-1+\mathring{p}_{I}+m\delta+\alpha_{j}$ $>0;$
for ${\check{\omega}}^{I}$,
$\big|E_{\mathrm{\mathbf{I}}}\big({\check{\omega}}_{i}^{I}\big)\big|$ $\leq C_{j}t^{-1-\mathring{p}_{I}+m\delta+\alpha_{j}},$ $\text{if}\quad-1-\mathring{p}_{I}+m\delta+\alpha_{j}$ $<0,$
$\big|E_{\mathrm{\mathbf{I}}}\big({\check{\omega}}_{i}^{I}\big)\big|$ $\leq C_{j},$ $\text{if}\quad-1-\mathring{p}_{I}+m\delta+\alpha_{j}$ $>0;$
and for $\Lambda_{IJ}^{K}$,
$\big|E_{\mathrm{\mathbf{I}}}\big(\Lambda_{IJ}^{K}\big)\big|$ $\leq C_{j}t^{-1+\mathring{p}_{I}+\mathring{p}_{J}-\mathring{p}_{K}+m\delta+\alpha_{j}},$ $\text{if}\quad-1+\mathring{p}_{I}+\mathring{p}_{J}-\mathring{p}_{K}+m\delta+\alpha_{j}$ $<0,$
$\big|E_{\mathrm{\mathbf{I}}}\big(\Lambda_{IJ}^{K}\big)\big|$ $\leq C_{j},$ $\text{if}\quad-1+\mathring{p}_{I}+\mathring{p}_{J}-\mathring{p}_{K}+m\delta+\alpha_{j}$ $>0.$
Then the same estimates hold for $|\mathrm{\mathbf{I}}|\leq k-2$ and $t\leq t_{0}$, with $m$ replaced by $m+1$.
Proof.
Fix one of the open sets $U_{j}$. For the rest of the proof it is understood that all functions on $\Sigma$ are restricted to $U_{j}$. For each ${\check{e}}_{I}$, we say that case 1 holds if $1+\mathring{p}_{I}+m\delta+\alpha_{j}<0$, and we say that case 2 holds if $1+\mathring{p}_{I}+m\delta+\alpha_{j}>0$. The conventions concerning the ${\check{\omega}}^{I}$ and the $\Lambda_{IJ}^{K}$ are similar. Moreover, we use the notation $\langle\ln t\rangle^{*}$ introduced in the proof of Lemma 54.
Consider $\Lambda_{IJ}^{K}$. The idea is to estimate $\partial_{t}\Lambda_{IJ}^{K}$ by estimating all the terms on the right-hand side of (126c). Let $|\mathrm{\mathbf{I}}|\leq k-2$. By Lemma 53, if case 1 holds for $\Lambda_{IJ}^{K}$,
$\big|E_{\mathrm{\mathbf{I}}}\big(\mathcal{B}_{L}^{L}\Lambda_{IJ}^{K}\big)\big|\leq C_{j}t^{-2+\mathring{p}_{I}+\mathring{p}_{J}-\mathring{p}_{K}+(m+1)\delta+\alpha_{j}},$
(no summation on $L$). Next, we look at $\sum_{L\neq K}\theta^{2(p_{K}-p_{L})}\mathcal{B}_{K}^{L}\Lambda_{IJ}^{L}$. If case 1 holds for $\Lambda_{IJ}^{L}$, then Lemma 53 implies
$\begin{split}\big|E_{\mathrm{\mathbf{I}}}\big(\theta^{2(p_{K}-p_{L})}\mathcal{B}_{K}^{L}\Lambda_{IJ}^{L}\big)\big|&\leq C_{j}\langle\ln t\rangle^{*}t^{-2(\mathring{p}_{K}-\mathring{p}_{L})}t^{-1+\mathring{p}_{K}-\mathring{p}_{L}+2\delta}t^{-1+\mathring{p}_{I}+\mathring{p}_{J}-\mathring{p}_{L}+m\delta+\alpha_{j}}\\ &=C_{j}\langle\ln t\rangle^{*}t^{-2+\mathring{p}_{I}+\mathring{p}_{J}-\mathring{p}_{K}+(m+2)\delta+\alpha_{j}};\end{split}$
and if case 2 holds for $\Lambda_{IJ}^{L}$, then Lemma 54 and (34) yield ($I\neq J$)
$\begin{split}\big|E_{\mathrm{\mathbf{I}}}\big(\theta^{2(p_{K}-p_{L})}\mathcal{B}_{K}^{L}\Lambda_{IJ}^{L}\big)\big|&\leq C_{j}\langle\ln t\rangle^{*}t^{-2(\mathring{p}_{K}-\mathring{p}_{L})}\big(t^{-1+\mathring{p}_{K}-\mathring{p}_{L}+(m+1)\delta}+t^{-1+2\delta}\min\{1,t^{2(\mathring{p}_{K}-\mathring{p}_{L})}\}\big)\\ &\leq C_{j}\langle\ln t\rangle^{*}\big(t^{-2+\mathring{p}_{I}+\mathring{p}_{J}-\mathring{p}_{K}+(m+3)\delta}+t^{-1+2\delta}\big).\end{split}$
The reader may verify that the remaining terms on the right-hand side of (126c) can be estimated in a similar way. We conclude that
$\big|E_{\mathrm{\mathbf{I}}}\big(\partial_{t}\Lambda_{IJ}^{K}\big)\big|\leq C_{j}t^{-2+\mathring{p}_{I}+\mathring{p}_{J}-\mathring{p}_{K}+(m+1)\delta+\alpha_{j}}+C_{j}t^{-1+\delta}.$
Note that Lemma 42 ensures that (118a) holds in $U_{j}$ for $I\neq J$. Therefore, for $s<t_{0}$,
$\begin{split}\big|E_{\mathrm{\mathbf{I}}}\big(\Lambda_{IJ}^{K}\big)(t_{0})-E_{\mathrm{\mathbf{I}}}\big(\Lambda_{IJ}^{K}\big)(s)\big|&\leq\textstyle\int_{s}^{t_{0}}\big|E_{\mathrm{\mathbf{I}}}\big(\partial_{t}\Lambda_{IJ}^{K}\big)(t)\big|{\mathrm{d}}t\\ &\leq C_{j}\big(t_{0}^{\delta}-s^{\delta}\big)\\ &\quad+C_{j}\frac{t_{0}^{-1+\mathring{p}_{I}+\mathring{p}_{J}-\mathring{p}_{K}+(m+1)\delta+\alpha_{j}}-s^{-1+\mathring{p}_{I}+\mathring{p}_{J}-\mathring{p}_{K}+(m+1)\delta+\alpha_{j}}}{-1+\mathring{p}_{I}+\mathring{p}_{J}-\mathring{p}_{K}+(m+1)\delta+\alpha_{j}}.\end{split}$
There is thus a constant $C_{j}$ such that, if $1+\mathring{p}_{I}+\mathring{p}_{J}-\mathring{p}_{K}+(m+1)\delta+\alpha_{j}>0$,
$\big|E_{\mathrm{\mathbf{I}}}\big(\Lambda_{IJ}^{K}\big)(s)\big|\leq C_{j}$
in $U_{j}$ and for all $s<t_{0}$. On the other hand, if $1+\mathring{p}_{I}+\mathring{p}_{J}-\mathring{p}_{K}+(m+1)\delta+\alpha_{j}<0$, then
$\big|E_{\mathrm{\mathbf{I}}}\big(\Lambda_{IJ}^{K}\big)(s)\big|\leq C_{j}s^{-1+\mathring{p}_{I}+\mathring{p}_{J}-\mathring{p}_{K}+(m+1)\delta+\alpha_{j}}$
in $U_{j}$ and for all $s<t_{0}$. This proves the result for the $\Lambda_{IJ}^{K}$. The result for the ${\check{e}}_{I}$ and the ${\check{\omega}}^{I}$ follows similarly by using (126a) and (126b). ∎
Now we are ready to deduce asymptotics for ${\check{e}}_{I}$ and ${\check{\omega}}^{I}$.
Theorem 56.
Let $3\leq n\in\mathbb{N}$, $\ell\in\mathbb{N}$, $r>0$, $\delta\in(0,1)$ and fix a $\delta$-admissible potential $V\in C^{\infty}(\mathbb{R})$. If $k_{0}$ is large enough, depending only on $\delta$ and $\ell$; $(M,g,\varphi)$ is a solution to the Einstein–nonlinear scalar field equations satisfying the $(\delta,n,k_{0},r)$-assumptions; and there is an eigenframe $\{\widehat{e}_{I}\}_{I=1}^{n}$ with the properties listed at the beginning of Subsection 3.2, then there is a $C^{\ell}$ frame $\{\mathring{e}_{I}\}_{I=1}^{n}$ on $\Sigma$, with dual frame $\{\mathring{\omega}^{I}\}_{I=1}^{n}$, such that
$\|{\check{e}}_{I}(t)-\mathring{e}_{I}\|_{C^{\ell}(\Sigma)}+\|{\check{\omega}}^{I}(t)-\mathring{\omega}^{I}\|_{C^{\ell}(\Sigma)}\leq Ct^{\delta}$
[…]
(134) $\mathring{\mathcal{H}}:=\textstyle\sum_{I}\mathring{\omega}^{I}\otimes\mathring{\omega}^{I},\qquad\mathring{\mathcal{K}}:=\textstyle\sum_{I}\mathring{p}_{I}\mathring{\omega}^{I}\otimes\mathring{e}_{I}.$
Then
$\|\mathcal{H}(t)-\mathring{\mathcal{H}}\|_{C^{\ell}(\Sigma)}+\|\mathcal{K}(t)-\mathring{\mathcal{K}}\|_{C^{\ell}(\Sigma)}\leq Ct^{\delta}$
for all $t\leq t_{0}$. Finally, $(\Sigma,\mathring{\mathcal{H}},\mathring{\mathcal{K}},\mathring{\Phi},\mathring{\Psi})$ are $C^{\ell}$ robust nondegenerate quiescent initial data on the singularity for the Einstein–nonlinear scalar field equations.
Proof.
First, note that if $I$, $J$ and $K$ are distinct, then
$1-\mathring{p}_{I}-\mathring{p}_{J}+\mathring{p}_{K}\leq 1+\sqrt{3}\big(\mathring{p}_{I}^{2}+\mathring{p}_{J}^{2}+\mathring{p}_{K}^{2}\big)^{1/2}\leq 1+\sqrt{3}.$
Moreover, by Lemma 37, $1\pm\mathring{p}_{I}<2-2\delta$ for all $I$. Hence, there is a large enough positive integer $N$, depending only on $\delta$, such that
$1+\mathring{p}_{I}+\mathring{p}_{J}-\mathring{p}_{K}+N\delta>0,\qquad-1-\mathring{p}_{I}+N\delta>0,$
where $I\neq J$ in the first inequality. Fix an open cover $U_{j}$ of $\Sigma$ as in Lemma 42. Starting from the bounds of Lemma 50, we can now apply Lemma 55 repeatedly, at most N1N-1 times, to obtain
$\big|E_{\mathrm{\mathbf{I}}}\big({\check{e}}_{I}^{i}\big)\big|+\big|E_{\mathrm{\mathbf{I}}}\big({\check{\omega}}_{i}^{I}\big)\big|+\big|E_{\mathrm{\mathbf{I}}}\big(\Lambda_{IJ}^{K}\big)\big|\leq C_{j}$
in $(0,t_{0}]\times U_{j}$, and for $|\mathrm{\mathbf{I}}|\leq\ell+2$, assuming $k_{0}\geq 2N+\ell+2$. By taking the maximum of the constants $C_{j}$, we get a constant $C$ such that
$\big|E_{\mathrm{\mathbf{I}}}\big({\check{e}}_{I}^{i}\big)\big|+\big|E_{\mathrm{\mathbf{I}}}\big({\check{\omega}}_{i}^{I}\big)\big|+\big|E_{\mathrm{\mathbf{I}}}\big(\Lambda_{IJ}^{K}\big)\big|\leq C,$
for $|\mathrm{\mathbf{I}}|\leq\ell+2$, in $(0,t_{0}]\times\Sigma$.
Now knowing that all the ${\check{e}}_{I}$ and the $\Lambda_{IJ}^{K}$ are bounded in $C^{\ell+1}(\Sigma)$ for all $t\in(0,t_{0}]$, Lemma 54 yields
$\big|E_{\mathrm{\mathbf{I}}}\big(\mathcal{B}_{I}^{J}\big)\big|\leq Ct^{-1+2\delta}\min\{1,t^{2(\mathring{p}_{I}-\mathring{p}_{J})}\},$
for $I\neq J$ and $|\mathrm{\mathbf{I}}|\leq\ell+1$. Combining this estimate with Lemmas 39, 51 and 53 yields
$\|\partial_{t}{\check{e}}_{I}^{i}\|_{C^{\ell}(\Sigma)}+\|\partial_{t}{\check{\omega}}_{i}^{I}\|_{C^{\ell}(\Sigma)}+\|\partial_{t}\Lambda_{IJ}^{K}\|_{C^{\ell}(\Sigma)}\leq Ct^{-1+\delta}.$
But then, for $u<s\leq t_{0}$,
$\big|E_{\mathrm{\mathbf{I}}}\big({\check{e}}_{I}^{i}\big)(s)-E_{\mathrm{\mathbf{I}}}\big({\check{e}}_{I}^{i}\big)(u)\big|\leq\textstyle\int_{u}^{s}\big|E_{\mathrm{\mathbf{I}}}\big(\partial_{t}{\check{e}}_{I}^{i}\big)(t)\big|{\mathrm{d}}t\leq C(s^{\delta}-u^{\delta}).$
for all $t\leq t_{0}$ and all $|\mathrm{\mathbf{I}}|\leq\ell$. Thus, there are functions $\mathring{e}_{I}^{i}\in C^{\ell}(\Sigma)$ such that
$\|{\check{e}}_{I}^{i}(t)-\mathring{e}_{I}^{i}\|_{C^{\ell}(\Sigma)}\leq Ct^{\delta},$
and we define $\mathring{e}_{I}:=\mathring{e}_{I}^{i}E_{i}$. Similarly, there are functions $\mathring{\omega}_{i}^{I},\mathring{\Lambda}_{IJ}^{K}\in C^{\ell}(\Sigma)$ such that
$\|{\check{\omega}}_{i}^{I}(t)-\mathring{\omega}_{i}^{I}\|_{C^{\ell}(\Sigma)}+\|\Lambda_{IJ}^{K}(t)-\mathring{\Lambda}_{IJ}^{K}\|_{C^{\ell}(\Sigma)}\leq Ct^{\delta},$
and we define $\mathring{\omega}^{I}:=\mathring{\omega}_{i}^{I}\eta^{i}$. Letting $t\downarrow 0$ in the relation ${\check{\omega}}_{i}^{J}{\check{e}}_{I}^{i}=\delta_{I}^{J}$, it is clear that $\{\mathring{e}_{I}\}_{I=1}^{n}$ is a well defined frame on $\Sigma$ with dual frame $\{\mathring{\omega}^{I}\}_{I=1}^{n}$. Moreover, since
$\Lambda_{IJ}^{K}={\check{\omega}}^{K}\big([{\check{e}}_{I},{\check{e}}_{J}]\big)={\check{e}}_{I}\big({\check{e}}_{J}^{j}\big){\check{\omega}}_{j}^{K}-{\check{e}}_{J}\big({\check{e}}_{I}^{i}\big){\check{\omega}}_{i}^{K}+{\check{e}}_{I}^{i}{\check{e}}_{J}^{j}{\check{\omega}}^{K}\big([E_{i},E_{j}]\big),$
we can let $t\downarrow 0$ to conclude that $\mathring{\Lambda}_{IJ}^{K}$ are the structure coefficients of the frame $\{\mathring{e}_{I}\}_{I=1}^{n}$. It only remains to prove the statement about the initial data on the singularity. Note, to this end, that
$\mathcal{H}=\textstyle\sum_{I}{\check{\omega}}^{I}\otimes{\check{\omega}}^{I},\qquad\mathcal{K}=\textstyle\sum_{I}p_{I}{\check{\omega}}^{I}\otimes{\check{e}}_{I}.$
[…]
$\|\mathcal{H}(t)-\mathring{\mathcal{H}}\|_{C^{\ell}(\Sigma)}+\|\mathcal{K}(t)-\mathring{\mathcal{K}}\|_{C^{\ell}(\Sigma)}\leq Ct^{\delta}.$
Now, we verify that the conditions of initial data on the singularity hold. First,
$\mathrm{tr}\mathring{\mathcal{K}}=\textstyle\sum_{I}\mathring{p}_{I}=1,\qquad\mathrm{tr}\mathring{\mathcal{K}}^{2}+\mathring{\Psi}^{2}=\textstyle\sum_{I}\mathring{p}_{I}^{2}+\mathring{\Psi}^{2}=1.$
[…]
$\mathrm{div}_{h}K({\check{e}}_{I})$
$=\textstyle\sum_{J}\theta^{2p_{J}}h\big(\overline{\nabla}_{{\check{e}}_{J}}K({\check{e}}_{I}),{\check{e}}_{J}\big)$
$=\textstyle\sum_{J}\theta^{2p_{J}}h\big(\overline{\nabla}_{{\check{e}}_{J}}(\theta p_{I}{\check{e}}_{I})-\Gamma_{JI}^{K}K({\check{e}}_{K}),{\check{e}}_{J}\big)$
$=\textstyle\sum_{J}\theta^{2p_{J}}h\big({\check{e}}_{J}(\theta p_{I}){\check{e}}_{I}+\theta p_{I}\Gamma_{JI}^{L}{\check{e}}_{L}-\textstyle\sum_{K}\theta p_{K}\Gamma_{JI}^{K}{\check{e}}_{K},{\check{e}}_{J}\big)$
$={\check{e}}_{I}(\theta p_{I})+\textstyle\sum_{J}\theta(p_{I}-p_{J})\Gamma_{JI}^{J}$
$=\theta{\check{e}}_{I}(p_{I})+p_{I}{\check{e}}_{I}(\theta)+\textstyle\sum_{J}\theta(p_{I}-p_{J})\big(-{\check{e}}_{I}(p_{J}\ln\theta)+\Lambda_{JI}^{J}\big)$
$=\theta{\check{e}}_{I}(p_{I})+p_{I}{\check{e}}_{I}(\theta)-\theta p_{I}{\check{e}}_{I}\big(\textstyle\sum_{J}p_{J}\ln\theta\big)+\displaystyle\textstyle\sum_{J}\theta p_{J}{\check{e}}_{I}(p_{J}\ln\theta)+\textstyle\sum_{J}\theta(p_{I}-p_{J})\Lambda_{JI}^{J}$
$=\theta{\check{e}}_{I}(p_{I})+\theta{\check{e}}_{I}(\ln\theta)\textstyle\sum_{J}p_{J}^{2}+\tfrac{\theta\ln\theta}{2}{\check{e}}_{I}\big(\textstyle\sum_{J}p_{J}^{2}\big)+\displaystyle\textstyle\sum_{J}\theta(p_{I}-p_{J})\Lambda_{JI}^{J}$
(no summation on $I$). On the other hand,
$e_{0}(\varphi){\check{e}}_{I}(\varphi)=\theta\Psi{\check{e}}_{I}(\Phi-\Psi\ln\theta)=\theta\Psi{\check{e}}_{I}(\Phi)-\theta{\check{e}}_{I}(\ln\theta)\Psi^{2}-\tfrac{\theta\ln\theta}{2}{\check{e}}_{I}\big(\Psi^{2}\big).$
Hence, after multiplying by $1/\theta$, the momentum constraint becomes
${\check{e}}_{I}(p_{I})+{\check{e}}_{I}(\ln\theta)\big(\textstyle\sum_{J}p_{J}^{2}+\Psi^{2}-1\big)+\frac{1}{2}\ln\theta{\check{e}}_{I}\big(\textstyle\sum_{J}p_{J}^{2}+\Psi^{2}\big)+\textstyle\sum_{J}(p_{I}-p_{J})\Lambda_{JI}^{J}=\Psi{\check{e}}_{I}(\Phi).$
Now, after multiplication by $\theta^{-2}$, the Hamiltonian constraint reads
$\sum_{J}p_{J}^{2}+\Psi^{2}-1=\theta^{-2}(\mathrm{Scal}_{h}-|{\mathrm{d}}\varphi|_{h}^{2}-2V\circ\varphi).$
Thus, by the assumptions and Lemmas 41 and 46,
$\|\textstyle\sum_{J}p_{J}^{2}+\Psi^{2}-1\|_{C^{k_{0}-1}(\Sigma)}\leq Ct^{\delta}.$
Going back to the momentum constraint, we can now let $t\downarrow 0$ to conclude that
$\mathring{e}_{I}(\mathring{p}_{I})+\textstyle\sum_{J}(\mathring{p}_{I}-\mathring{p}_{J})\mathring{\Lambda}_{JI}^{J}=\mathring{\Psi}\mathring{e}_{I}(\mathring{\Phi}).$
Since the left-hand side of this equation coincides with $\mathrm{div}_{\mathring{\mathcal{H}}}\mathring{\mathcal{K}}(\mathring{e}_{I})$, the lemma follows. ∎
Proof of Theorem 24.
The statement follows essentially immediately from Theorem 56. The only difference is that we do not assume the existence of an eigenframe in the statement of Theorem 24. However, this is easily achieved by taking a finite covering space; cf. the comments at the beginning of Subsection 3.2. On this covering space, the conclusions of Theorem 56 hold. Moreover, since the only ambiguity in the choice of frame is a sign, the objects introduced in (134) descend to the quotient. Moreover, the conclusions of Theorem 24 follow from the conclusions of Theorem 56. ∎
4. Proof of the main result
At this point, we are ready to prove Theorem 11 by combining the main theorem of [33] with Theorems 19 and 24.
Proof of Theorem 11.
Given admissibility thresholds $\sigma_{V},\sigma_{p}\in(0,1)$ as in the statement of the theorem, $\sigma$ is defined as in [33, (11), p. 6]:
(135) $\sigma=\min\big(\tfrac{\sigma_{V}}{3},\tfrac{\sigma_{p}}{5}\big).$
In [33, (12), p. 6], lower bounds on $k_{0}$ and $k_{1}$ are specified. Here we let
$k_{0}:=\left\lceil\tfrac{n+1}{2}\right\rceil;$
i.e., $k_{0}$ is the minimal number such that [33, (12a), p. 6] holds. Note, however, that $k_{0}\geq 2$. We also assume that $k_{1}$ is the minimal integer such that [33, (12b), p. 6] holds. In particular, $\sigma$ is determined by $\sigma_{V}$ and $\sigma_{p}$; $k_{0}$ is determined by $n$; and $k_{1}$ is determined by $\sigma_{V}$, $\sigma_{p}$ and $n$. We define $\kappa:=k_{1}+2$.
Next, note that the assumptions include the requirement that the eigenvalues of $\mathcal{K}_{0}$ be distinct. Due to the arguments presented in the proof of [42, Lemma A.1, p. 223], it follows that a finite (at most a $2^{n}$-fold) covering space of $\Sigma$ has a global frame. Going to this finite cover, we can assume that there is a global orthonormal frame $\{E_{i}\}_{i=1}^{n}$ of $(\Sigma,h_{\mathrm{ref}})$. Considering the Sobolev norms on the left hand side of (8), they are at most increased by a factor of $2^{n}$ when going to the finite cover. Moreover, the norms used in the statement of [33, Theorem 12, pp. 6-7] are defined using a frame $\{E_{i}\}_{i=1}^{n}$ and the norms in (8) are defined using the metric $h_{\mathrm{ref}}$. However, the two norms are equivalent up to constants depending only on $(\Sigma,h_{\mathrm{ref}})$ and the frame. In other words, this difference does not affect the statement. Fixing a $\zeta_{0}$ as in the statement of the theorem, there is a $\zeta_{0}$ as in the statement of [33, Theorem 12, pp. 6-7] (obtained by compensating for the covering space and the equivalence of the norms), say $\zeta_{0}^{\prime}$. Due to [33, Theorem 12, pp. 6-7], we then obtain a $\zeta_{1}$. Given that the assumptions of the theorem are satisfied with this $\zeta_{1}$ (and the original $\zeta_{0}$), the conclusions of [33, Theorem 12, pp. 6-7] hold. Moreover, the estimates derived in the proof of this theorem hold.
Due to [33, Proposition 74, p. 29], there is a constant $\rho_{0}>0$ depending only on $\zeta_{0}$, $\sigma$, $k_{0}$, $k_{1}$, $(\Sigma,h_{\mathrm{ref}})$ and $\{E_{i}\}_{i=1}^{n}$. In our setting, $\rho_{0}$ thus only depends on $\zeta_{0}$, $\sigma_{V}$, $\sigma_{p}$, $(\Sigma,h_{\mathrm{ref}})$ and $\{E_{i}\}_{i=1}^{n}$. In accordance with [33, Notation 75, p. 29], we then say that $C$ is a standard constant if it is strictly positive and only depends on $\rho_{0}$, $\sigma_{p}$, $\sigma_{V}$, $k_{0}$, $k_{1}$, $c_{k_{1}+2}$ (see Definition 9), $(\Sigma,h_{\mathrm{ref}})$ and the orthonormal frame $\{E_{i}\}_{i=1}^{n}$.
Next, we wish to appeal to Theorem 19. Note, therefore, that we have a solution to (11)–(19) as described in Subsection 1.4; the equations solved in [33] are the ones written down in [33, Proposition 70, pp. 23-24], and these equations coincide with (11)–(19). Next, note that (25) with $\varepsilon=\sigma$ is an immediate consequence of [33, (110), p. 37] and [33, Lemma 100, p. 39]; recall [33, Definition 87, p. 36]. The estimate (26) is an immediate consequence of [33, (147), p. 49]. Note also that $3\sigma\leq\sigma_{V}$ due to (135). This means that the conclusions of Theorem 19 hold.
Next, we wish to combine the above conclusions with Theorem 24. From [33, Proposition 70, pp. 23-24], it follows that $g$ takes the form (29) and that (32) holds. Next, let $\mathring{\Phi}$ and $\mathring{\Psi}$ be as in the statement of Theorem 19 and let $\mathring{p}_{I}$ be the eigenvalues of $\mathring{\mathfrak{K}}$ (the matrix with components $\mathring{\mathfrak{K}}_{IJ}$). Then the $\mathring{p}_{I}$ are distinct due to [33, Theorem 12, pp. 6-7] and smooth since the $\mathring{\mathfrak{K}}_{IJ}$ are smooth. Due to [33, (14), p. 7], it is clear that (33) holds. Next, due to [33, (226), p. 70], we can, by demanding that $\zeta_{1}$ be large enough, ensure that $p_{I}+p_{J}-p_{K}\leq 1-\sigma_{p}/2$ for all $t\leq t_{0}$ and $I$, $J$, $K$ such that $I\neq J$. This means that (34) holds for any $\delta\leq\sigma$. Next, due to (28), the estimate (35) holds for all $k_{0}$, assuming $\delta<\sigma/4$. Assuming $\delta\leq\sigma/3$, the first three estimates in (36) follow immediately from (27). The final estimate follows from (27) by demanding that $T$ be small enough. By demanding that $T$ be small enough, we can also assume that $|p_{I}-p_{J}|>r$ for $I\neq J$ and some $r>0$. Fixing such $r$, $T$ and $\delta<\sigma/4$, it follows that $(M,g,\varphi)$ is a solution to the Einstein–nonlinear scalar field equations satisfying the $(\delta,n,k_{0},r)$-assumptions for any $k_{0}\in\mathbb{N}$.
At this point we can appeal to Theorem 24. This yields the conclusion that there are smooth data on the singularity, as desired. Moreover, (9) holds. The remaining conclusions of the theorem follow from [33, Theorem 12, pp. 6-7]. This finishes the proof of the theorem, with one caveat: we have demonstrated the desired conclusions for the finite covering space. On the other hand, the covering maps are local isometries, so that the conclusions descend to the quotient. The theorem follows. ∎
Appendix A Notation, conventions, estimates and identities
A.1. Conventions concerning number systems
Throughout this article, we use the convention that $\mathbb{N}=\{1,2,\dots\}$ and that $\mathbb{N}_{0}=\{0,1,2,3,\dots\}$.
A.2. Symmetrisation and antisymmetrisation
Parentheses around indices in an expression denote symmetrisation and brackets denote antisymmetrisation. For example, in (13),
$2N^{-1}e_{[I}(Nk_{J]K})=-N^{-1}e_{I}(Nk_{JK})+N^{-1}e_{J}(Nk_{IK}).$
Moreover, in (14),
$\gamma_{K(IJ)}=\tfrac{1}{2}(\gamma_{KIJ}+\gamma_{KJI}).$
A.3. Conventions concerning the geometry
A spacetime is a time oriented Lorentz manifold. Let $(M,g)$ be a spacetime with Levi-Civita connection $\nabla$ and let $\Sigma\subset M$ a spacelike hypersurface with future pointing unit normal $e_{0}$ and induced metric $h$. Then the second fundamental form of $\Sigma$ is given by $k(X,Y):=g(\nabla_{X}e_{0},Y)$ for $X$ and $Y$ tangential to $\Sigma$. If $s$ is a symmetric covariant $\ell$-tensor field on $\Sigma$, $s^{\sharp}$ denotes the $(1,\ell-1)$–tensor field on $\Sigma$ obtained by raising one index with $h$. In particular, the Weingarten map is the $(1,1)$-tensor field defined by $K:=k^{\sharp}$. The mean curvature is defined by $\theta:=\mathrm{tr}K$.
Let $(M,g)$ be an n+1n+1–dimensional spacetime. Then $(M,g)$ is said to have a smooth foliation by spacelike hypersurfaces if there is a smooth nn–dimensional manifold $\Sigma$, an interval $\mathcal{I}$ and a diffeomorphism $\Xi:\Sigma\times\mathcal{I}\rightarrow M$ such that the hypersurfaces $\Sigma_{t}:=\Sigma\times\{t\}$ are spacelike with respect to $\Xi^{*}g$. The foliation is said to have lapse $N>0$ and vanishing shift vector field if
(136) $\Xi^{*}g=-N^{2}{\mathrm{d}}t\otimes{\mathrm{d}}t+h,$
where $h$ denotes the family of induced (Riemannian) metrics on the hypersurfaces $\Sigma_{t}$. When we are in this situation, we usually omit reference to the diffeomorphism $\Xi$ and identify $M$ with $\Sigma\times\mathcal{I}$. In particular, we normally replace $\Xi^{*}g$ on the left hand side of (136) by $g$. Next, note that $e_{0}:=N^{-1}\partial_{t}$ is the future pointing unit normal of the $\Sigma_{t}$. If $\{E_{i}\}_{i=1}^{n}$ is a smooth global frame for $\Sigma$, it can be extended to $\Sigma\times\mathcal{I}$ by requiring that $\partial_{t},E_{i}]=0$ for all $i=1,\ldots,n$. Let $\{\eta^{i}\}_{i=1}^{n}$ be the corresponding dual frame. Then, the normal Lie derivative of $K$ is given in terms of this frame by
$\mathcal{L}_{e_{0}}K=N^{-1}\partial_{t}(K^{i}_{\phantom{i}j})E_{i}\otimes\eta^{j};$
see, e.g., [42, Subsection A.2, pp. 225-227]. Moreover, the Levi-Civita connection of $h$ is denoted by $\overline{\nabla}$. We also use the notation
$\langle u,v\rangle_{h}:=h_{p_{1}r_{1}}\cdots h_{p_{k}r_{k}}h^{q_{1}s_{1}}\cdots h^{q_{\ell}s_{\ell}}u^{p_{1}\cdots p_{k}}_{q_{1}\cdots q_{\ell}}v^{r_{1}\cdots r_{k}}_{s_{1}\cdots s_{\ell}}$
for $(k,\ell)$-tensor fields $u$ and $v$ on $\Sigma_{t}$. Moreover,
$|u|_{h}^{2}:=\langle u,u\rangle_{h}.$
The notation $|u|_{g}^{2}$ is defined similarly, but now using the Lorentz metric $g$ and assuming $u$ to be a $(k,\ell)$-tensor field on $M$. Note, however, that $|u|_{g}^{2}$ need not be non-negative in this case.
A.4. Norms
Let $(\Sigma,h_{\mathrm{ref}})$ be a closed Riemannian manifold with an orthonormal frame $\{E_{i}\}_{i=1}^{n}$. Denote the dual frame by $\{\eta^{i}\}_{i=1}^{n}$. In this article, bold capitalised Latin indices denote elements of $\{1,\dots,n\}^{\ell}$ for some $\ell\in\mathbb{N}_{0}$. If $\mathbf{I}\in\{1,\dots,n\}^{\ell}$, we let $|\mathbf{I}|:=\ell$ and if $\mathbf{I}=(i_{1},\dots,i_{\ell})$, then $E_{\mathbf{I}}:=E_{i_{1}}\cdots E_{i_{\ell}}$. If $k\in\mathbb{N}_{0}$ and $\psi\in C^{\infty}(\Sigma)$, we define the $C^{k}$- and $H^{k}$-norms of $\psi$ on $\Sigma$ as follows
$\|\psi\|_{C^{k}(\Sigma)}$ $:=\textstyle{\sum}_{|\mathbf{I}|\leq k}\sup_{x\in\Sigma}|(E_{\mathbf{I}}\psi)(x)|,$
$\|\psi\|_{H^{k}(\Sigma)}$ $:=\big(\textstyle{\sum}_{|\mathbf{I}|\leq k}\int_{\Sigma}|E_{\mathbf{I}}\psi|^{2}{\mathrm{d}}\mu_{h_{\mathrm{ref}}}\big)^{1/2}.$
If $u$ is a tensor field on $\Sigma$, the $C^{k}$-norm of $u$ on $\Sigma$ is defined as the sum of the $C^{k}$-norms of the components of $u$, with respect to the frame $\{E_{i}\}_{i=1}^{n}$. Just as in [33, (71), p. 28], we use the notation
(137a) $\|e\|_{C^{k}(\Sigma)}$ $:=\textstyle{\sum}_{I,i}\|e_{I}^{i}\|_{C^{k}(\Sigma)},$
(137b) $\|\gamma\|_{H^{k}(\Sigma)}$ $:=\big(\textstyle{\sum}_{I,J,K}\|\gamma_{IJK}\|_{H^{k}(\Sigma)}^{2}\big)^{1/2}$
and similarly for $\omega^{I}_{i}$ and $k_{IJ}$. We also use the notation
$\|\vec{e}\varphi\|_{C^{k}(\Sigma)}:=\textstyle{\sum}_{I}\|e_{I}\varphi\|_{C^{k}(\Sigma)},\ \ \ \|\vec{e}\varphi\|_{H^{k}(\Sigma)}:=\big(\textstyle{\sum}_{I}\|e_{I}\varphi\|_{H^{k}(\Sigma)}^{2}\big)^{1/2}$
and similarly when $\varphi$ is replaced by $N$.
A.5. Basic estimates and identities
Next, we recall the Gagliardo-Nirenberg estimates, which will be of central importance in what follows.
Lemma 57 (Gagliardo-Nirenberg estimates).
Let $(\Sigma,h_{\mathrm{ref}})$ be a closed Riemannian manifold and $\{E_{i}\}_{i=1}^{n}$ be a global orthonormal frame with respect to $h_{\mathrm{ref}}$. Let $\ell_{i}\in\mathbb{N}_{0}$, $i=1,\dots,j$, and $\ell=\ell_{1}+\dots+\ell_{j}$. If $|\mathbf{I}_{i}|=\ell_{i}$ and $\psi_{i}\in C^{\infty}(\Sigma)$, $i=1,\dots,j$, then
(138) $\|E_{\mathbf{I}_{1}}\psi_{1}\cdots E_{\mathbf{I}_{j}}\psi_{j}\|_{L^{2}(\Sigma)}\leq C\textstyle{\sum}_{i=1}^{j}\|\psi_{i}\|_{H^{\ell}(\Sigma)}\prod_{m\neq i}\|\psi_{m}\|_{C^{0}(\Sigma)},$
where $C$ only depends on $\ell$, $(\Sigma,h_{\mathrm{ref}})$ and $\{E_{i}\}_{i=1}^{n}$. In particular,
(139) $\|\psi_{1}\cdots\psi_{j}\|_{H^{\ell}(\Sigma)}\leq C\textstyle{\sum}_{i=1}^{j}\|\psi_{i}\|_{H^{\ell}(\Sigma)}\prod_{m\neq i}\|\psi_{m}\|_{C^{0}(\Sigma)}.$
Proof.
The first statement is a special case of [42, Corollary B.8]. The second statement is an immediate consequence of the first. ∎
We also need the following interpolation estimate.
Lemma 58 (Lemma 135, p. 74, [33]).
With $(\Sigma,h_{\mathrm{ref}})$ and $\{E_{i}\}_{i=1}^{n}$ as in Subsection 1.4, let $k$, $\ell\in\mathbb{N}_{0}$ be such that $0\leq\ell\leq k$ and $k\geq 1$. Then there is a constant $C$, depending only on $k$, $\ell$, $(\Sigma,h_{\mathrm{ref}})$ and $(E_{i})_{i=1}^{n}$, such that for all $\psi\in C^{\infty}(\Sigma)$,
(140) $\|\psi\|_{H^{\ell}(\Sigma)}\leq C\|\psi\|_{L^{2}(\Sigma)}^{1-\ell/k}\|\psi\|_{H^{k}(\Sigma)}^{\ell/k}.$
Finally, in the derivation of the energy estimates, we make use of the following elementary observation. If $\psi\in C^{\infty}(\Sigma)$ and $X$ is a smooth vector field on $\Sigma$, then
(141) ${\int}_{\Sigma}X(\psi){\mathrm{d}}\mu_{h_{\mathrm{ref}}}=-\textstyle{\int}_{\Sigma}\psi(\mathrm{div}_{h_{\mathrm{ref}}}X){\mathrm{d}}\mu_{h_{\mathrm{ref}}};$
see, e.g., [42, (B.7), p. 233].